### v1

# From Lipschitz Embedding to Lipschitz Equivalence between Dust-like Self-similar Sets

Abstract.
Let $K,F\subset\mathbb{R}^{d}$ be two dust-like self-similar sets sharing the same Hausdorff dimension. We consider when the mere existence of a Lipschitz embedding from $K$ to $F$ already implies their Lipschitz equivalence. Our main result is threefold: (1) if the Lipschitz image of $K$ intersects $F$ in a set of positive Hausdorff measure, then $K$ admits a Lipschitz surjection onto $F$; (2) if $F$ is in addition homogeneous, then the generating iterated function systems of $K,F$ should have algebraically dependent ratios and consequently, $K$ and $F$ are Lipschitz equivalent; (3) the Lipschitz equivalence can fail without the homogeneity assumption. This answers two questions in Balka and Keleti [Adv. Math. 446 (2024), 109669].
Key words and phrases:
Lipschitz equivalence, self-similar sets, Lipschitz embedding, algebraic dependence.
2020 Mathematics Subject Classification:
Primary 28A80; Secondary 28A78, 51F30
Corresponding author.
1. Introduction
Self-similar sets, which arise as attractors of self-similar iterated function systems (IFS), exhibit intricate geometric and topological properties that make them fundamental objects in the study of fractals. A key question in this context is to complete the Lipschitz classification of these sets, as Lipschitz equivalence preserves a number of crucial geometric and measure-theoretic properties, including fractal dimensions, regularity and metric scaling structures. Recall that $A,B\subset\mathbb{R}^{d}$ are said to be Lipschitz equivalent if there exists a bijection $f:A\to B$ such that
$c^{-1}|x-y|\leq|f(x)-f(y)|\leq c|x-y|,\quad\forall x,y\in A$
for some constant $c>0$, that is, both of $f$ and its inverse $f^{-1}$ are Lipschitz maps.
Unlike topological equivalence, the Lipschitz equivalence requires a bounded geometric distortion, making the classification far more rigid and heavily reliant on the precise local patterns of self-similar sets. A natural starting point is to examine dust-like self-similar sets, where the absence of overlaps yields a clear geometric structure. Recall that a self-similar set $K$ generated by an iterated function system (IFS) $\Phi=\{\varphi_{i}\}_{i=1}^{n}$, where each $\varphi_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a contractive similitude, is called dust-like when the images $\{\varphi_{i}(K)\}_{i=1}^{n}$ are pairwise disjoint. In this case, we also say that $\Phi$ satisfies the strong separation condition (SSC). For dust-like self-similar sets, it is proved that the Lipschitz equivalence reduces entirely to the algebraic compatibility of scaling ratios in their associated IFSs (see e.g. [15]), thus simplifying many technical challenges. The situation becomes particularly tractable ((see Lemma 2.4)) when dealing with homogeneous self-similar sets, where all similitudes in the associated IFS share a common contraction ratio.
One of the foundational studies toward this algebraic direction was conducted by Falconer and Marsh [8]. They demonstrated that the Lipschitz equivalence between two dust-like self-similar sets requires specific multiplicative relationships among the corresponding scaling ratios, thereby revealing a deep connection to the algebraic properties of the IFSs. This line was further developed in [14] where the first author, in collaboration with Rao and Wang, introduced a necessary condition for the Lipschitz equivalence called matchable condition, which enables a complete classification for two important cases: when the contraction vectors have full ranks, or when the IFSs consist of exactly two maps. Later, Rao and Zhang [17] linked the classification with a higher-dimensional Frobenius problem, yielding criteria for equivalence when the scaling ratios are coplanar in a certain sense. More recently, Xi and Xiong [22] gave a checkable characterization for systems with commensurable ratios. For related results, see [3, 4, 12, 20] and the references therein.
Another direction investigates whether a dust-like self-similar set admits a Lipschitz embedding into another, and how such embeddability relates to the Lipschitz equivalence. Since Lipschitz maps cannot increase Hausdorff dimension, if $K,F$ are two dust-like self-similar sets with $\operatorname{dim_{H}}K<\operatorname{dim_{H}}F$ (where $\operatorname{dim_{H}}$ denotes the Hausdorff dimension), then no Lipschitz maps can send $K$ onto $F$. On the other hand, if $\operatorname{dim_{H}}K>\operatorname{dim_{H}}F$, a nontrivial Lipschitz embedding of $K$ into $F$ always exists (see [2]). The critical case occurs when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$. For this scenario, Deng, Wen, Xiong and Xi [5] proved that $K$ admits a bi-Lipschitz embedding into $F$ if and only if these sets are Lipschitz equivalent.
In 2024, Balka and Keleti [2] made significant progress. They discovered that in certain circumstances, the mere existence of a Lipschitz surjection from one dust-like self-similar set to another already implies the Lipschitz equivalence. Their investigation focused particularly on the case when $K$ (as the domain) is homogeneous.
Theorem 1.1 ([2]).
Let $K,F\subset\mathbb{R}^{d}$ be two dust-like self-similar sets with the same Hausdorff dimension. If $K$ is homogeneous and there is a Lipschitz surjection from $K$ to $F$, then $K$ and $F$ are Lipschitz equivalent.
This result naturally raises the question of whether either the surjectivity condition or the homogeneity assumption can be weakened. More precisely, Balka and Keleti proposed the following two questions.
Question 1.2 ([2, Section 8]).
Let $K,F\subset\mathbb{R}^{d}$ be two dust-like self-similar sets with the same Hausdorff dimension $s$. Denote by $\mathcal{H}^{s}$ the $s$-dimensional Hausdorff measure.
  1. (1)
    If $K$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, must $K$ and $F$ be Lipschitz equivalent?
  2. (2)
    Under no homogeneity assumptions, if $K$ can be sent onto $F$ by a Lipschitz map, must $K$ and $F$ be Lipschitz equivalent?
The main aim of this paper is to address these questions. Our first result is as follows.
Theorem 1.3.
Let $K,F\subset\mathbb{R}^{d}$ be two dust-like self-similar sets with the same Hausdorff dimension $s$. If there is a Lipschitz map $f:K\to\mathbb{R}^{d}$ with $\mathcal{H}^{s}(f(K)\cap F)>0$, then $K$ admits a Lipschitz surjection onto $F$.
The construction of the desired surjection proceeds via three steps. First, we identify a regular local region within the intersection $f(K)\cap F$ that is not too condensed in an appropriate sense. Next, we apply a rescaling argument based on a standard density theorem in geometric measure theory. Finally, we establish the existence of the required map by passing to an appropriate limit. Combining this result with Theorem 1.1, we immediately obtain an affirmative answer to Question 1.2(1).
Corollary 1.4.
Let $K,F$ be as in Question 1.2. If $K$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K,F$ are Lipschitz equivalent.
Regarding the second question, our findings present a dichotomy of results. On the one hand, we prove that the Lipschitz equivalence holds when $F$ (as the range of the Lipschitz embedding) is homogeneous, establishing a converse to Balka and Keleti’s original theorem. On the other hand, we demonstrate that the homogeneity assumption on $K$ or $F$ is indeed essential: without it, there exists an explicit example (detailed in the last section) that provides a negative answer to Question 1.2(2). We summarize these results as follows.
Theorem 1.5.
Let $K,F\subset\mathbb{R}^{d}$ be two dust-like self-similar sets with the same Hausdorff dimension $s$.
  1. (1)
    If $F$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K,F$ are Lipschitz equivalent.
  2. (2)
    Assume that $K$ can be sent onto $F$ by a Lipschitz map. Without the homogeneity assumption, $K,F$ need not to be Lipschitz equivalent.
Here is an outline of the key steps of the proof. For part (1), it suffices to build the algebraic dependence between the corresponding contraction ratios (see Lemma 2.4). Note that Theorem 1.3 guarantees a Lipschitz surjection from $K$ to $F$. From this, we construct an alternative map which is almost injective and enjoys a crucial measure-preserving property. It is worth pointing out that the measure-preserving property is enough to derive the algebraic dependence when $K$ is homogeneous, thus providing an alternative way to recover Theorem 1.1. When $F$ is homogeneous instead, the argument admits a certain “reversal” (in some sense) once we prove that the image of every cell of $K$ under the alternative map contains a cell of $F$ of comparable size. For part (2), we rely on two powerful tools from [22]: a criterion of the equivalence in the commensurable case and a mass decomposition lemma. Using these, we explicitly construct a Lipschitz surjection between two dust-like self-similar sets that fail to be Lipschitz equivalent, thereby completing the proof.
Our results are also related to a notable conjecture by Feng, Huang and Rao [9], which states that an affine embedding from $K$ into $F$ may enforce the algebraic dependence between their associated ratios even without assuming that $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$. Significant progress has been made toward this conjecture, see [1, 6, 10]. In particular, a resolution of the special case when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$ was obtained as a corollary in an ongoing work [16] of the second author joint with Rao. The proof of Theorems 1.1, 1.3 and 1.5(1) reveals that if in addition that either $K$ or $F$ is homogeneous, the algebraic dependence follows already from the existence of a nontrivial Lipschitz embedding—a weaker condition than the original affinity requirement. However, as shown in [5], the ratios may remain independent without the dimensional restriction.
The paper is organized as follows. Section 2 introduces some necessary preliminaries. Section 3 establishes our notational conventions and presents the proof of Theorem 1.3. Building on these foundations, Sections 4 and 5 develop the construction of a well-behaved Lipschitz surjection from $K$ to $F$ and present the proof of Theorem 1.5(1). Finally, Section 6 provides an explicit example that demonstrates the second part of Theorem 1.5.
Notation
We write $X\lesssim Y$ when $X\leq CY$ for some constant $C>0$ and write $X\approx Y$ when $X\lesssim Y$ and $Y\lesssim X$ hold simultaneously. For $A,B\subset\mathbb{R}^{d}$, we denote by $\operatorname{diam}(A)$ the diameter of $A$ and write $\operatorname{dist}(A,B):=\inf\{|a-b|:a\in A,b\in B\}$.
2. Preliminaries
In this section we prove some useful preliminaries. We start with recalling the classical Lipschitz extension theorem of Kirszbraun.
Lemma 2.1 ([11]).
Let $E\subset\mathbb{R}^{n}$ and let $f:E\to\mathbb{R}^{m}$ be any Lipschitz map. Then $f$ can be extended to a Lipschitz map $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with $\operatorname{Lip}(g)=\operatorname{Lip}(f)$, where $\operatorname{Lip}(f)$ denotes the Lipschitz constant of $f$, that is,
$\operatorname{Lip}(f)=\sup\Big{\{}\frac{|f(x)-f(y)|}{|x-y|}:x\neq y\in E\Big{\}}.$
In what follows, we will fix two self-similar sets $K,F\subset\mathbb{R}^{d}$ with the same Hausdorff dimension $s$, generating by self-similar IFSs $\Phi=\{\varphi_{i}\}_{i\in I}$ and $\Psi=\{\psi_{j}\}_{j\in J}$ satisfying the SSC, respectively. Without loss of generality, we always assume that both of $K,F$ contain the origin and have diameter $1$.
One of the principal advantages of self-similarity is that it allows us, via an appropriate limiting process, to extract universal geometric or analytic features that persist across scales. A fundamental step in our subsequent arguments is as follows.
Lemma 2.2.
Let $P\geq 1$, $\gamma,L>0$ and let $u,v]\subset(0,\infty)$ be an interval. Set $\mathcal{K}=\mathcal{K}(P,\gamma,u,v)$ to be the collection of all sets of the form $\bigcup_{t=1}^{p}(c_{t}O_{t}K+a_{t})$, where $1\leq p\leq P$, $c_{t}\in[u,v]$, $O_{t}$ is a $d\times d$ orthogonal matrix and $a_{t}\in\mathbb{R}^{d}$, such that
$\min_{t\neq t^{\prime}}\operatorname{dist}(c_{t}O_{t}K+a_{t},c_{t^{\prime}}O_{t^{\prime}}K+a_{t^{\prime}})\geq\gamma.$
For any sequence $\{K_{n}\}_{n=1}^{\infty}\subset\mathcal{K}$ and Lipschitz maps $f_{n}:K_{n}\to F$ with $\sup_{n}\operatorname{Lip}(f_{n})\leq L$, there exists $E\in\mathcal{K}$ and a Lipschitz map $g:E\to F$ with $\operatorname{Lip}(g)\leq L$ such that
$\mathcal{H}^{s}(E)=\lim_{j\to\infty}\mathcal{H}^{s}(K_{n_{j}})\quad\text{and}\quad\mathcal{H}^{s}(g(E))\geq\limsup_{j\to\infty}\mathcal{H}^{s}(f_{n_{j}}(K_{n_{j}}))$
for some subsequence $\{n_{j}\}$.
Proof.
Since $K_{n}\in\mathcal{K}$, we can write it as $K_{n}=\bigcup_{t=1}^{p_{n}}(c_{n,t}O_{n,t}K+a_{n,t})$. Passing to a subsequence if necessary, we may assume that $p_{n}\equiv p$ for some $1\leq p\leq P$. Put $\widetilde{\gamma}:=v+L^{-1}+\gamma$ and $b_{t}:=(3t\widetilde{\gamma},0,\ldots,0)\in\mathbb{R}^{d}$ for $1\leq t\leq p$. For $n\geq 1$, it is easy to see that when $t\neq t^{\prime}$,
$\operatorname{dist}(c_{n,t}O_{n,t}K\,$ $+b_{t},c_{n,t^{\prime}}O_{n,t^{\prime}}K+b_{t^{\prime}})$
$\geq|b_{t}-b_{t^{\prime}}|-\operatorname{diam}(c_{n,t}O_{n,t}K+b_{t})-\operatorname{diam}(c_{n,t^{\prime}}O_{n,t^{\prime}}K+b_{t^{\prime}})$
$\geq 3\widetilde{\gamma}-v-v$
(2.1) $>\widetilde{\gamma}>\max\{L^{-1},\gamma\}.$
Shifting $f$ slightly on each piece, we obtain a new sequence of maps given by
$g_{n}:\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})$ $\to F$
$c_{n,t}O_{n,t}x+b_{t}$ $\mapsto f_{n}(c_{n,t}O_{n,t}x+a_{n,t}),$
which is well defined because the domain is a disjoint union (recall (2.1)). Clearly,
(2.2) $g_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\Big{)}=f_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+a_{n,t})\Big{)}=f_{n}(K_{n}).$
Fix any $n\geq 1$. Let $x,y\in\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})$, say $x=c_{n,t}O_{n,t}\alpha+b_{t}$ and $y=c_{n,t^{\prime}}O_{n,t^{\prime}}\beta+b_{t^{\prime}}$. If $t=t^{\prime}$ then
$|g_{n}(x)-g_{n}(y)|$ $=|f_{n}(c_{n,t}O_{n,t}\alpha+a_{n,t})-f_{n}(c_{n,t}O_{n,t}\beta+a_{n,t})|$
$\leq\operatorname{Lip}(f_{n})\cdot c_{n,t}|\alpha-\beta|$
$\leq L|x-y|.$
If $t\neq t^{\prime}$ then
$|g_{n}(x)-g_{n}(y)|$ $\leq 1$ (since $\operatorname{diam}(F)=1$)
$\leq L\operatorname{dist}(c_{n,t}O_{n,t}K+b_{t},c_{n,t^{\prime}}O_{n,t^{\prime}}K+b_{t^{\prime}})$ (by (2.1))
$\leq L|x-y|.$
Thus $g_{n}$ is Lipschitz as well with $\operatorname{Lip}(g_{n})\leq L$.
Passing to a subsequence and rearranging if necessary, we may assume that for every $1\leq t\leq p$, both of the sequences $\{c_{n,t}\}_{n},\{O_{n,t}\}_{n}$ are convergent, say $c_{n,t}\to c_{t}\in[u,v]$ and $O_{n,t}\to O_{t}$. Writing $E:=\bigcup_{t=1}^{p}(c_{t}O_{t}K+b_{t})$, we have by (2.1) and the convergence that $E\in\mathcal{K}$. Also,
$\mathcal{H}^{s}(E)=\sum_{t=1}^{p}c_{t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\sum_{t=1}^{p}c_{n,t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\mathcal{H}^{s}(K_{n}).$
To find a limit of $g_{n}$, we first pick a large $R>0$ so that $\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\subset B(0,R)$ for all $n\geq 1$. By Lemma 2.1, we can extend each $g_{n}$ to a Lipschitz map $\widetilde{g}_{n}:B(0,R)\to\mathbb{R}^{d}$ with $\operatorname{Lip}(\widetilde{g}_{n})=\operatorname{Lip}(g_{n})\leq L$. In particular, $\{\widetilde{g}_{n}\}$ is equicontinuous. Furthermore, since $\widetilde{g}_{n}(B(0,R))\cap F\neq\varnothing$ and
$\operatorname{diam}(\widetilde{g}_{n}(B(0,R)))\leq\operatorname{Lip}(\widetilde{g}_{n})\cdot\operatorname{diam}(B(0,R))\leq 2LR,$
$\widetilde{g}_{n}(B(0,R))$ is contained in the $2LR$-neighborhood of $F$ for all $n$, that is, $\{\widetilde{g}_{n}\}$ is pointwise bounded. Then by Arzelá-Ascoli theorem (e.g., see [19, Corollary 1.8.25]), there exists a subsequence $\{\widetilde{g}_{n_{j}}\}_{j=1}^{\infty}$ which converges uniformly to a limit function, say $g$. Clearly, $g$ is also Lipschitz with $\operatorname{Lip}(g)\leq L$.
Finally, since $\widetilde{g}_{n}$ extends $g_{n}$,
(2.3) $\widetilde{g}_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\Big{)}=g_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\Big{)}\subset F,\quad\forall n\geq 1.$
Then we have by the compactness of $F$ and all the convergences that
$g(E)=g\Big{(}\bigcup_{t=1}^{p}(c_{t}O_{t}K+b_{t})\Big{)}\subset F$
and for every $\varepsilon>0$,
$\widetilde{g}_{n_{j}}\Big{(}\bigcup_{t=1}^{p}(c_{n_{j},t}O_{n_{j},t}K+b_{t})\Big{)}\subset\mathcal{N}_{\varepsilon}(g(E))\cap F\quad\text{ for all large $j$},$
where $\mathcal{N}_{\varepsilon}(\cdot)$ denotes the open $\varepsilon$-neighborhood. Since $\mathcal{H}^{s}(F)<\infty$, we see that
(2.4) $\mathcal{H}^{s}(g(E))=\lim_{\varepsilon\to 0}\mathcal{H}^{s}(\mathcal{N}_{\varepsilon}(g(E))\cap F)$
and thus
$\mathcal{H}^{s}(g(E))$ $\geq\limsup_{j\to\infty}\mathcal{H}^{s}\Big{(}\widetilde{g}_{n_{j}}\Big{(}\bigcup_{t=1}^{p}(c_{{n_{j}},t}O_{{n_{j}},t}K+b_{t})\Big{)}\Big{)}$
$=\limsup_{j\to\infty}\mathcal{H}^{s}\Big{(}g_{n_{j}}\Big{(}\bigcup_{t=1}^{p}(c_{{n_{j}},t}O_{{n_{j}},t}K+b_{t})\Big{)}\Big{)}$
$=\limsup_{j\to\infty}\mathcal{H}^{s}(f_{n_{j}}(K_{n_{j}})).$ (by (2.2))
This completes the proof. ∎
Remark 2.3.
Under the assumptions of the above lemma, if $f_{n}(E_{n})$ is eventually dense in $F$, i.e., $f_{n}(E_{n})$ converges to $F$ under the Hausdorff distance $\operatorname{d_{H}}$, then $g(E)=F$. This is because $f_{n_{j}}(E_{n_{j}})=\widetilde{g}_{n_{j}}\big{(}\bigcup_{t=1}^{p}(c_{n_{j},t}O_{n_{j},t}K+b_{t})\big{)}\to g(E)$ under $\operatorname{d_{H}}$ (recall (2.2),  (2.3) and all the previous convergences).
Write $r_{i}$ (resp. $\lambda_{j}$) to be the contraction ratio of $\varphi_{i}$, $i\in I$ (resp. $\psi_{j}$, $j\in J$). When either $\Phi$ or $\Psi$ is homogeneous, the algebraic dependence of the corresponding ratios is in fact sufficient to establish the Lipschitz equivalence between $K$ and $F$. While this fact is implicitly contained in the work of Balka and Keleti [2], we state it explicitly as a separate lemma for the reader’s convenience.
Lemma 2.4.
Assume that $r_{i}\equiv r$ for some $0<r<1$. If $\frac{\log\lambda_{j}}{\log r}\in\mathbb{Q}$ for all $j\in J$, then $K,F$ are Lipschitz equivalent.
Proof.
Write $\lambda_{j}=r^{\alpha_{j}}$, where $\alpha_{j}\in\mathbb{Q}$ for $j\in J$. The SSC implies that $I\cdot r^{s}=1$ and $\sum_{j\in J}\lambda_{j}^{s}=1$. If $I=m^{k}$ for some integers $m,k>1$, then $K$ is Lipschitz equivalent to any self-similar set $E=\bigcup_{t=1}^{m}g_{t}(E)$ with the SSC such that each $g_{t}$ has similarity ratio $r^{1/k}$ (e.g., see [15]). So we may assume that $k=1$. Then $r^{s}=m^{-1}$ and thus
$1=\sum_{j\in J}\lambda_{j}^{s}=\sum_{j\in J}r^{\alpha_{j}s}=\sum_{j\in J}m^{-\alpha_{j}}.$
It follows from [2, Lemma 7.1] that $\alpha_{j}$ are all integers. Then by [2, Theorem 7.7], $K$ and $F$ are Lipschitz equivalent. ∎
3. Proof of Theorems 1.3
In this section, we present the proof of Theorem 1.3. Regarding $\Phi=\{\varphi_{i}\}_{i\in I}$ and $\Psi=\{\psi_{j}\}_{j\in J}$, we adopt the following customary notations in the rest of this paper.
  • $I^{*}:=\bigcup_{n=1}^{\infty}I^{n}$ and $K_{\bm{i}}:=\varphi_{\bm{i}}(K)$ for $\bm{i}\in I^{*}$;
  • $J^{*}:=\bigcup_{n=1}^{\infty}J^{n}$ and $F_{\bm{j}}:=\psi_{\bm{j}}(F)$ for $\bm{j}\in J^{*}$;
  • For $\bm{i}=i_{1}\cdots i_{n}\in I^{*}$, $\varphi_{\bm{i}}:=\varphi_{i_{1}}\circ\cdots\circ\varphi_{i_{n}}$ and $r_{\bm{i}}=r_{i_{1}}\cdots r_{i_{n}}$;
  • For $\bm{j}=j_{1}\cdots j_{n}\in J^{*}$, $\psi_{\bm{j}}:=\psi_{j_{1}}\circ\cdots\circ\psi_{j_{n}}$ and $\lambda_{\bm{j}}:=\lambda_{j_{1}}\cdots\lambda_{j_{n}}$;
  • $\underline{r}:=\min_{i}r_{i}$ and $\underline{\lambda}:=\min_{j}\lambda_{j}$;
  • $\wedge$: the operation between any two words in $I^{*}$ or $J^{*}$ which produces their longest common prefix;
  • For $\bm{i}\in I^{*}$ (resp. $\bm{j}\in J^{*}$), write $\bm{i}^{-}$ (resp. $\bm{j}^{-}$) to be the word obtained by removing the last digit of $\bm{i}$ (resp. $\bm{j}$);
  • $|\cdot|$: either the modulus of a vector in $\mathbb{R}^{d}$ or the length of a finite word in $I^{*}\cup J^{*}$;
  • $\Delta_{K}:=\min_{i\neq i^{\prime}}\operatorname{dist}(K_{i},K_{i^{\prime}})$ and $\Delta_{F}:=\min_{j\neq j^{\prime}}\operatorname{dist}(F_{j},F_{j^{\prime}})$, which are both positive numbers due to the SSC.
Our first observation is a reduction of the problem to its embedding case.
Lemma 3.1.
For any Lipschitz map $f:K\to\mathbb{R}^{d}$ with $f(K)\cap F\neq\varnothing$, there is a Lipschitz map $g:K\to F$ with $g(K)=f(K)\cap F$.
Proof.
We begin with constructing a map $g:K\to F$ with $g(K)=f(K)\cap F$. For any $\bm{i}\in I^{*}$ such that $f(K_{\bm{i}})\cap F=\varnothing$ but $f(K_{\bm{i}^{-}})\cap F\neq\varnothing$, we choose an arbitrary point $y_{\bm{i}}$ in $f(K_{\bm{i}^{-}})\cap F$ and define $g(K_{\bm{i}})=\{y_{\bm{i}}\}$. So if $x\in K$ is such that $f(x)\notin F$, then $g(x)$ is defined as $y_{\omega}$, where $\omega$ is the unique word in $I^{*}$ such that $x\in K_{\omega}$, $f(K_{\omega})\cap F=\varnothing$ and $f(K_{\omega^{-}})\cap F\neq\varnothing$. Note that the existence of $\omega$ is due to the compactness of $F$. If $f(x)\in F$ we simply set $g(x)=f(x)$. In this way, the map $g:K\to F$ is well defined and it is clear that $g(K)=f(K)\cap F$.
We claim that $\operatorname{diam}(g(K_{\bm{i}}))\leq\operatorname{diam}(f(K_{\bm{i}}))$ for all $\bm{i}\in I^{*}$. In fact, if $f(K_{\bm{i}})\cap F=\varnothing$, then $g(K_{\bm{i}})$ is a singleton and the inequality clearly holds. If $f(K_{\bm{i}})\cap F\neq\varnothing$, then by the definition of $g$,
$g(K_{\bm{i}})\subset\bigcup_{i\in I:f(K_{\bm{i}i})\cap F\neq\varnothing}f(K_{\bm{i}i})\subset f(K_{\bm{i}}),$
which also implies the desired inequality.
It remains to prove that $g$ is Lipschitz. For any distinct $x,x^{\prime}\in K$, let $\omega\in I^{*}$ be such that $|\omega|=\max\{|\bm{i}|:\bm{i}\in I^{*},x,x^{\prime}\in K_{\bm{i}}\}$. So
$|x-x^{\prime}|\geq\min_{i\neq i^{\prime}}\operatorname{dist}(K_{\omega i},K_{\omega i^{\prime}})=r_{\omega}\Delta_{K}=\Delta_{K}\operatorname{diam}(K_{\omega}).$
Then by the previous claim,
$|g(x)-g(x^{\prime})|\leq\operatorname{diam}(g(K_{\omega}))$ $\leq\operatorname{diam}(f(K_{\omega}))$
$\leq\operatorname{Lip}(f)\cdot\operatorname{diam}(K_{\omega})\leq\operatorname{Lip}(f)\cdot\Delta_{K}^{-1}|x-x^{\prime}|.$
So $g$ is Lipschitz. ∎
Based on this, the proof of Theorem 1.3 reduces to establishing the following result.
Theorem 3.2.
If there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K$ admits a Lipschitz surjection onto $F$.
In the rest of this section, let us fix such a Lipschitz embedding $f:K\to F$. For notational simplicity, write $L:=\operatorname{Lip}(f)$.
3.1. Localization for rescaling
Fix any small $0<\delta<1$ and set
(3.1) $\left\{\begin{array}[]{l}I_{n}:=\{\bm{i}\in I^{*}:r_{\bm{i}}\leq\delta^{n}<r_{\bm{i}^{-}}\},\\ J_{n}:=\{\bm{j}\in J^{*}:\lambda_{\bm{j}}\leq\delta_{n}<\lambda_{\bm{j}^{-}}\},\end{array}\right$.
where $\delta_{n}:=\frac{L}{\Delta_{F}}\cdot\delta^{n}$. The following lemma records several basic facts that will be used frequently later. Recall that $\underline{r}:=\min_{i\in I}r_{i}$ and $\underline{\lambda}:=\min_{j\in J}\lambda_{j}$.
Lemma 3.3.
Let $n\geq 1$.
  1. (1)
    We have $\underline{r}\delta^{n}<r_{\bm{i}}\leq\delta^{n}$ for $\bm{i}\in I_{n}$ and $\underline{\lambda}\delta_{n}<\lambda_{\bm{j}}\leq\delta_{n}$ for $\bm{j}\in J_{n}$. In particular, $I_{n}\approx\#J_{n}\approx\delta^{-ns}$, where the implicit constants are independent of $n$.
  2. (2)
    For every distinct $\bm{i},\bm{i}^{\prime}\in I_{n}$, $\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})>\Delta_{K}\delta^{n}$; for every distinct $\bm{j},\bm{j}^{\prime}\in J_{n}$, $\operatorname{dist}(F_{\bm{j}},F_{\bm{j}^{\prime}})>L\delta^{n}$.
  3. (3)
    For every $\bm{i}\in I_{n}$, there is a unique word $\bm{j}\in J_{n}$ such that $f(K_{\bm{i}})\subset F_{\bm{j}}$.
Proof.
The first statement follows directly from (3.1). For (2), note that
$\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})\geq r_{\bm{i}\wedge\bm{i}^{\prime}}\Delta_{K}\geq r_{\bm{i}^{-}}\Delta_{K}>\Delta_{K}\delta^{n};$
similarly,
$\operatorname{dist}(F_{\bm{j}},F_{\bm{j}^{\prime}})\geq\lambda_{\bm{j}^{-}}\Delta_{F}>\delta_{n}\Delta_{F}=\frac{L}{\Delta_{F}}\delta^{n}\Delta_{F}=L\delta^{n}.$
For (3), just note that when $\bm{i}\in I_{n}$,
$\operatorname{diam}(f(K_{\bm{i}}))\leq L\operatorname{diam}(K_{\bm{i}})=Lr_{\bm{i}}\leq L\delta^{n}<\min_{\bm{j}\neq\bm{j}^{\prime}\in J_{n}}\operatorname{dist}(F_{\bm{j}},F_{\bm{j}^{\prime}}),$
where the last inequality comes from (2). ∎
Lemma 3.4.
For $n\geq 1$ and $y\in F$, let $\bm{j}_{n}(y)$ be the unique word in $J_{n}$ such that $y\in F_{\bm{j}_{n}(y)}$. Then
$\lim_{n\to\infty}\frac{\mathcal{H}^{s}(f(K)\cap F_{\bm{j}_{n}(y)})}{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})}=1,\quad\mathcal{H}^{s}\text{-a.e. }y\in f(K).$
Proof.
Since the restriction of $\mathcal{H}^{s}$ on $F$ is a Radon measure, by a standard density theorem (see [13, Corollary 2.14]),
(3.2) $\lim_{\rho\to 0}\frac{\mathcal{H}^{s}(f(K)\cap B(y,\rho))}{\mathcal{H}^{s}(F\cap B(y,\rho))}=\lim_{\rho\to 0}\frac{\mathcal{H}^{s}(F\cap f(K)\cap B(y,\rho))}{\mathcal{H}^{s}(F\cap B(y,\rho))}=1$
for $\mathcal{H}^{s}$-a.e. $y\in f(K)$, where the first equality follows from $f(K)\subset F$. Fix any $y$ satisfying (3.2). For any $0<\varepsilon<1$, we have for all small $\rho$ that
(3.3) $\mathcal{H}^{s}(f(K)\cap B(y,\rho))\geq(1-\varepsilon)\mathcal{H}^{s}(F\cap B(y,\rho)).$
Pick the smallest $n$ so that $\delta_{n}\leq\rho$. Since $\bm{j}_{n}(y)\in J_{n}$, $\operatorname{diam}(F_{\bm{j}_{n}(y)})=\lambda_{\bm{j}_{n}(y)}\leq\delta_{n}$ and thus $F_{\bm{j}_{n}(y)}\subset B(y,\rho)$. On the other hand, by Lemma 3.3(1) and the minimality of $n$ (so $\delta_{n}\approx\rho$), there is a constant $C>0$ (independent of $y,n,\varepsilon$) such that
(3.4) $\mathcal{H}^{s}(F\cap B(y,\rho))\leq\sum_{\bm{j}\in J_{n}:F_{\bm{j}}\cap B(y,\rho)\neq\varnothing}\mathcal{H}^{s}(F_{\bm{j}})\leq C\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}).$
Combining all these observations and noticing that
$\big{(}F\cap B(y,\rho)\big{)}\setminus f(K)\supset\big{(}F_{\mathbf{j}_{n}(y)}\cap B(y,\rho)\big{)}\setminus f(K)=F_{\mathbf{j}_{n}(y)}\setminus f(K),$
we have
$\mathcal{H}^{s}(f(K)\cap B(y,\rho))$ $\leq\mathcal{H}^{s}(F\cap B(y,\rho))-\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))$
$=\mathcal{H}^{s}(F\cap B(y,\rho))-\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))}{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})}\cdot\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})$
$\leq\Big{(}1-C^{-1}\cdot\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))}{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})}\Big{)}\mathcal{H}^{s}(F\cap B(y,\rho)),$
where the last inequality follows from (3.4). Recalling (3.3), we immediately obtain
$\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))}{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})}\leq C\varepsilon.$
Since $\varepsilon$ is arbitrary, the proof is finished. ∎
To obtain an effective rescaling, we also need to identify a non-condensed local area of the image set $f(K)$.
Proposition 3.5.
There is an integer $Q\geq 1$ satisfying the following property. For any $0<\varepsilon<1$, we can find some $n$ and $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(F_{\bm{j}}\cap f(K))}{\mathcal{H}^{s}(F_{\bm{j}})}>1-\varepsilon$ and
$\{\bm{i}\in I_{n}:f(K_{\bm{i}})\subset F_{\bm{j}}\}=\#\{\bm{i}\in I_{n}:f(K_{\bm{i}})\cap F_{\bm{j}}\neq\varnothing\}\leq Q.$
Proof.
By Lemma 3.3(3), the first equality holds trivially since the two sets are identical. It remains to verify the cardinality. Let $0<\varepsilon<1$. By Lemma 3.4, there exists $E\subset f(K)$ with $\mathcal{H}^{s}(E)=\mathcal{H}^{s}(f(K))$ such that for all $y\in E$,
$\mathcal{H}^{s}(f(K)\cap F_{\bm{j}_{n}(y)})>(1-\varepsilon)\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})$
when $n$ is large enough. This naturally gives us a Vitali covering of $E$ consisting of specific cells of $F$. By a version of Vitali covering theorem (see [7, Theorem 1.10]), we can find $\{y_{t}\}_{t=1}^{\infty}\subset E$ and $\{n_{t}\}_{t=1}^{\infty}$ such that $\mathcal{H}^{s}(E\setminus\bigcup_{t=1}^{\infty}F_{\bm{j}_{n_{t}}(y_{t})})=0$, where $\{F_{\bm{j}_{n_{t}(y_{t})}}\}_{t}$ is a disjoint family and
(3.5) $\mathcal{H}^{s}(f(K)\cap F_{\bm{j}_{n_{t}}(y_{t})})>(1-\varepsilon)\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}(y_{t})}),\quad\forall t\geq 1.$
For convenience, below we simply write $\bm{j}_{n_{t}}:=\bm{j}_{n_{t}}(y_{t})$.
Pick any $p\geq 1$ such that
(3.6) $\mathcal{H}^{s}\Big{(}\bigcup_{t=1}^{p}F_{\bm{j}_{n_{t}}}\Big{)}\geq\frac{1}{2}\mathcal{H}^{s}(E)=\frac{1}{2}\mathcal{H}^{s}(f(K)).$
Pick an integer $P\geq 2\max\{n_{1},\ldots,n_{p}\}$ and a constant $C>\underline{\lambda}^{-2s}$. Set
(3.7) $\mathcal{M}_{t}:=\Big{\{}\eta\in J_{P}:F_{\eta}\subset F_{\bm{j}_{n_{t}}},\frac{\mathcal{H}^{s}(f(K)\cap F_{\eta})}{\mathcal{H}^{s}(F_{\eta})}>1-C\varepsilon\Big{\}},\quad 1\leq t\leq p$
and $\mathcal{M}:=\bigcup_{t=1}^{p}\mathcal{M}_{t}$. Since $F_{\bm{j}_{n_{1}}},\ldots,F_{\bm{j}_{n_{p}}}$ are distinct, $\mathcal{M}_{1},\ldots,\mathcal{M}_{p}$ are disjoint. We claim that $\mathcal{M}\gtrsim\delta_{N}^{-s}$.
To this end, fix any $1\leq t\leq p$. From $\bm{j}_{n_{t}}\in J_{n_{t}}$ and (3.5),
(3.8) $\varepsilon\delta_{n_{t}}^{s}\mathcal{H}^{s}(F)\geq\varepsilon\lambda_{\bm{j}_{n_{t}}}^{s}\mathcal{H}^{s}(F)=\varepsilon\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})>\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}}\setminus f(K)).$
On the other hand, by (3.7), for any $\eta\in J_{P}\setminus\mathcal{M}_{t}$ with $F_{\eta}\subset F_{\bm{j}_{n_{t}}}$, we have
$\mathcal{H}^{s}(F_{\eta}\setminus f(K))\geq C\varepsilon\mathcal{H}^{s}(F_{\eta})\geq C\varepsilon\cdot(\underline{\lambda}\delta_{P})^{s}\mathcal{H}^{s}(F).$
So
$\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}}\setminus f(K))$ $\geq\sum_{\eta\in J_{P}\setminus\mathcal{M}_{t}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}}\mathcal{H}^{s}(F_{\eta}\setminus f(K))$
$\geq\#\{\eta\in J_{P}\setminus\mathcal{M}_{t}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}\cdot C\varepsilon\cdot(\underline{\lambda}\delta_{P})^{s}\mathcal{H}^{s}(F).$
Combining this with (3.8), we see that
$\{\eta\in J_{P}\setminus\mathcal{M}_{t}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}\leq C^{-1}\underline{\lambda}^{-s}\delta_{n_{t}}^{s}\delta_{P}^{-s}$
and hence
$\#\mathcal{M}_{t}$ $=\#\{\eta\in J_{P}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}-\#\{\eta\in J_{P}\setminus\mathcal{M}_{t}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}$
$\geq\#\{\eta\in J_{P}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}-C^{-1}\underline{\lambda}^{-s}\delta_{n_{t}}^{s}\delta_{P}^{-s}$
$\geq\frac{\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})}{\max_{\eta\in J_{P}}\mathcal{H}^{s}(F_{\eta})}-C^{-1}\underline{\lambda}^{-s}\delta_{n_{t}}^{s}\delta_{P}^{-s}$
$\geq\frac{(\underline{\lambda}\delta_{n_{t}})^{s}}{\delta_{P}^{s}}-C^{-1}\underline{\lambda}^{-s}\delta_{n_{t}}^{s}\delta_{P}^{-s}$
(3.9) $=(\underline{\lambda}^{s}-C^{-1}\underline{\lambda}^{-s})\delta_{n_{t}}^{s}\delta_{P}^{-s}=:C_{1}\delta_{n_{t}}^{s}\delta_{P}^{-s}.$
It follows from our choice of $C$ that $C_{1}>0$. Since the above estimate holds for all $1\leq t\leq p$,
$\#\mathcal{M}$ $=\sum_{t=1}^{p}\#\mathcal{M}_{t}$ (by the disjointness)
$\geq C_{1}\delta_{N}^{-s}\sum_{t=1}^{p}\delta_{n_{t}}^{s}$ (by (3.9))
$=C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\delta_{n_{t}}^{s}\mathcal{H}^{s}(F)$
$\geq C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})$ (since $\lambda_{\bm{j}_{n_{t}}}\leq\delta_{n_{t}}$)
$\geq C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\cdot\frac{1}{2}\mathcal{H}^{s}(f(K))$ (by (3.6))
$=:C_{2}\delta_{P}^{-s},$
as claimed.
Now, by Lemma 3.3(3), $\{\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\}_{\eta\in\mathcal{M}}$ is a disjoint family. By the pigeonhole principle, there must exist a word $\eta\in\mathcal{M}$ such that
$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\leq\frac{\#I_{P}}{\#\mathcal{M}}.$
By Lemma 3.3(1), $I_{P}\approx\delta^{-Ps}$. Hence
$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\lesssim\frac{\delta^{-Ps}}{C_{2}\delta_{P}^{-s}}=\frac{\delta^{-Ps}}{C_{2}}\cdot\Big{(}\frac{L\delta^{P}}{\Delta_{F}}\Big{)}^{s}=\frac{L^{s}}{C_{2}\Delta_{F}^{s}},$
Combining this with (3.5), we prove the proposition. ∎
3.2. Proof of Theorem 3.2
Now let us prove Theorem 3.2. Applying Proposition 3.5 to $\varepsilon=n^{-1}$, we find $\{k_{n}\}_{n=1}^{\infty}\subset\mathbb{Z}^{+}$ and $\{\bm{j}_{n}\}_{n=1}^{\infty}$ (say $\bm{j}_{n}\in J_{k_{n}}$) such that
(3.10) $\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap f(K))>(1-n^{-1})\mathcal{H}^{s}(F_{\bm{j}_{n}})$
whereas
$P_{n}:=\#\{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\cap F_{\bm{j}_{n}}\neq\varnothing\}=\#\{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}\}$
are uniformly bounded. Passing to a subsequence if necessary, we may assume that $P_{n}\equiv P$ for some $P\in\mathbb{Z}^{+}$. Note that by Lemma 3.3(3), (3.10) becomes
(3.11) $\mathcal{H}^{s}\Big{(}F_{\bm{j}_{n}}\cap\bigcup_{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}}f(K_{\bm{i}})\Big{)}>(1-n^{-1})\mathcal{H}^{s}(F_{\bm{j}_{n}}).$
For convenience, enumerate $\{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}\}=:\{\bm{i}_{n,q}\}_{q=1}^{P}$.
With the notation, for every $n\geq 1$, $f$ sends each $K_{\bm{i}_{n,q}}$ into $F_{\bm{j}_{n}}=\psi_{\bm{j}_{n}}(F)$. This naturally generates a blow-up map
$f_{n}:\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})$ $\to F$
$\psi_{\bm{j}_{n}}^{-1}(x)$ $\mapsto\psi_{\bm{j}_{n}}^{-1}f(x).$
It follows that
$\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})\Big{)}\Big{)}$ $=\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,q}})\Big{)}$
$=\mathcal{H}^{s}(F)\cdot\lim_{n\to\infty}\frac{\mathcal{H}^{s}(F\cap\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,q}}))}{\mathcal{H}^{s}(F)}$
$=\mathcal{H}^{s}(F)\cdot\lim_{n\to\infty}\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\bigcup_{q=1}^{P}f(K_{\bm{i}_{n,q}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n}})}$
(3.12) $=\mathcal{H}^{s}(F),$
where the last equality follows from (3.11).
For $1\leq q\leq P$, write $\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})=:c_{n,q}O_{n,q}K+t_{n,q}$, where $c_{n,q}:=\lambda_{\bm{j}_{n}}^{-1}r_{\bm{i}_{n,q}}$. Recall that $\bm{j}_{n}\in J_{k_{n}}$ and $\bm{i}_{n,q}\in I_{k_{n}}$. By Lemma 3.3, it is easy to compute that $c_{n,q}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$ and
$\min_{q\neq q^{\prime}}\operatorname{dist}(\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}}),\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q^{\prime}}}))$ $=\lambda_{\bm{j}_{n}}^{-1}\min_{q\neq q^{\prime}}\operatorname{dist}(K_{\bm{i}_{n,q}},K_{\bm{i}_{n,q^{\prime}}})$
$>\lambda_{\bm{j}_{n}}^{-1}\Delta_{K}\delta^{k_{n}}$
$\geq L^{-1}\Delta_{K}\Delta_{F}.$
Applying Lemma 2.2, we can find a Lipschitz map $g:\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q})\to F$, where $c_{q}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$, $O_{q}$ is a $d\times d$ orthogonal matrix and $a_{q}\in\mathbb{R}^{d}$, such that $\operatorname{Lip}(g)\leq L$ and (recalling (3.12))
$\mathcal{H}^{s}\Big{(}g\Big{(}\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q})\Big{)}\Big{)}\geq\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})\Big{)}\Big{)}=\mathcal{H}^{s}(F).$
So $g(\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q}))$ is a closed subset of $F$ with full measure. As a result, $F=g(\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q}))$.
By Baire’s theorem, there is some $q$ such that $g(cO_{q}K+a_{q})$ contains an interior part of $F$, say $F_{\bm{j}}\subset g(cO_{q}K+a_{q})$. Thus, similarly as in the proof of Lemma 3.1, it is easy to construct a Lipschitz map sending $cO_{q}K+a_{q}$ onto $F_{\bm{j}}$. Since $cO_{q}K+a_{q}$ and $F_{\bm{j}}$ are nothing but scaled copies of $K$ and $F$, respectively, this establishes Theorem 3.2.
4. Lipschitz surjections with measure-preserving properties
In this section, we fix a Lipschitz embedding $f:K\to F$. By Theorem 3.2, we may assume $f$ to be surjective. Again write $L:=\operatorname{Lip}(f)$ for simplicity. A key insight, noted in [3, 21], asserts that any bi-Lipschitz map between $K$ and $F$ (if there is any) should preserve the $s$-dimensional Hausdorff measure in some local area. See [14, Lemma 2.4] for a short proof. While the given map $f$ may fall short of this ideal, we can nevertheless produce an alternative Lipschitz surjection from $K$ to $F$ that enjoys some nice measure linearity.
Let $\mathcal{K}:=\mathcal{K}(P,\gamma,u,v)$ be the collection as in Lemma 2.2 with $u,v]:=[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$, $P:=\lfloor\frac{L^{s}}{\underline{r}^{s}\Delta_{F}^{s}}\rfloor$ and $\gamma:=L^{-1}\Delta_{F}\Delta_{K}$.
Lemma 4.1.
The collection
$\{E\in\mathcal{K}:\exists\text{ a Lipschitz}\text{ surjection $g:E\to F$ with $\operatorname{Lip}(g)\leq L$}\}$
is non-empty.
Proof.
Let $c_{1}:=\min_{\bm{j}\in J_{1}}\frac{\mathcal{H}^{s}(f^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}$. Since $f(K)=F$,
(4.1) $\frac{\mathcal{H}^{s}(K)}{\mathcal{H}^{s}(F)}=\sum_{\eta\in J_{1}}\frac{\mathcal{H}^{s}(F_{\eta})}{\mathcal{H}^{s}(F)}\cdot\frac{\mathcal{H}^{s}(f^{-1}(F_{\eta}))}{\mathcal{H}^{s}(F_{\eta})}\geq c_{1}\sum_{\eta\in J_{1}}\frac{\mathcal{H}^{s}(F_{\eta})}{\mathcal{H}^{s}(F)}=c_{1}.$
Pick $\bm{j}\in J_{1}$ with $\frac{\mathcal{H}^{s}(f^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=c_{1}$. By Lemma 3.3(3), $f^{-1}(F_{\bm{j}})$ is a finite union $\bigcup_{\bm{i}\in\mathcal{I}}K_{\bm{i}}$ for some $\mathcal{I}\subset I_{1}$. Then we can blow up $f:\bigcup_{\bm{i}\in\mathcal{I}}K_{\bm{i}}\to F_{\bm{j}}$ and get a Lipschitz surjection
$g:\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}K_{\bm{i}}$ $\to F$
$\psi_{\bm{j}}^{-1}(x)$ $\mapsto\psi_{\bm{j}}^{-1}f(x)$
with $\operatorname{Lip}(g)\leq\operatorname{Lip}(f)=L$.
It remains to check that $\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}(K_{\bm{i}})\in\mathcal{K}$. By Lemma 3.3(1), we have
$\frac{\underline{r}\Delta_{F}}{L}\leq\lambda_{\bm{j}}^{-1}r_{\bm{i}}\leq\frac{\Delta_{F}}{\underline{\lambda}L},\quad\forall\bm{i}\in\mathcal{I}$
and
$\operatorname{dist}(\psi_{\bm{j}}^{-1}(K_{\bm{i}}),\psi_{\bm{j}}^{-1}(K_{\bm{i}^{\prime}}))=\lambda_{\bm{j}}^{-1}\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})\geq L^{-1}\Delta_{F}\Delta_{K}=\gamma,\quad\forall\bm{i}\neq\bm{i}^{\prime}\in\mathcal{I}.$
Moreover, since
$\#\mathcal{I}\cdot(\underline{r}\delta)^{s}\mathcal{H}^{s}(K)$ $\leq\#\mathcal{I}\cdot\min_{\bm{i}\in I_{1}}\mathcal{H}^{s}(K_{\bm{i}})$ (by Lemma 3.3(1))
$\leq\mathcal{H}^{s}\Big{(}\bigcup_{\bm{i}\in\mathcal{I}}K_{\bm{i}}\Big{)}$
$=\mathcal{H}^{s}(f^{-1}(F_{\bm{j}}))$
$=c_{1}\mathcal{H}^{s}(F_{\bm{j}})$
$\leq\frac{\mathcal{H}^{s}(K)}{\mathcal{H}^{s}(F)}\cdot\Big{(}\frac{L\delta}{\Delta_{F}}\Big{)}^{s}\mathcal{H}^{s}(F),$ (by (4.1) and Lemma 3.3(1))
we have $\mathcal{I}\leq\lfloor\frac{L^{s}}{\underline{r}^{s}\Delta_{F}^{s}}\rfloor=P$. So $\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}(K_{\bm{i}})\in\mathcal{K}$.
We remark that
(4.2) $\frac{\mathcal{H}^{s}(\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}(K_{\bm{i}}))}{\mathcal{H}^{s}(F)}=\frac{\mathcal{H}^{s}(\bigcup_{\bm{i}\in\mathcal{I}}K_{\bm{i}})}{\mathcal{H}^{s}(F_{\bm{j}})}=\frac{\mathcal{H}^{s}(f^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=c_{1}\leq\frac{\mathcal{H}^{s}(K)}{\mathcal{H}^{s}(F)},$
which will be used later. ∎
Corollary 4.2.
We can find $\widetilde{K}\in\mathcal{K}$ and a Lipschitz surjection $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$ such that
$\mathcal{H}^{s}(\widetilde{K})=\inf\{\mathcal{H}^{s}(E):E\in\mathcal{K},\exists\text{ a Lipschitz}$ $\text{ surjection $g:E\to F$ with $\operatorname{Lip}(g)\leq L$}\}.$
Moreover, $\mathcal{H}^{s}(\widetilde{K})\leq\mathcal{H}^{s}(K)$.
Proof.
The above lemma tells us that the infimum is well defined. Then one can take $\{E_{n}\}\subset\mathcal{K}$ and a sequence $\{g_{n}\}$, where each $g_{n}:E_{n}\to F$ is a Lipschitz surjection with $\operatorname{Lip}(g_{n})\leq L$, such that $\lim_{n}\mathcal{H}^{s}(E_{n})$ equals the infimum. By Lemma 2.2, we can find $\widetilde{K}\in\mathcal{K}$ and a Lipschitz map $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$ such that
$\mathcal{H}^{s}(\widetilde{K})=\lim_{j\to\infty}\mathcal{H}^{s}(E_{n_{j}})\quad\text{and}\quad\mathcal{H}^{s}(\widetilde{f}(K))\geq\limsup_{j\to\infty}\mathcal{H}^{s}(g_{n_{j}}(E_{n_{j}}))=\mathcal{H}^{s}(F)$
[…]
In the sequel, let us fix such an optimal domain $\widetilde{K}\in\mathcal{K}$, say $\widetilde{K}=\bigcup_{t=1}^{p}(a_{t}O_{t}K+b_{t})$, and a Lipschitz surjection $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$. Write $\widetilde{c}:=\frac{\mathcal{H}^{s}(\widetilde{K})}{\mathcal{H}^{s}(F)}$.
[…]
(4.3) $I^{\prime}_{t,n}:=\{\bm{i}\in I^{*}:a_{t}r_{\bm{i}}\leq\delta^{n}<a_{t}r_{\bm{i}^{-}}\},\quad 1\leq t\leq p.$
instead of $I_{n}$ (defined in (3.1)).
Lemma 4.3.
For every $n$ and $\bm{j}\in J_{n}$, $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for some $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$. Moreover, if $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$, then $\psi_{\bm{j}}^{-1}(\widetilde{f}^{-1}(F_{\bm{j}}))\in\mathcal{K}$ and $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$.
Proof.
Fix $n$ and $\bm{j}\in J_{n}$. Firstly, for each $1\leq t\leq p$, if $\bm{i}\in I^{\prime}_{t,n}$ then by Lemma 3.3(2),
$\operatorname{diam}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$ $\leq L\operatorname{diam}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
$=La_{t}r_{\bm{i}}\leq L\delta^{n}<\min_{\eta\neq\eta^{\prime}\in J_{n}}\operatorname{dist}(F_{\eta},F_{\eta^{\prime}}),$
that is, $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ should be contained in a unique $F_{\eta}$. Therefore, one can write
(4.4) $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
for some $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$.
Secondly, we show that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{m}\bigcup_{\bm{i}\in\mathcal{I}_{t}}\psi_{\bm{j}}^{-1}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ is a member of $\mathcal{K}$ provided that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$. Let $1\leq t,t^{\prime}\leq p$, $\bm{i}\in\mathcal{I}_{t}$ and $\bm{i}^{\prime}\in\mathcal{I}_{t^{\prime}}$. By Lemma 3.3(1), every ratio $\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$. On the other hand, if $t\neq t^{\prime}$ then since $\widetilde{K}\in\mathcal{K}$,
$\operatorname{dist}(\psi_{\bm{j}}^{-1}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$ $\psi_{\bm{j}}^{-1}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))$
$>\operatorname{dist}(a_{t}O_{t}K+b_{t},a_{t^{\prime}}O_{t^{\prime}}K+b_{t^{\prime}})\geq\gamma=L^{-1}\Delta_{F}\Delta_{K};$
while if $t=t^{\prime}$ and $\bm{i}\neq\bm{i}^{\prime}$ then
$\operatorname{dist}(\psi_{\bm{j}}^{-1}(a_{t}O_{t}$ $K_{\bm{i}}+b_{t}),\psi_{\bm{j}}^{-1}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))$
$=\lambda_{\bm{j}}^{-1}a_{t}\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})\geq\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}^{-}}\Delta_{K}\geq\frac{\Delta_{F}}{L\delta^{n}}\cdot\delta^{n}\cdot\Delta_{K}=L^{-1}\Delta_{F}\Delta_{K}$
as well.
Assume that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$. Then
$\Big{(}\sum_{t=1}^{p}\#\mathcal{I}_{t}\Big{)}\cdot(\underline{r}\delta^{n})^{s}\mathcal{H}^{s}(K)$ $\leq\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))$ (recall (4.4))
$\leq\widetilde{c}\mathcal{H}^{s}(F_{\bm{j}})$
$\leq\frac{\mathcal{H}^{s}(K)}{\mathcal{H}^{s}(F)}\cdot\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(F)$ (by Corollary 4.2)
$=\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(K)$
$\leq\frac{(L\delta^{n})^{s}}{\Delta_{F}^{s}}\mathcal{H}^{s}(K),$ (since $\bm{j}\in J_{n}$)
which implies that $\sum_{t=1}^{p}\#\mathcal{I}_{t}\leq\lfloor\frac{L^{s}}{\underline{r}^{s}\Delta_{F}^{s}}\rfloor=P$. So $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
Note that the inclusion $\widetilde{f}:\widetilde{f}^{-1}(F_{\bm{j}})\to F_{\bm{j}}$ naturally generates a Lipschitz surjection
$g:\psi_{\bm{j}}^{-1}(\widetilde{f}^{-1}(F_{\bm{j}}))$ $\to F$
$\psi_{\bm{j}}^{-1}(x)$ $\mapsto\psi_{\bm{j}}^{-1}\widetilde{f}(x)$
with $\operatorname{Lip}(g)\leq\operatorname{Lip}(\widetilde{f})\leq L$. Since $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$, it follows from our choice of $\widetilde{K}$ that
$\mathcal{H}^{s}(\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}}))\geq\mathcal{H}^{s}(\widetilde{K})=\widetilde{c}\mathcal{H}^{s}(F),$
which implies that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$ as desired. ∎
Proposition 4.4.
For all $n\geq 1$ and $\bm{j}\in J_{n}$, $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$ and $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
Proof.
Let $n\geq 1$. Since $\widetilde{f}:\widetilde{K}\to F$ is onto,
$\widetilde{c}=\frac{\mathcal{H}^{s}(\widetilde{K})}{\mathcal{H}^{s}(F)}=\sum_{\bm{j}\in J_{n}}\frac{\mathcal{H}^{s}(F_{\bm{j}})}{\mathcal{H}^{s}(F)}\cdot\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})},$
which is a convex combination. This identity should be $\widetilde{c}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}\widetilde{c}$. Otherwise, since every term in that combination is positive, there exists $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}<\widetilde{c}$, which contradicts Lemma 4.3. In particular, $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$ for every $\bm{j}\in J_{n}$. This in turn implies by Lemma 4.3 that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$. ∎
Remark 4.5.
It is standard to extend the above measure-preserving property to all Borel subsets of $F$, i.e., $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(B))}{\mathcal{H}^{s}(B)}=\widetilde{c}$ for all Borel $B\subset F$. This fact will be used later.
In other words, the Lipschitz surjection $\widetilde{f}$ is measure-preserving in some rough sense. This immediately yields the following corollary, recovering the aforementioned result of Balka and Keleti (Theorem 1.1).
Corollary 4.6 ([2]).
If $\Phi$ is homogeneous, then $K$ and $F$ are Lipschitz equivalent.
Proof.
Write $r$ to be the common ratio of maps in $\Phi$. By Lemma 2.4, it suffices to show that $\frac{\log\lambda_{j}}{\log r}\in\mathbb{Q}$ for $j\in J$. Let $j\in J$ and pick for $n\geq 1$ a positive integer $k_{n}$ so that
$j^{k_{n}}:=\underbrace{jj\cdots j}_{k_{n}\text{ terms}}\in J_{n}.$
Since $r_{i}\equiv r$, there exists $q_{t,n}\in\mathbb{N}$ such that $I^{\prime}_{t,n}\subset I^{q_{t,n}}$, where $I^{\prime}_{t,n}$ was defined in (4.3). Moreover, for $t\neq t^{\prime}$, we always have $a_{t}r^{q_{t,n}}\leq\delta^{n}<a_{t^{\prime}}r^{q_{t^{\prime},n}-1}$ and hence $r^{q_{t^{\prime},n}-q_{t,n}}>a_{t}r/a_{t^{\prime}}$. So there is a constant $C\in\mathbb{N}$ such that
(4.5) $|q_{t,n}-q_{t^{\prime},n}|\leq C,\quad\forall 1\leq t,t^{\prime}\leq m,\forall n\in\mathbb{Z}^{+}.$
By (4.4) and Proposition 4.4,
(4.6) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\Big{\{}\mathcal{H}^{s}(K)\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}:c_{n,t}\in\mathbb{N}\text{ with }\sum_{t=1}^{p}c_{n,t}\leq P\Big{\}}.$
Note that every inner sum on the right hand side above satisfies
$\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}$ $=r^{q_{1,n}s}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{(q_{t,n}-q_{1,n})s}$
$\in\Big{\{}r^{q_{1,n}s}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{\beta s}:-C\leq\beta\leq C\text{ is an integer}\Big{\}}.$ (by (4.5))
Combining this with (4.6), we see that
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\{\mathcal{H}^{s}(K)r^{q_{1,n}s}b:b\in\mathcal{B}\},\quad\forall n\geq 1,$
where
$\mathcal{B}:=\Big{\{}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{\beta s}:c_{n,t}\in\mathbb{N},\beta\in\mathbb{Z}\text{ are such that}\sum_{t=1}^{p}c_{n,t}\leq P\text{ and }-C\leq\beta\leq C\Big{\}}$
is a finite collection. Therefore, we can find (by the pigeonhole principle) a pair of distinct integers $n,\tilde{n}$ such that
$\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))}{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{\tilde{n}}}}))}=r^{(q_{1,n}-q_{1,\tilde{n}})s}$
whereas $q_{1,n}\neq q_{1,\tilde{n}}$ and $k_{n}\neq k_{\tilde{n}}$ (these automatically hold when |nn~||n-$\tilde${n}| is large enough). By Proposition 4.4, we also have
$\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))}{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{\tilde{n}}}}))}=\frac{\mathcal{H}^{s}(F_{j^{k_{n}}})}{\mathcal{H}^{s}(F_{j^{k_{\tilde{n}}}})}=\lambda_{j}^{(k_{n}-k_{\tilde{n}})s}.$
Thus $\frac{\log\lambda_{j}}{\log r}\in\mathbb{Q}$. ∎
Unfortunately, the above argument fails without the homogeneity of $\Phi$, and even granting the homogeneity of $\Psi$ seems to offer little help. This reveals a subtle asymmetry: to upgrade a Lipschitz embedding to a Lipschitz isomorphism, assuming the homogeneity on $K$ is markedly easier than assuming that on $F$. It is noteworthy that Proposition 4.4 may relate to a characterization given by Xi [20] via finding appropriate graph-directed subsystems, but we do not know whether the measure-preserving property would directly ensure that.
A natural step forward would be to ask how far the optimal map $\widetilde{f}$ deviates from being bi-Lipschitz. The next lemma shows that $\widetilde{f}$ is almost injective and locally nondegenerate.
Lemma 4.7.
Let $n\geq 1$, $\bm{j}\in J_{n}$ and write $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ (where $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$) as in Lemma 4.3. Then the following properties hold.
  1. (1)
    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0$ unless $(t,\bm{i})=(t^{\prime},\bm{i}^{\prime})$;
  2. (2)
    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\gtrsim\mathcal{H}^{s}(F_{\bm{j}})$ for all $1\leq t\leq m$ and $\bm{i}\in\mathcal{I}_{t}$.
Proof.
For (1), write $E:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})$ and suppose on the contrary that $\mathcal{H}^{s}(E)>0$. By Lemma 3.4, we can find $\eta_{k}\in J_{k}$ with $F_{\eta_{k}}\subset F_{\bm{j}}$ such that
(4.7) $\lim_{k\to\infty}\frac{\mathcal{H}^{s}(E\cap F_{\eta_{k}})}{\mathcal{H}^{s}(F_{\eta_{k}})}=1.$
For all large $k$, consider $A_{k}:=\widetilde{f}^{-1}(F_{\eta_{k}})$ and let
$A_{k,1}:=\widetilde{f}^{-1}(F_{\eta_{k}})\cap(a_{t}O_{t}K_{\bm{i}}+b_{t})\quad\text{and}\quad A_{k,2}:=\widetilde{f}^{-1}(F_{\eta_{k}})\cap(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}).$
Clearly, $A_{k,1}$, $A_{k,2}$ are both non-empty (since $\mathcal{H}^{s}(E\cap F_{\eta_{k}})>0$), mutually disjoint (since $(t,\bm{i})\neq(t^{\prime},\bm{i}^{\prime})$) and $A_{k,1}\cup A_{k,2}\subset A_{k}$. Moreover, it follows from (4.7) that
(4.8) $1\geq\lim_{k\to\infty}\frac{\mathcal{H}^{s}(\widetilde{f}(A_{k,1}))}{\mathcal{H}^{s}(F_{\eta_{k}})}=\lim_{k\to\infty}\frac{\mathcal{H}^{s}(F_{\eta_{k}}\cap\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))}{\mathcal{H}^{s}(F_{\eta_{k}})}\geq\lim_{k\to\infty}\frac{\mathcal{H}^{s}(F_{\eta_{k}}\cap E)}{\mathcal{H}^{s}(F_{\eta_{k}})}=1,$
that is, (4.8) is an equality.
On the other hand, by Lemma 4.3, we can write $\widetilde{f}^{-1}(F_{\eta_{k}})=\bigcup_{\ell=1}^{p}\bigcup_{\omega\in\mathcal{I}_{\ell,k}}(a_{\ell}O_{\ell}K_{\omega}+b_{\ell})$ for some $\mathcal{I}_{\ell,k}\subset I^{\prime}_{\ell,k}$. So
$A_{k,1}=\bigcup_{\omega\in\mathcal{I}_{t,k}:K_{\omega}\subset K_{\bm{i}}}(a_{t}O_{t}K_{\omega}+b_{t})\quad\text{and}\quad A_{k,2}=\bigcup_{\omega\in\mathcal{I}_{t^{\prime},k}:K_{\omega}\subset K_{\bm{i}^{\prime}}}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}}).$
By Proposition 4.4, $\psi_{\eta_{k}}^{-1}\widetilde{f}^{-1}(F_{\eta_{k}})\in\mathcal{K}$. So $\psi_{\eta_{k}}^{-1}(A_{k,1}),\psi_{\eta_{k}}^{-1}(A_{k,2})\in\mathcal{K}$ (because they are sub-unions). Also,
$\mathcal{H}^{s}(A_{k,2})$ $\geq\min_{\omega\in I^{\prime}_{t^{\prime},k}}\mathcal{H}^{s}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}})$
$\geq(\underline{r}\delta^{k})^{s}\mathcal{H}^{s}(K)$ (recall (4.3))
$\geq\frac{\underline{r}^{s}\Delta_{F}^{s}}{L^{s}}\cdot\lambda_{\eta_{k}}^{s}\mathcal{H}^{s}(K)$ (since $\eta_{k}\in J_{k}$)
$=\frac{\underline{r}^{s}\Delta_{F}^{s}}{L^{s}\mathcal{H}^{s}(F)}\mathcal{H}^{s}(F_{\eta_{k}})\mathcal{H}^{s}(K).$
Recalling that $A_{k,1}\cup A_{k,2}\subset A_{k}$ (which is a disjoint union) and $\frac{\mathcal{H}^{s}(A_{k})}{\mathcal{H}^{s}(F_{\eta_{k}})}=\widetilde{c}$ (Proposition 4.4), we have
(4.9) $\frac{\mathcal{H}^{s}(A_{k,1})}{\mathcal{H}^{s}(F_{\eta_{k}})}\leq\frac{\mathcal{H}^{s}(A_{k})-\mathcal{H}^{s}(A_{k,2})}{\mathcal{H}^{s}(F_{\eta_{k}})}\leq\widetilde{c}-\frac{\underline{r}^{s}\Delta_{F}^{s}\mathcal{H}^{s}(K)}{L^{s}\mathcal{H}^{s}(F)}<\widetilde{c}$.
Similarly as before, the restriction $\widetilde{f}:A_{k,1}\to F_{\eta_{k}}$ naturally generates a Lipschitz map
$g_{k}:\psi_{\eta_{k}}^{-1}(A_{k,1})$ $\to F$
$\psi_{\eta_{k}}^{-1}(x)$ $\mapsto\psi_{\eta_{k}}^{-1}\widetilde{f}(x).$
Since $\psi_{\eta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ and $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$, we can find by Lemma 2.2 a Lipschitz map $g:E\to F$ with $E\in\mathcal{K}$, $\operatorname{Lip}(g)\leq L$ such that
$\mathcal{H}^{s}(E)\leq\sup_{k}\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}(A_{k,1}))=\mathcal{H}^{s}(F)\cdot\sup_{k}\frac{\mathcal{H}^{s}(A_{k,1})}{\mathcal{H}^{s}(F_{\eta_{k}})}<\widetilde{c}\mathcal{H}^{s}(F)$
(where the last inequality is due to (4.9)) and
$\mathcal{H}^{s}(g(E))$ $\geq\limsup_{k\to\infty}\mathcal{H}^{s}(g_{k}(\psi_{\eta_{k}}^{-1}(A_{k,1})))$
$=\limsup_{k\to\infty}\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}\widetilde{f}(A_{k,1}))$ (by definition)
$=\mathcal{H}^{s}(F)\cdot\limsup_{k\to\infty}\frac{\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}\widetilde{f}(A_{k,1}))}{\mathcal{H}^{s}(F)}$
$=\mathcal{H}^{s}(F).$ (by (4.8))
The latter implies that $g(E)=F$. However, by our choice of $\widetilde{K}$ (recall Corollary 4.2),
$\widetilde{c}\mathcal{H}^{s}(F)>\mathcal{H}^{s}(E)\geq\mathcal{H}^{s}(\widetilde{K})=\widetilde{c}\mathcal{H}^{s}(F),$
which is a contradiction. This proves (1).
For (2), let $1\leq t\leq p$, $\bm{i}\in\mathcal{I}_{t}\subset I^{\prime}_{t,n}$ and let
$B:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\Big{(}\bigcup_{(\bm{i}^{\prime},t^{\prime})\neq(\bm{i},t)}\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})\Big{)},$
which is a closed set. We have seen in (1) that $\mathcal{H}^{s}(B)=0$. By Remark 4.5, $\mathcal{H}^{s}(\widetilde{f}^{-1}(B))=\widetilde{c}\mathcal{H}^{s}(B)=0$ and
$\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$ $=\widetilde{c}^{-1}\mathcal{H}^{s}\big{(}\widetilde{f}^{-1}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\big{)}$
$=\widetilde{c}^{-1}\mathcal{H}^{s}\big{(}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cup\widetilde{f}^{-1}(B)\big{)}$
$=\widetilde{c}^{-1}\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
$\gtrsim\mathcal{H}^{s}(F_{\bm{j}}),$
where the last estimate again comes from (4.3) and Lemma 3.3(1). This proves (2). ∎
5. Lipschitz equivalence via the homogeneity of $F$
In this section, we restrict our attention to the case when $\Psi$ is homogeneous, say $\lambda_{j}\equiv\lambda$ for $j\in J$, and establish Theorem 1.5(1). By Lemma 2.4, the proof again reduces to showing that $\frac{\log r_{i}}{\log\lambda}\in\mathbb{Q}$ for $i\in I$. The following key observation will enable us to, in some sense, interchange the roles of $K$ and $F$ in adapting the proof of Corollary 4.6.
Proposition 5.1.
There exists $\alpha\in\mathbb{Z}^{+}$ such that for all $1\leq t\leq m$, $n\geq 1$ and $\bm{i}\in I^{\prime}_{t,n}$, $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ contains at least one cell in $\{F_{\bm{j}}:\bm{j}\in J_{n+\alpha}\}$.
In other words, the Lipschitz surjection $\widetilde{f}$ is not distributed excessively fragmented in $F$. Since the proof of this proposition is rather involved, let us prove Theorem 1.5(1) upon it first and then check its validation in the next subsection.
5.1. From Proposition 5.1 to Theorem 1.5(1)
We first explain how Proposition 5.1 leads to the logarithmic commensurability. Assume Proposition 5.1 holds and let $i\in I$ be arbitrary.
For $n\geq 1$, let $k_{n}$ be the unique integer such that $i^{k_{n}}\in I^{\prime}_{1,n}$. So $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ (recall (4.3)). By Proposition 5.1, $\widetilde{f}(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})\supset F_{\bm{j}_{n}}$ for some $\bm{j}_{n}\in J_{n+\alpha}$. Again, we can write as in Lemma 4.3 that
(5.1) $\widetilde{f}^{-1}(F_{\bm{j}_{n}})=\bigcup_{t=1}^{m}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$
where $\mathcal{I}_{t}\subset I^{\prime}_{t,n+\alpha}$. Notice that by Lemma 4.7 and our choice of $\bm{j}_{n}$,
$\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0\;\textrm{ unless }(t^{\prime},\bm{i}^{\prime})=(1,i^{k_{n}}).$
Thus, from Remark 4.5,
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}})\cap(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0\;\textrm{ unless }(t^{\prime},\bm{i}^{\prime})=(1,i^{k_{n}}),$
which implies that
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}}))=\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}})\cap(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})).$
Combining this with (5.1), we have
(5.2) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}}))=\mathcal{H}^{s}\Big{(}\bigcup_{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}}(a_{1}O_{1}K_{\bm{i}}+b_{1})\Big{)}.$
Note that $\{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}\}$ can be written as $\{i^{k_{n}}\omega:\omega\in\Gamma_{n}\}$ for some $\Gamma_{n}\subset I^{*}$. By definition, $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ and $a_{1}r_{i}^{k_{n}}r_{\omega}\approx\delta^{n+\alpha}$ for all $\omega\in\Gamma_{n}$. So there exists a uniform integer $q$ such that $\Gamma_{n}\subset\bigcup_{k=1}^{q}I^{k}$. Then by the pigeonhole principle, we can find $n_{1}<n_{2}$ such that $\Gamma_{n_{1}}=\Gamma_{n_{2}}$. Since $\Psi$ is homogeneous, there are integers $z_{1},z_{2}$ such that $J_{n_{1}+\alpha}=J^{z_{1}}$ and $J_{n_{2}+\alpha}=J^{z_{2}}$. Recalling Proposition 4.4 and (5.2), we have
$\widetilde{c}=\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n_{\ell}}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(a_{1}O_{1}K_{i^{k_{n_{\ell}}}\omega}+b_{1})}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{a_{1}^{s}r_{i}^{k_{n_{\ell}}s}\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(K_{\omega})}{\lambda^{z_{1}s}\mathcal{H}^{s}(F)}$
for $\ell=1,2$. Since $\Gamma_{n_{1}}=\Gamma_{n_{2}}$, we have $r_{i}^{k_{n_{1}}-k_{n_{2}}}=\lambda^{z_{1}-z_{2}}$, i.e., $\frac{\log r_{i}}{\log\lambda}\in\mathbb{Q}$ (choosing $n_{1},n_{2}$ with a large difference, we can require that $z_{1}\neq z_{2}$). Since $i$ is arbitrary, this establish Theorem 1.5.
5.2. Proof of Proposition 5.1
The purpose of this subsection is to prove Proposition 5.1 by contradiction. The following landmark result of Szemerédi will be useful when establishing a denseness property later in our argument.
Lemma 5.2 ([18]).
Let $\gamma>0$. For every $k\geq 1$, there exists a threshold number $n(k)$ such that for $n\geq n(k)$, every subset of $\{1,2,\ldots,n\}$ with cardinality larger than $\gamma n$ contains a $k$-term arithmetic progression.
Since $\Psi$ is homogeneous, for each $n$ we can assign a unique integer $\tau(n)$ such that $J_{n}=J^{\tau(n)}$. For $1\leq t\leq m$ and $n\geq 1$, let
(5.3) $p(t,n):=\max\{p\geq 1:\exists\bm{i}\in I^{\prime}_{t,n}\text{ such that $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for all $\bm{j}\in J^{\tau(n)+p}$}\}$
Suppose on the contrary that Proposition 5.1 fails. Then $\max_{1\leq t\leq m}\sup_{n}p(t,n)=\infty$. Without loss of generality, assume that $\sup_{n}p(1,n)=\infty$.
Let $\gamma\in(0,1)$ be a fixed constant (will be specified later). For $k\geq(\#J)^{6}$, by Szemerédi’s theorem, there is a large integer $n_{k}$ such that any subset of $\{1,\ldots,(\#J)^{p(1,n_{k})}\}$ with cardinality $\geq\gamma(\#J)^{p(1,n_{k})}$ contains a $k$-term arithmetic progression. For notational simplicity, in the rest of the proof we write $p_{k}:=p(1,n_{k})$ and enumerate $J^{p_{k}}$ as $\{\eta_{k,t}:1\leq t\leq(\#J)^{p_{k}}\}$ according to the lexicographical order, that is,
$\eta_{k,1}=\underbrace{1\cdots 1}_{p_{k}\text{ terms}},\,\eta_{k,2}=\underbrace{1\cdots 1}_{p_{k}-1\text{ terms}}\!\!\!\!2,\,\ldots,\,\eta_{k,(\#J)^{p_{k}}}=\underbrace{(\#J)\cdots(\#J)}_{p_{k}\text{ terms}}.$
It will be helpful to note that for any $\bm{j}\in J^{*}$ with $|\bm{j}|\leq p_{k}$,
(5.4) $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}=\{z(\#J)^{p_{k}-|\bm{j}|}+1,z(\#J)^{p_{k}-|\bm{j}|}+2,\ldots,(z+1)(\#J)^{p_{k}-|\bm{j}|}\}$
for some integer $z\geq 0$, i.e., it is a (#J)(\#J)-adic interval of level-$(p_{k}-|\bm{j}|)$.
By (5.3), one can find a word $\bm{i}_{k}\in I^{\prime}_{1,n_{k}}$ with $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ for all $\bm{j}\in J^{\tau(n_{k})+p_{k}}$. On the other hand, recalling Lemma 3.3(3), $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\subset F_{\bm{j}_{k}}$ for some $\bm{j}_{k}\in J_{n_{k}}=J^{\tau(n_{k})}$. Since
$F_{\bm{j}_{k}}=\bigcup_{\theta\in J^{p_{k}}}F_{\bm{j}_{k}\theta}=\bigcup_{1\leq t\leq(\#J)^{p_{k}}}F_{\bm{j}_{k}\eta_{k,t}},$
we have
$\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ $=\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap\bigcup_{1\leq t\leq(\#J)^{p_{k}}}F_{\bm{j}_{k}\eta_{k,t}}$
(5.5) $=\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap\bigcup_{t\in T_{k}}F_{\bm{j}_{k}\eta_{k,t}},$
where $T_{k}:=\{1\leq t\leq(\#J)^{p_{k}}:\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap F_{\bm{j}_{k}\eta_{k,t}}\neq\varnothing\}$. Observe that
$\delta^{n_{k}s}$ $\lesssim\mathcal{H}^{s}(\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1}))$ (since $\bm{i}_{k}\in I^{\prime}_{1,n_{k}}$)
$\leq\mathcal{H}^{s}\Big{(}\bigcup_{t\in T_{k}}F_{\bm{j}_{k}\eta_{k,t}}\Big{)}$ (by (5.5))
$=\sum_{t\in T_{k}}\mathcal{H}^{s}(F_{\bm{j}_{k}\eta_{k,t}})$
$=\sum_{t\in T_{k}}\lambda_{\bm{j}_{k}}^{s}\mathcal{H}^{s}(F_{\eta_{k,t}})$
$\lesssim(\#T_{k})\cdot\delta^{n_{k}s}\cdot\lambda^{p_{k}s}\mathcal{H}^{s}(F)$ (since $\bm{j}_{k}\in J_{n_{k}}$ and $\eta_{k,t}\in J^{p_{k}}$)
and hence $T_{k}\gtrsim\lambda^{-p_{k}s}=(\#J)^{p_{k}}$. In summary, although every cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{p_{k}}\}$ is not entirely contained in $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$, the latter does intersect a large proportion of these cells.
Picking $0<\gamma<1$ properly at the beginning, we have $T_{k}\geq\gamma(\#J)^{p_{k}}$ for all $k$. As a consequence, $T_{k}$ should contain a $k$-term arithmetic progression, say $AP_{k}$. For the purpose of blowing up on a good sub-cell of $F_{\bm{j}_{k}}$, we perform the following truncation: pick $\theta_{k}\in J^{*}$ such that $F_{\bm{j}_{k}\theta_{k}}$ is a largest cell contained entirely in $\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\bm{j}_{k}\eta_{k,t}}$ and set
(5.6) $AP^{\prime}_{k}:=\{t\in AP_{k}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\},$
which is a subprogression of $AP_{k}$. By (5.4), $\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}$ is simply a #J\#J-adic interval of the form
(5.7) $\{z_{k}(\#J)^{p_{k}-|\theta_{k}|}+1,z_{k}(\#J)^{p_{k}-|\theta_{k}|}+2,\ldots,(z_{k}+1)(\#J)^{p_{k}-|\theta_{k}|}\}$
for some $z_{k}\geq 0$. Write $d_{k}$ to be the common difference of $AP_{k}$. Clearly,
(5.8) $AP^{\prime}_{k}\geq\frac{\#\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{d_{k}}-2=\frac{(\#J)^{p_{k}-|\theta_{k}|}}{d_{k}}-2.$
It is straightforward to obtain the following lower bound estimate of $AP_{k}^{\prime}$.
Lemma 5.3.
$AP^{\prime}_{k}\geq(\#J)^{-4}k$.
Proof.
Write $\ell_{k}:=\lfloor\frac{\log(k-1)d_{k}}{\log\#J}\rfloor$. Then
$\max AP_{k}-\min AP_{k}\geq(\#AP_{k}-1)d_{k}=(k-1)d_{k}\geq(\#J)^{\ell_{k}}>3\cdot(\#J)^{\ell_{k}-2}.$
Therefore, $\min AP_{k},\max AP_{k}]$ must contain a #J\#J-adic interval
$\{z(\#J)^{\ell_{k}-2}+1,\ldots,(z+1)(\#J)^{\ell_{k}-2}\}$
for some $z\geq 0$. By (5.4), this is just $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}$ for some $\bm{j}\in J^{p_{k}-(\ell_{k}-2)}$. In particular, $F_{\bm{j}}\subset\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\eta_{k,t}}$. Recalling our choice of $\theta_{k}$ (above (5.6)), we have $|\theta_{k}|\leq|\bm{j}|=p_{k}-(\ell_{k}-2)$ and hence
$\#AP^{\prime}_{k}$ $\geq\frac{(\#J)^{\ell_{k}-2}-2d_{k}}{d_{k}}$ (by (5.8))
$=(\#J)^{-2}\cdot(\#J)^{\lfloor\frac{\log(k-1)d_{k}}{\log\#J}\rfloor}\cdot d_{k}^{-1}-2$
$\geq(\#J)^{-3}(k-1)d_{k}\cdot d_{k}^{-1}-2$
$>k(\#J)^{-3}-3\geq k(\#J)^{-4},$
where the last inequality holds since $k\geq(\#J)^{6}$. ∎
To derive a contradiction, write
(5.9) $A_{k,1}:=\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}})\cap(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1}).$
Lemma 5.4.
The compact set $A_{k,1}$ satisfies the following two properties.
  1. (1)
    $\psi_{\bm{j}_{k}\theta_{k}}^{-1}\widetilde{f}(A_{k,1})$ converges to $F$ in the Hausdorff distance as $k\to\infty$.
  2. (2)
    There exists $0<c_{*}<1$ such that $\mathcal{H}^{s}(A_{k,1})\leq(1-c_{*})\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}}))$ for all $k$.
Proof.
For (1), it follows from Lemma 5.3 that
$d_{k}$ $\leq\frac{\#\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{\#AP^{\prime}_{k}-1}$ (by (5.6))
$=\frac{(\#J)^{p_{k}-|\theta_{k}|}}{\#AP^{\prime}_{k}-1}$ (by (5.7))
$\leq(\#J)^{p_{k}-|\theta_{k}|-\lfloor\frac{\log k}{\log\#J}\rfloor-c}$
[…]
$\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})\leq\lambda^{|\bm{j}_{k}|+p_{k}-\widetilde{p}_{k}}=\lambda_{\bm{j}_{k}}\cdot\lambda^{|\theta_{k}|+\lfloor\frac{\log k}{\log\#J}\rfloor+c}=\lambda_{\bm{j}_{k}\theta_{k}}\cdot\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor+c},$
where $\operatorname{d_{H}}$ denotes the Hausdorff distance. Then
$\operatorname{d_{H}}(\psi_{\bm{j}_{k}\theta_{k}}\widetilde{f}(A_{k,1}),F)=\lambda_{\bm{j}_{k}\theta_{k}}^{-1}\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})=\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor+c},$
which vanishes as $k\to\infty$. This proves (1).
For (2), recall that $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ does not contain any cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{p_{k}}\}$ and $|\theta_{k}|\leq p_{k}$. So there are $1\leq t\leq m$, $\bm{i}^{\prime}_{k}\in I^{\prime}_{t,n_{k}}$ with $(t,\bm{i}^{\prime}_{k})\neq(1,\bm{i}_{k})$ such that
$\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}}+b_{t})\cap F_{\bm{j}_{k}\theta_{k}}\neq\varnothing.$
Pick  [^1] For general $\delta$ the word $\bm{j}_{k}\theta_{k}$ might not belong to any $J_{n}$. But one can simply pick $\delta=\lambda$ at the beginning to fix this problem. an integer $q_{k}\in\mathbb{Z}^{+}$ so that $\bm{j}_{k}\theta_{k}\in J_{q_{k}}$. In the spirit of Lemma 4.3, there exists a word $\omega_{k}$ such that $\bm{i}^{\prime}_{k}\omega_{k}\in I^{\prime}_{t,q_{k}}$ and $\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\subset F_{\bm{j}_{k}\theta_{k}}$. Note that
(5.10) $\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\approx\delta^{q_{k}s}\approx\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})=\widetilde{c}^{-1}\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}})),$
where the last equality comes from Proposition 4.4. Then
$\mathcal{H}^{s}(A_{k,1})$ $=\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}})\cap(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1}))$ (by (5.9))
$\leq\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}}))-\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})$
$\leq(1-c_{*})\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}}))$ (by (5.10))
for some $0<c_{*}<1$ independent of $k$. This proves (2). ∎
Now we are ready to obtain a contradiction. Again, each Lipschitz inclusion $\widetilde{f}:A_{k,1}\to F_{\bm{j}_{k}\theta_{k}}$ naturally generates a blow-up map
$g_{k}:\psi_{\bm{j}_{k}\theta_{k}}^{-1}(A_{k,1})$ $\to F$
$\psi_{\bm{j}_{k}\theta_{k}}^{-1}(x)$ $\mapsto\psi_{\bm{j}_{k}\theta_{k}}^{-1}\widetilde{f}(x)$
with $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$. Similar to the second paragraph of the proof of Lemma 4.7, we have $\psi_{\bm{j}_{k}\theta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ for all $k$. Then, combining Lemma 2.2, Remark 2.3 and Lemma 5.4(2), we obtain a Lipschitz surjection $g:E\to F$ for some $E\in\mathcal{K}$ such that $\operatorname{Lip}(g)\leq L$ and
$\mathcal{H}^{s}(E)$ $\leq\sup_{k}\mathcal{H}^{s}(\psi_{\bm{j}_{k}\theta_{k}}^{-1}(A_{k,1}))$
$=\mathcal{H}^{s}(F)\cdot\sup_{k}\frac{\mathcal{H}^{s}(A_{k,1})}{\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})}$
$\leq\mathcal{H}^{s}(F)\cdot(1-c_{*})\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}}))}{\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})}$ (by Lemma 5.4(1))
[…]
which contradicts our choice of $\widetilde{K}$ (recall Lemma 4.2).
6. A counterexample in the inhomogeneous case
We finally present a concrete example supporting Theorem 1.5(2). Following [22], we say that $\Phi$ is commensurable if $\frac{\log r_{i}}{\log r_{i^{\prime}}}\in\mathbb{Q}$ for all $i,i^{\prime}\in I$ and call $r$ the ratio root, where $r$ is the unique number in $(0,1)$ such that the multiplicative group generated by $\{r_{i}:i\in I\}$ equals $r^{\mathbb{Z}}$. Let $\mu$ be the normalized measure of $\mathcal{H}^{s}$ restricted on $K$. In particular, $\mu(K_{\bm{i}})=r_{\bm{i}}^{s}$ for all $\bm{i}\in I^{*}$.
To state a mass decomposition lemma in [22], write $\mathbb{Z}[r^{s}]$ be the ring generated by $r^{s}$ and the integer set $\mathbb{Z}$. With a slight abuse of notation, here we pick $\delta=r$ and let
$\left\{\begin{array}[]{l}I_{n}:=\{\bm{i}\in I^{*}:r_{\bm{i}}\leq\delta^{n}<r_{\bm{i}^{-}}\},\\ J_{n}:=\{\bm{j}\in J^{*}:\lambda_{\bm{j}}\leq\delta^{n}<\lambda_{\bm{j}^{-}}\}.\end{array}\right$.
Lemma 6.1 ([22, Lemma 3.4]).
Assume that $\Phi$ is commensurable and $r$ is the associated ratio root. Let $\Omega\subset\mathbb{Z}[r^{s}]\cap(0,+\infty)$ be a finite set. For $k\geq 1$, write $\mathcal{B}_{k}$ be the collection of all finite union of cells in $\{K_{\omega}:\omega\in I_{k}\}$. Then there exists some $c\geq 1$ satisfying the following property: for any $k\geq 1$ and any set $E\in\mathcal{B}_{k}$, if $\mu(E)=b_{1}+\cdots+b_{n}$ for some $b_{1},\ldots,b_{n}\in r^{sk}\Omega$, then we can find a partition $E=\bigcup_{t=1}^{n}E_{t}$ such that every $E_{t}$ is a member in $\mathcal{B}_{k+c}$ with $\mu(E_{t})=b_{t}$.
The desired example is as follows.
Example 6.2.
Consider [22, Example 1.1], where $\Phi$ has ratios $r_{1}=r_{2}=3^{-1}$, $r_{3}=r_{4}=3^{-2}$ and $\Psi$ has ratios $\lambda_{1}=\cdots=\lambda_{20}=3^{-3}$, $\lambda_{21}=\cdots=\lambda_{28}=3^{-6}$. As stated there, $K$ and $F$ share an identical Hausdorff dimension $s$ but fail to be Lipschitz equivalent. However, by employing the above lemma, it is not hard to construct a Lipschitz surjection from $K$ to $F$. This process can be viewed as a hierarchical mass distribution operation.
For convenience, write $x:=3^{-s}$ and let $\nu$ be the normalized $\mathcal{H}^{s}$-measure on $F$. The dimension formula implies that $2x^{2}+2x=1$ and $8x^{6}+20x^{3}=1$. So in $\{F_{j}:j\in J\}$ there are $8$ cells of $\nu$-mass $x^{6}$ and $20$ cells of $\nu$-mass $x^{3}$. Since
$(2x^{2}+2x)^{3}=8x^{6}+8x^{3}+24x^{5}+24x^{4},$
in $\{K_{\bm{i}}:\bm{i}\in I^{3}\}$ there are $8$ cells of $\mu$-mass $x^{6}$, $8$ cells of $\mu$-mass $x^{3}$, $24$ cells of $\mu$-mass $x^{5}$ and $24$ cells of $\mu$-mass $x^{4}$. Noting that $2x^{5}+2x^{4}=x^{3}(2x^{2}+2x)=x^{3}$, we can perform the following decomposition based on the $\mu$-mass of cells in $\{K_{\bm{i}}:\bm{i}\in I^{3}\}$:
$\underbrace{\{x^{6}\},\ldots\{x^{6}\}}_{8\text{ terms}},\underbrace{\{x^{3}\},\ldots,\{x^{3}\}}_{8\text{ terms}},\underbrace{\{x^{5},x^{5},x^{4},x^{4}\},\ldots,\{x^{5},x^{5},x^{4},x^{4}\}}_{12\text{ terms}}.$
Thus there exist $8$ subgroups with $\mu$-mass $x^{6}$ and $20$ subgroups with $\mu$-mass $x^{3}$. This induces a natural partition of $I^{3}$ that admits a “mass-preserving” bijection to $J$. More precisely, we can decompose $I^{3}=\bigcup_{j\in J}\tilde{I}_{j}$ such that $\sum_{\bm{i}\in\tilde{I}_{j}}r_{\bm{i}}^{s}=\lambda_{j}^{s}$ for each $j\in J$.
In this example, the ratio root $r$ equals $3^{-1}$. So $\delta=3^{-1}$ and $J_{1}=J$. Write
$\Omega:=\{1,3^{-s},3^{-2s},\ldots,3^{-5s}\}.$
For such parameters $r$ and $\Omega$, there exists a large integer $c$ as specified in Lemma 6.1. We claim that for every $k\geq 1$, there exists a partition $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$ such that $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1+(k-1)c}$.
The claim can be proved by induction. Clearly, picking $c$ large enough at the beginning, every cell in $\{K_{\bm{i}}:\bm{i}\in I^{3}\}$ is a finite union of cells in $\{K_{\omega}:\omega\in I_{1+c}\}$. So we can transform the previous partition of $I^{3}$ into another partition $I_{1+c}=\bigcup_{\bm{j}\in J_{1}}\tilde{I}_{\bm{j}}$ with $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1}$. This settles the case $k=1$.
Suppose we have established the claim for some $k\geq 1$. By definition, $\lambda_{\eta}\leq 3^{-(1+kc)}<\lambda_{\eta^{-}}$ for each $\eta\in J_{1+kc}$. Since each map in $\Psi$ has ratio either $3^{-3}$ or $3^{-6}$, we have
$\lambda_{\eta}\in\{3^{-(1+kc)},3^{-(1+kc)-1},\ldots,3^{-(1+kc)-5}\},\quad\forall\eta\in J_{1+kc}$
and thus
(6.1) $\lambda_{\eta}^{s}\in 3^{-(1+kc)s}\{1,3^{-s},\ldots,3^{-5s}\}=3^{-(1+kc)s}\Omega=r^{(1+kc)s}\Omega,\quad\forall\eta\in J_{1+kc}.$
Let $\bm{j}\in J_{1+(k-1)c}$. Note that by the induction hypothesis, $\bm{j}$ corresponds to some $\tilde{I}_{\bm{j}}\subset I_{1+kc}$ with
$\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}=\sum_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\lambda_{\eta}^{s},$
where $\bm{j}\prec\eta$ means that $\bm{j}$ is a prefix of $\eta$. In other words, writing $\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}, EE$ is simply an element in $\mathcal{B}_{1+kc}$ (as in Lemma 5.4) with $\mu(E)=\sum_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\lambda_{\eta}^{s}$, where each $\lambda_{\eta}^{s}$ is a member of $r^{(1+kc)s}\Omega$ (recall (6.1)). It then follows from Lemma 6.1 that there exists a partition $E=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}$, where every $E_{\eta}$ is a member of $\mathcal{B}_{1+(k+1)c}$ with $\mu(E_{\eta})=\lambda_{\eta}^{s}$. By definition, there is some $\tilde{I}_{\eta}\subset I_{1+(k+1)c}$ such that $E_{\eta}=\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}$. We conclude that
(6.2) $\bigcup_{\omega\in\tilde{I}_{\bm{j}}}K_{\omega}=E=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa},$
where each equality represents a partition. It follows that
$\bigcup_{\kappa\in I_{1+(k+1)c}}K_{\kappa}=K$ $=\bigcup_{\omega\in I_{1+kc}}K_{\omega}$
$=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\omega\in\tilde{I}_{\bm{j}}}K_{\omega}$ (since $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$)
$=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}$ (by (6.2))
$=\bigcup_{\eta\in J_{1+kc}}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa},$
which in turn induces a partition (since each $\tilde{I}_{\eta}\subset I_{1+(k+1)c}$)
$I_{1+(k+1)c}=\bigcup_{\eta\in J_{1+kc}}\tilde{I}_{\eta}$
such that $\sum_{\kappa\in\tilde{I}_{\eta}}r_{\kappa}^{s}=\mu(E_{\eta})=\lambda_{\eta}^{s}$ for all $\eta\in J_{1+kc}$. This completes the induction.
The mass distribution in our claim naturally induces a Lipschitz surjection from $K$ to $F$ as follows. For $x\in K$ and $k\geq 1$, let $\bm{i}_{k}$ be the unique word in $I_{1+kc}$ such that $x\in K_{\bm{i}_{k}}$. Note that $\bm{i}_{1}\prec\bm{i}_{2}\prec\cdots$ and $\{x\}=\bigcap_{k}K_{\bm{i}_{k}}$. By our claim, there exists $\bm{j}_{k}\in J_{1+(k-1)c}$ such that $\bm{i}_{k}\in\tilde{I}_{\bm{j}_{k}}$. Moreover, it follows from (6.2) that
$K_{\bm{i}_{k+1}}\subset K_{\bm{i}_{k}}\subset\bigcup_{\omega\in\tilde{I}_{\bm{j}_{k}}}K_{\omega}=\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}.$
In other words, $\bm{i}_{k+1}\in\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\tilde{I}_{\eta}$. So $\bm{j}_{k}\prec\bm{j}_{k+1}$, that is, $\bm{j}_{1}\prec\bm{j}_{2}\prec\cdots$ as well. Then we just set $g(x)$ to be the single point in $\bigcap_{k}F_{\bm{j}_{k}}$. In this way, it is not hard to see that $g(K)=F$. It remains to prove that $g$ is Lipschitz.
Let $x,x^{\prime}\in K$ be distinct and let $k_{*}$ be the maximal integer such that $x,x^{\prime}\in K_{\omega_{*}}$ for a common $\omega_{*}\in I_{1+k_{*}c}$. So $|x-x^{\prime}|\approx\operatorname{diam}(K_{\omega_{*}})=r_{\omega_{*}}\approx 3^{-(1+k_{*}c)}$. Since $g$ sends $x,x^{\prime}$ into a specific cell in $\{F_{\bm{j}}:\bm{j}\in J_{1+(k_{*}-1)c}\}$, we have
$|g(x)-g(x^{\prime})|\leq\max\{\operatorname{diam}(F_{\bm{j}}):\bm{j}\in J_{1+(k_{*}-1)c}\}\approx 3^{-1-(k_{*}-1)c}\approx|x-x^{\prime}|.$
So $g$ is Lipschitz.
Acknowledgements. The research of the authors is partially supported by NSFC grant 12371089, and the Fundamental Research Funds for the Central Universities of China grant 2024FZZX02-01-01.

### v2

@@
 Self-similar sets, which arise as attractors of self-similar iterated function systems (IFS), exhibit intricate geometric and topological properties that make them fundamental objects in the study of fractals.
-A key question in this context is to complete the Lipschitz classification of these sets, as Lipschitz equivalence preserves a number of crucial geometric and measure-theoretic properties, including fractal dimensions, regularity and metric scaling structures.
+A key question in this context is to complete the Lipschitz classification of such sets, as Lipschitz equivalence preserves a number of crucial geometric and measure-theoretic properties, including fractal dimensions, regularity and metric scaling structures.
 Recall that $A,B\subset\mathbb{R}^{d}$ are said to be Lipschitz equivalent if there exists a bijection $f:A\to B$ such that
@@
 A natural starting point is to examine dust-like self-similar sets, where the absence of overlaps yields a clear geometric structure.
-Recall that a self-similar set $K$ generated by an iterated function system (IFS) $\Phi=\{\varphi_{i}\}_{i=1}^{n}$, where each $\varphi_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a contractive similitude, is called dust-like when the images $\{\varphi_{i}(K)\}_{i=1}^{n}$ are pairwise disjoint.
+Recall that a self-similar set $K$ generated by an iterated function system (IFS) $\Phi=\{\varphi_{i}\}_{i=1}^{n}$, where each $\varphi_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a contractive similitude, is called dust-like when $n\geq 2$ and the images $\{\varphi_{i}(K)\}_{i=1}^{n}$ are pairwise disjoint.
 In this case, we also say that $\Phi$ satisfies the strong separation condition (SSC).
@@
 [15]), thus simplifying many technical challenges.
-The situation becomes particularly tractable ((see Lemma 2.4)) when dealing with homogeneous self-similar sets, where all similitudes in the associated IFS share a common contraction ratio.
-One of the foundational studies toward this algebraic direction was conducted by Falconer and Marsh [8].
+The situation becomes particularly tractable (see Lemma 2.4) when dealing with homogeneous self-similar sets, where all similitudes in the associated IFS share a common contraction ratio.
+One of the foundational studies toward the algebraic direction was conducted by Falconer and Marsh [8].
 They demonstrated that the Lipschitz equivalence between two dust-like self-similar sets requires specific multiplicative relationships among the corresponding scaling ratios, thereby revealing a deep connection to the algebraic properties of the IFSs.
@@
 For related results, see [3, 4, 12, 20] and the references therein.
-Another direction investigates whether a dust-like self-similar set admits a Lipschitz embedding into another, and how such embeddability relates to the Lipschitz equivalence.
+Another direction investigates Lipschitz embeddings between dust-like self-similar sets and how such embeddability relates to the Lipschitz equivalence.
 Since Lipschitz maps cannot increase Hausdorff dimension, if $K,F$ are two dust-like self-similar sets with $\operatorname{dim_{H}}K<\operatorname{dim_{H}}F$ (where $\operatorname{dim_{H}}$ denotes the Hausdorff dimension), then no Lipschitz maps can send $K$ onto $F$.
-On the other hand, if $\operatorname{dim_{H}}K>\operatorname{dim_{H}}F$, a nontrivial Lipschitz embedding of $K$ into $F$ always exists (see [2]).
+On the other hand, if $\operatorname{dim_{H}}K>\operatorname{dim_{H}}F$, a Lipschitz embedding of $K$ onto $F$ always exists (see [2]).
 The critical case occurs when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$.
-For this scenario, Deng, Wen, Xiong and Xi [5] proved that $K$ admits a bi-Lipschitz embedding into $F$ if and only if these sets are Lipschitz equivalent.
+For this scenario, Deng, Wen, Xiong and Xi [5] proved that $K$ admits a bi-Lipschitz embedding into $F$ if and only if $K$ and $F$ are Lipschitz equivalent.
 In 2024, Balka and Keleti [2] made significant progress.
@@
 If there is a Lipschitz map $f:K\to\mathbb{R}^{d}$ with $\mathcal{H}^{s}(f(K)\cap F)>0$, then $K$ admits a Lipschitz surjection onto $F$.
-The construction of the desired surjection proceeds via three steps.
-First, we identify a regular local region within the intersection $f(K)\cap F$ that is not too condensed in an appropriate sense.
+The construction of such a surjection proceeds via three steps.
+We begin by determining a regular local region of $F$ where the map $f$ is not overly concentrated.
 Next, we apply a rescaling argument based on a standard density theorem in geometric measure theory.
-Finally, we establish the existence of the required map by passing to an appropriate limit.
+The desired surjection is then constructed by passing to an appropriate limit.
 Combining this result with Theorem 1.1, we immediately obtain an affirmative answer to Question 1.2(1).
 Corollary 1.4.
-Let $K,F$ be as in Question 1.2. If $K$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K,F$ are Lipschitz equivalent.
+Let $K,F$ be as in Question 1.2. If $K$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K$ and $F$ are Lipschitz equivalent.
 Regarding the second question, our findings present a dichotomy of results.
-On the one hand, we prove that the Lipschitz equivalence holds when $F$ (as the range of the Lipschitz embedding) is homogeneous, establishing a converse to Balka and Keleti’s original theorem.
-On the other hand, we demonstrate that the homogeneity assumption on $K$ or $F$ is indeed essential: without it, there exists an explicit example (detailed in the last section) that provides a negative answer to Question 1.2(2).
+On the one hand, we prove that the Lipschitz equivalence holds when $F$ (as the target space of the embedding) is homogeneous, establishing a converse to Balka and Keleti’s original theorem.
+On the other hand, we demonstrate that the homogeneity assumption on $K$ or $F$ is indeed essential: without it, there exists an explicit example (see Example 6.2) that provides a negative answer to Question 1.2(2).
 We summarize these results as follows.
@@
   1. (1)
-    If $F$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K,F$ are Lipschitz equivalent.
+    If $F$ is homogeneous and there is a Lipschitz map $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$, then $K$ and $F$ are Lipschitz equivalent.
   2. (2)
-    Assume that $K$ can be sent onto $F$ by a Lipschitz map. Without the homogeneity assumption, $K,F$ need not to be Lipschitz equivalent.
+    Assume that $K$ can be sent onto $F$ by a Lipschitz map. Without the homogeneity assumption, $K$ and $F$ need not to be Lipschitz equivalent.
 Here is an outline of the key steps of the proof.
@@
 When $F$ is homogeneous instead, the argument admits a certain “reversal” (in some sense) once we prove that the image of every cell of $K$ under the alternative map contains a cell of $F$ of comparable size.
-For part (2), we rely on two powerful tools from [22]: a criterion of the equivalence in the commensurable case and a mass decomposition lemma.
+For part (2), we rely on a powerful mass decomposition lemma in [22].
 Using these, we explicitly construct a Lipschitz surjection between two dust-like self-similar sets that fail to be Lipschitz equivalent, thereby completing the proof.
-Our results are also related to a notable conjecture by Feng, Huang and Rao [9], which states that an affine embedding from $K$ into $F$ may enforce the algebraic dependence between their associated ratios even without assuming that $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$.
+Our results are also related to a notable conjecture by Feng, Huang and Rao [9], which states that the existence of an affine embedding from $K$ into $F$ may enforce the algebraic dependence between their associated ratios even without assuming that $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$.
 Significant progress has been made toward this conjecture, see [1, 6, 10].
-In particular, a resolution of the special case when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$ was obtained as a corollary in an ongoing work [16] of the second author joint with Rao.
-The proof of Theorems 1.1, 1.3 and 1.5(1) reveals that if in addition that either $K$ or $F$ is homogeneous, the algebraic dependence follows already from the existence of a nontrivial Lipschitz embedding—a weaker condition than the original affinity requirement.
-However, as shown in [5], the ratios may remain independent without the dimensional restriction.
+In particular, a resolution of the special case when $\operatorname{dim_{H}}K=\operatorname{dim_{H}}F$ was recently obtained as a corollary in an ongoing work [16] of the second author joint with Rao.
+The proofs of Theorems 1.1, 1.3 and 1.5(1) reveal that if in addition that either $K$ or $F$ is homogeneous, the algebraic dependence follows already from the existence of a nontrivial Lipschitz embedding—a weaker condition than the affinity requirement.
+However, as shown in [5], such dependence is not guaranteed without the dimensional restriction.
 The paper is organized as follows.
@@
 Section 3 establishes our notational conventions and presents the proof of Theorem 1.3.
-Building on these foundations, Sections 4 and 5 develop the construction of a well-behaved Lipschitz surjection from $K$ to $F$ and present the proof of Theorem 1.5(1).
+Building on these foundations, Sections 4 and 5 develop the construction of a well-behaved Lipschitz surjection and present the proof of Theorem 1.5(1).
 Finally, Section 6 provides an explicit example that demonstrates the second part of Theorem 1.5.
@@
 2. Preliminaries
-In this section we prove some useful preliminaries. We start with recalling the classical Lipschitz extension theorem of Kirszbraun.
+We begin with a classical Lipschitz extension theorem of Kirszbraun.
 Lemma 2.1 ([11]).
@@
 $\min_{t\neq t^{\prime}}\operatorname{dist}(c_{t}O_{t}K+a_{t},c_{t^{\prime}}O_{t^{\prime}}K+a_{t^{\prime}})\geq\gamma.$
-For any sequence $\{K_{n}\}_{n=1}^{\infty}\subset\mathcal{K}$ and Lipschitz maps $f_{n}:K_{n}\to F$ with $\sup_{n}\operatorname{Lip}(f_{n})\leq L$, there exists $E\in\mathcal{K}$ and a Lipschitz map $g:E\to F$ with $\operatorname{Lip}(g)\leq L$ such that
-$\mathcal{H}^{s}(E)=\lim_{j\to\infty}\mathcal{H}^{s}(K_{n_{j}})\quad\text{and}\quad\mathcal{H}^{s}(g(E))\geq\limsup_{j\to\infty}\mathcal{H}^{s}(f_{n_{j}}(K_{n_{j}}))$
+For any sequence $\{E_{n}\}_{n=1}^{\infty}\subset\mathcal{K}$ and Lipschitz maps $f_{n}:E_{n}\to F$ with $\sup_{n}\operatorname{Lip}(f_{n})\leq L$, there exists $E\in\mathcal{K}$ and a Lipschitz map $g:E\to F$ with $\operatorname{Lip}(g)\leq L$ such that
+$\mathcal{H}^{s}(E)=\lim_{j\to\infty}\mathcal{H}^{s}(E_{n_{j}})\quad\text{and}\quad\mathcal{H}^{s}(g(E))\geq\limsup_{j\to\infty}\mathcal{H}^{s}(f_{n_{j}}(E_{n_{j}}))$
 for some subsequence $\{n_{j}\}$.
 Proof.
-Since $K_{n}\in\mathcal{K}$, we can write it as $K_{n}=\bigcup_{t=1}^{p_{n}}(c_{n,t}O_{n,t}K+a_{n,t})$.
+Since $E_{n}\in\mathcal{K}$, we can write it as $E_{n}=\bigcup_{t=1}^{p_{n}}(c_{n,t}O_{n,t}K+a_{n,t})$.
 Passing to a subsequence if necessary, we may assume that $p_{n}\equiv p$ for some $1\leq p\leq P$.
@@
 which is well defined because the domain is a disjoint union (recall (2.1)). Clearly,
-(2.2) $g_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\Big{)}=f_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+a_{n,t})\Big{)}=f_{n}(K_{n}).$
+(2.2) $g_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\Big{)}=f_{n}\Big{(}\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+a_{n,t})\Big{)}=f_{n}(E_{n}).$
 Fix any $n\geq 1$. Let $x,y\in\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})$, say $x=c_{n,t}O_{n,t}\alpha+b_{t}$ and $y=c_{n,t^{\prime}}O_{n,t^{\prime}}\beta+b_{t^{\prime}}$. If $t=t^{\prime}$ then
@@
 Also,
-$\mathcal{H}^{s}(E)=\sum_{t=1}^{p}c_{t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\sum_{t=1}^{p}c_{n,t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\mathcal{H}^{s}(K_{n}).$
+$\mathcal{H}^{s}(E)=\sum_{t=1}^{p}c_{t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\sum_{t=1}^{p}c_{n,t}^{s}\mathcal{H}^{s}(K)=\lim_{n\to\infty}\mathcal{H}^{s}(E_{n}).$
 To find a limit of $g_{n}$, we first pick a large $R>0$ so that $\bigcup_{t=1}^{p}(c_{n,t}O_{n,t}K+b_{t})\subset B(0,R)$ for all $n\geq 1$.
-By Lemma 2.1, we can extend each $g_{n}$ to a Lipschitz map $\widetilde{g}_{n}:B(0,R)\to\mathbb{R}^{d}$ with $\operatorname{Lip}(\widetilde{g}_{n})=\operatorname{Lip}(g_{n})\leq L$.
+By Lemma 2.1, each $g_{n}$ extends to a Lipschitz map $\widetilde{g}_{n}:B(0,R)\to\mathbb{R}^{d}$ with $\operatorname{Lip}(\widetilde{g}_{n})=\operatorname{Lip}(g_{n})\leq L$.
 In particular, $\{\widetilde{g}_{n}\}$ is equicontinuous.
@@
 $\widetilde{g}_{n}(B(0,R))$ is contained in the $2LR$-neighborhood of $F$ for all $n$, that is, $\{\widetilde{g}_{n}\}$ is pointwise bounded.
-Then by Arzelá-Ascoli theorem (e.g., see [19, Corollary 1.8.25]), there exists a subsequence $\{\widetilde{g}_{n_{j}}\}_{j=1}^{\infty}$ which converges uniformly to a limit function, say $g$.
+Then by Arzelá-Ascoli theorem (e.g., see [19, Corollary 1.8.25]), there exists a subsequence $\{\widetilde{g}_{n_{j}}\}_{j=1}^{\infty}$ which converges uniformly to a limit map, say $g$.
 Clearly, $g$ is also Lipschitz with $\operatorname{Lip}(g)\leq L$.
@@
 $\widetilde{g}_{n_{j}}\Big{(}\bigcup_{t=1}^{p}(c_{n_{j},t}O_{n_{j},t}K+b_{t})\Big{)}\subset\mathcal{N}_{\varepsilon}(g(E))\cap F\quad\text{ for all large $j$},$
-where $\mathcal{N}_{\varepsilon}(\cdot)$ denotes the open $\varepsilon$-neighborhood. Since $\mathcal{H}^{s}(F)<\infty$, we see that
-(2.4) $\mathcal{H}^{s}(g(E))=\lim_{\varepsilon\to 0}\mathcal{H}^{s}(\mathcal{N}_{\varepsilon}(g(E))\cap F)$
-and thus
+where $\mathcal{N}_{\varepsilon}(\cdot)$ denotes the open $\varepsilon$-neighborhood. Note that the restriction of $\mathcal{H}^{s}$ on $F$ is a Radon measure and, since $g(E)\subset F$,
+$\mathcal{H}^{s}(g(E))=\mathcal{H}^{s}(g(E)\cap F)=\lim_{\varepsilon\to 0}\mathcal{H}^{s}(\mathcal{N}_{\varepsilon}(g(E))\cap F).$
+Therefore,
 $\mathcal{H}^{s}(g(E))$ $\geq\limsup_{j\to\infty}\mathcal{H}^{s}\Big{(}\widetilde{g}_{n_{j}}\Big{(}\bigcup_{t=1}^{p}(c_{{n_{j}},t}O_{{n_{j}},t}K+b_{t})\Big{)}\Big{)}$
 $=\limsup_{j\to\infty}\mathcal{H}^{s}\Big{(}g_{n_{j}}\Big{(}\bigcup_{t=1}^{p}(c_{{n_{j}},t}O_{{n_{j}},t}K+b_{t})\Big{)}\Big{)}$
-$=\limsup_{j\to\infty}\mathcal{H}^{s}(f_{n_{j}}(K_{n_{j}})).$ (by (2.2))
+[…]
 This completes the proof. ∎
@@
 If $I=m^{k}$ for some integers $m,k>1$, then $K$ is Lipschitz equivalent to any self-similar set $E=\bigcup_{t=1}^{m}g_{t}(E)$ with the SSC such that each $g_{t}$ has similarity ratio $r^{1/k}$ (e.g., see [14]).
-So we may assume that $k=1$.
+So we may assume that $k=1$ and $m$ is not of the form $p^{\ell}$ where $p,\ell$ are integers strictly larger than $1$.
 Then $r^{s}=m^{-1}$ and thus
 $1=\sum_{j\in J}\lambda_{j}^{s}=\sum_{j\in J}r^{\alpha_{j}s}=\sum_{j\in J}m^{-\alpha_{j}}.$
-It follows from [2, Lemma 7.1] that $\alpha_{j}$ are all integers. Then by [2, Theorem 7.7], $K$ and $F$ are Lipschitz equivalent. ∎
+It follows from [2, Lemma 7.1] that $\alpha_{j}$ are all integers. Then by [2, Theorem 7.7] or [14, Theorem 1.5], $K$ and $F$ are Lipschitz equivalent. ∎
 3. Proof of Theorems 1.3
 In this section, we present the proof of Theorem 1.3. Regarding $\Phi=\{\varphi_{i}\}_{i\in I}$ and $\Psi=\{\psi_{j}\}_{j\in J}$, we adopt the following customary notations in the rest of this paper.
-  • $I^{*}:=\bigcup_{n=1}^{\infty}I^{n}$ and $K_{\bm{i}}:=\varphi_{\bm{i}}(K)$ for $\bm{i}\in I^{*}$;
-  • $J^{*}:=\bigcup_{n=1}^{\infty}J^{n}$ and $F_{\bm{j}}:=\psi_{\bm{j}}(F)$ for $\bm{j}\in J^{*}$;
-  • For $\bm{i}=i_{1}\cdots i_{n}\in I^{*}$, $\varphi_{\bm{i}}:=\varphi_{i_{1}}\circ\cdots\circ\varphi_{i_{n}}$ and $r_{\bm{i}}=r_{i_{1}}\cdots r_{i_{n}}$;
-  • For $\bm{j}=j_{1}\cdots j_{n}\in J^{*}$, $\psi_{\bm{j}}:=\psi_{j_{1}}\circ\cdots\circ\psi_{j_{n}}$ and $\lambda_{\bm{j}}:=\lambda_{j_{1}}\cdots\lambda_{j_{n}}$;
-  • $\underline{r}:=\min_{i}r_{i}$ and $\underline{\lambda}:=\min_{j}\lambda_{j}$;
-  • $\wedge$: the operation between any two words in $I^{*}$ or $J^{*}$ which produces their longest common prefix;
-  • For $\bm{i}\in I^{*}$ (resp. $\bm{j}\in J^{*}$), write $\bm{i}^{-}$ (resp. $\bm{j}^{-}$) to be the word obtained by removing the last digit of $\bm{i}$ (resp. $\bm{j}$);
-  • $|\cdot|$: either the modulus of a vector in $\mathbb{R}^{d}$ or the length of a finite word in $I^{*}\cup J^{*}$;
-  • $\Delta_{K}:=\min_{i\neq i^{\prime}}\operatorname{dist}(K_{i},K_{i^{\prime}})$ and $\Delta_{F}:=\min_{j\neq j^{\prime}}\operatorname{dist}(F_{j},F_{j^{\prime}})$, which are both positive numbers due to the SSC.
+Let $I^{*}:=\bigcup_{n=1}^{\infty}I^{n}$ be the set of all finite words over the alphabet $I$ with the convention that $I^{0}=\{\vartheta\}$, where $\vartheta$ denotes the empty word.
+For $n\geq 0$ and $\bm{i}\in I^{n}$, we call $n$ the length of $\bm{i}$ and denote it by $|\bm{i}|$.
+When $n\geq 1$, we write for $\bm{i}=i_{1}\cdots i_{n}\in I^{n}$ that
+$\varphi_{\bm{i}}:=\varphi_{i_{1}}\circ\cdots\circ\varphi_{i_{n}},\quad K_{\bm{i}}:=\varphi_{\bm{i}}(K),\quad r_{\bm{i}}:=r_{i_{1}}\cdots r_{i_{n}},\quad\bm{i}^{-}:=i_{1}\cdots i_{n-1},$
+with the convention that $\bm{i}^{-}=\vartheta$ if $n=1$; when $n=0$, we write $\varphi_{\vartheta}$ to be the identity map and let $K_{\vartheta}:=K$ and $r_{\vartheta}:=1$.
+For any $\bm{i},\bm{i}^{\prime}\in I^{*}$, $\bm{i}\wedge\bm{i}^{\prime}$ denotes the longest common prefix of $\bm{i}$ and $\bm{i}^{\prime}$.
+Analogously, for the alphabet $J$, we write $J^{*}:=\bigcup_{n=1}^{\infty}J^{n}$, $J^{0}:=\{\vartheta\}$ and adopt the notation $\psi_{\bm{j}}$, $F_{\bm{j}}$, $\lambda_{\bm{j}}$, $\bm{j}^{-}$ and $\bm{j}\wedge\bm{j}^{\prime}$ in the same manner as above.
+Furthermore, let $\underline{r}:=\min_{i\in I}r_{i}$, $\underline{\lambda}:=\min_{j\in J}\lambda_{j}$, $\Delta_{K}:=\min_{i\neq i^{\prime}\in I}\operatorname{dist}(K_{i},K_{i^{\prime}})$ and $\Delta_{F}:=\min_{j\neq j^{\prime}\in J}\operatorname{dist}(F_{j},F_{j^{\prime}})$.
+The SSC ensures that $\Delta_{K}$ and $\Delta_{F}$ are both positive.
 Our first observation is a reduction of the problem to its embedding case.
@@
 Proof.
-We begin with constructing a map $g:K\to F$ with $g(K)=f(K)\cap F$.
-For any $\bm{i}\in I^{*}$ such that $f(K_{\bm{i}})\cap F=\varnothing$ but $f(K_{\bm{i}^{-}})\cap F\neq\varnothing$, we choose an arbitrary point $y_{\bm{i}}$ in $f(K_{\bm{i}^{-}})\cap F$ and define $g(K_{\bm{i}})=\{y_{\bm{i}}\}$.
-So if $x\in K$ is such that $f(x)\notin F$, then $g(x)$ is defined as $y_{\omega}$, where $\omega$ is the unique word in $I^{*}$ such that $x\in K_{\omega}$, $f(K_{\omega})\cap F=\varnothing$ and $f(K_{\omega^{-}})\cap F\neq\varnothing$.
-Note that the existence of $\omega$ is due to the compactness of $F$.
-If $f(x)\in F$ we simply set $g(x)=f(x)$.
+We begin by constructing a map $g:K\to F$ with $g(K)=f(K)\cap F$.
+For any $\bm{i}\in I^{*}$ such that $f(K_{\bm{i}})\cap F=\varnothing$ but $f(K_{\bm{i}^{-}})\cap F\neq\varnothing$, choose an arbitrary point $y_{\bm{i}}$ in $f(K_{\bm{i}^{-}})\cap F$ and define $g(K_{\bm{i}})=\{y_{\bm{i}}\}$.
+So for those $x\in K$ with $f(x)\notin F$, $g(x)$ is defined as $y_{\omega}$, where $\omega\in I^{*}$ is the unique word such that $x\in K_{\omega}$, $f(K_{\omega})\cap F=\varnothing$ and $f(K_{\omega^{-}})\cap F\neq\varnothing$.
+Note that the existence of such an $\omega$ is guaranteed by the closedness of $F$.
+For $x\in K$ with $f(x)\in F$, we simply set $g(x)=f(x)$.
 In this way, the map $g:K\to F$ is well defined and it is clear that $g(K)=f(K)\cap F$.
 We claim that $\operatorname{diam}(g(K_{\bm{i}}))\leq\operatorname{diam}(f(K_{\bm{i}}))$ for all $\bm{i}\in I^{*}$.
-In fact, if $f(K_{\bm{i}})\cap F=\varnothing$, then $g(K_{\bm{i}})$ is a singleton and the inequality clearly holds.
+Indeed, if $f(K_{\bm{i}})\cap F=\varnothing$, then $g(K_{\bm{i}})$ is a singleton and the inequality holds trivially.
 If $f(K_{\bm{i}})\cap F\neq\varnothing$, then by the definition of $g$,
@@
 which also implies the desired inequality.
-It remains to prove that $g$ is Lipschitz. For any distinct $x,x^{\prime}\in K$, let $\omega\in I^{*}$ be such that $|\omega|=\max\{|\bm{i}|:\bm{i}\in I^{*},x,x^{\prime}\in K_{\bm{i}}\}$. So
+It remains to prove that $g$ is Lipschitz. For any distinct $x,x^{\prime}\in K$, let $\omega\in I^{*}\cup I^{0}$ be the longest word such that $x,x^{\prime}\in K_{\omega}$. So
 $|x-x^{\prime}|\geq\min_{i\neq i^{\prime}}\operatorname{dist}(K_{\omega i},K_{\omega i^{\prime}})=r_{\omega}\Delta_{K}=\Delta_{K}\operatorname{diam}(K_{\omega}).$
@@
 So $g$ is Lipschitz. ∎
-Based on this, the proof of Theorem 1.3 reduces to establishing the following result.
+The proof of Theorem 1.3 then reduces to establishing the next result.
 Theorem 3.2.
@@
 (3.1) $\left\{\begin{array}[]{l}I_{n}:=\{\bm{i}\in I^{*}:r_{\bm{i}}\leq\delta^{n}<r_{\bm{i}^{-}}\},\\ J_{n}:=\{\bm{j}\in J^{*}:\lambda_{\bm{j}}\leq\delta_{n}<\lambda_{\bm{j}^{-}}\},\end{array}\right$.
-where $\delta_{n}:=\frac{L}{\Delta_{F}}\cdot\delta^{n}$.
-The following lemma records several basic facts that will be used frequently later.
+where $\delta_{n}:=\frac{L}{\Delta_{F}}\cdot\delta^{n}$ is a constant multiple of $\delta^{n}$.
+The next lemma records several basic facts that will be used frequently later.
 Recall that $\underline{r}:=\min_{i\in I}r_{i}$ and $\underline{\lambda}:=\min_{j\in J}\lambda_{j}$.
@@
 Proof.
-The first statement follows directly from (3.1). For (2), note that
+The first statement follows directly from (3.1) and the fact that $\sum_{\bm{i}\in I_{n}}r_{\bm{i}}^{s}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}=1$. For (2), note that
 $\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})\geq r_{\bm{i}\wedge\bm{i}^{\prime}}\Delta_{K}\geq r_{\bm{i}^{-}}\Delta_{K}>\Delta_{K}\delta^{n};$
 similarly,
-$\operatorname{dist}(F_{\bm{j}},F_{\bm{j}^{\prime}})\geq\lambda_{\bm{j}^{-}}\Delta_{F}>\delta_{n}\Delta_{F}=\frac{L}{\Delta_{F}}\delta^{n}\Delta_{F}=L\delta^{n}.$
+$\operatorname{dist}(F_{\bm{j}},F_{\bm{j}^{\prime}})\geq\lambda_{\bm{j}^{-}}\Delta_{F}>\delta_{n}\Delta_{F}=L\delta^{n}.$
 For (3), just note that when $\bm{i}\in I_{n}$,
@@
 Since $\bm{j}_{n}(y)\in J_{n}$, $\operatorname{diam}(F_{\bm{j}_{n}(y)})=\lambda_{\bm{j}_{n}(y)}\leq\delta_{n}$ and thus $F_{\bm{j}_{n}(y)}\subset B(y,\rho)$.
-On the other hand, by Lemma 3.3(1) and the minimality of $n$ (so $\delta_{n}\approx\rho$), there is a constant $C>0$ (independent of $y,n,\varepsilon$) such that
+On the other hand, by Lemma 3.3(1) and the minimality of $n$, we have $\rho/C_{1}\leq\delta_{n}\leq C_{1}\rho$ for some constant $C_{1}>1$.
+By Lemma 3.3(2), one can find a constant $C_{2}>0$ such that for all $n\geq 1$ and $\bm{j}\in J_{n}$, there exists $x_{\bm{j}}\in F_{\bm{j}}$ such that $\{B(x_{\bm{j}},C_{2}\rho)\}_{\bm{j}\in J_{n}}$ are mutually disjoint.
+Thus
+$\#\big{\{}\bm{j}\in J_{n}:F_{\bm{j}}\cap B(y,\rho)\not$$\emptyset\big{\}}$ $\leq\#\big{\{}\bm{j}\in J_{n}:B(x_{\bm{j}},C_{2}\rho)\subset B\big{(}y,(1+C_{1}+C_{2})\rho\big{)}\big{\}}$
+$\leq\Big{(}\frac{1+C_{1}+C_{2}}{C_{2}}\Big{)}^{d}.$
+As a result, there is a constant $C>0$ (independent of $y,n,\varepsilon$) such that
 (3.4) $\mathcal{H}^{s}(F\cap B(y,\rho))\leq\sum_{\bm{j}\in J_{n}:F_{\bm{j}}\cap B(y,\rho)\neq\varnothing}\mathcal{H}^{s}(F_{\bm{j}})\leq C\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}).$
-Combining all these observations and noticing that
-$\big{(}F\cap B(y,\rho)\big{)}\setminus f(K)\supset\big{(}F_{\mathbf{j}_{n}(y)}\cap B(y,\rho)\big{)}\setminus f(K)=F_{\mathbf{j}_{n}(y)}\setminus f(K),$
-we have
-$\mathcal{H}^{s}(f(K)\cap B(y,\rho))$ $\leq\mathcal{H}^{s}(F\cap B(y,\rho))-\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))$
+Recalling that $f(K)\subset F$ and $F_{\bm{j}_{n}(y)}\subset B(y,\rho)\cap F$, we have
+$\mathcal{H}^{s}(f(K)\cap B(y,\rho))$ $\leq\mathcal{H}^{s}(F\cap B(y,\rho))-\mathcal{H}^{s}((F\cap B(y,\rho))\setminus f(K))$
+$\leq\mathcal{H}^{s}(F\cap B(y,\rho))-\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))$
 $=\mathcal{H}^{s}(F\cap B(y,\rho))-\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))}{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})}\cdot\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})$
 $\leq\Big{(}1-C^{-1}\cdot\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))}{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})}\Big{)}\mathcal{H}^{s}(F\cap B(y,\rho)),$
-where the last inequality follows from (3.4). Recalling (3.3), we immediately obtain
+where the last inequality follows from (3.4). By (3.3), we immediately obtain
 $\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)}\setminus f(K))}{\mathcal{H}^{s}(F_{\bm{j}_{n}(y)})}\leq C\varepsilon.$
 Since $\varepsilon$ is arbitrary, the proof is finished. ∎
-To obtain an effective rescaling, we also need to identify a non-condensed local area of the image set $f(K)$.
+To obtain an effective rescaling, we also need to determine a local region within $F$ where the map $f$ is not overly concentrated.
 Proposition 3.5.
 There is an integer $Q\geq 1$ satisfying the following property.
-For any $0<\varepsilon<1$, we can find some $n$ and $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(F_{\bm{j}}\cap f(K))}{\mathcal{H}^{s}(F_{\bm{j}})}>1-\varepsilon$ and
+For any $0<\varepsilon<1$, we can find some $n$ and $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(f(K)\cap F_{\bm{j}})}{\mathcal{H}^{s}(F_{\bm{j}})}>1-\varepsilon$ and
 $\{\bm{i}\in I_{n}:f(K_{\bm{i}})\subset F_{\bm{j}}\}=\#\{\bm{i}\in I_{n}:f(K_{\bm{i}})\cap F_{\bm{j}}\neq\varnothing\}\leq Q.$
 Proof.
-By Lemma 3.3(3), the first equality holds trivially since the two sets are identical.
+The first equality is immediately from Lemma 3.3(3) as the two sets are identical.
 It remains to verify the cardinality.
@@
 (3.5) $\mathcal{H}^{s}(f(K)\cap F_{\bm{j}_{n_{t}}(y_{t})})>(1-\varepsilon)\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}(y_{t})}),\quad\forall t\geq 1.$
-For convenience, below we simply write $\bm{j}_{n_{t}}:=\bm{j}_{n_{t}}(y_{t})$.
-Pick any $p\geq 1$ such that
-(3.6) $\mathcal{H}^{s}\Big{(}\bigcup_{t=1}^{p}F_{\bm{j}_{n_{t}}}\Big{)}\geq\frac{1}{2}\mathcal{H}^{s}(E)=\frac{1}{2}\mathcal{H}^{s}(f(K)).$
+For convenience, below we abbreviate $\bm{j}_{n_{t}}:=\bm{j}_{n_{t}}(y_{t})$.
+Pick any $p\geq 1$ such that $\mathcal{H}^{s}(\bigcup_{t=1}^{p}F_{\bm{j}_{n_{t}}})\geq\mathcal{H}^{s}(E)/2=\mathcal{H}^{s}(f(K))/2$. So
+(3.6) $\sum_{t=1}^{p}\delta_{n_{t}}^{s}\geq\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\lambda_{n_{t}}^{s}\mathcal{H}^{s}(F)=\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})\geq\frac{\mathcal{H}^{s}(f(K))}{2\mathcal{H}^{s}(F)}.$
 Pick an integer $P\geq 2\max\{n_{1},\ldots,n_{p}\}$ and a constant $C>\underline{\lambda}^{-2s}$. Set
@@
 and $\mathcal{M}:=\bigcup_{t=1}^{p}\mathcal{M}_{t}$.
-Since $F_{\bm{j}_{n_{1}}},\ldots,F_{\bm{j}_{n_{p}}}$ are distinct, $\mathcal{M}_{1},\ldots,\mathcal{M}_{p}$ are disjoint.
-We claim that $\mathcal{M}\gtrsim\delta_{N}^{-s}$.
+Since $F_{\bm{j}_{n_{1}}},\ldots,F_{\bm{j}_{n_{p}}}$ are disjoint, $\mathcal{M}_{1},\ldots,\mathcal{M}_{p}$ are disjoint.
+We claim that $\mathcal{M}\gtrsim\delta_{P}^{-s}$.
 To this end, fix any $1\leq t\leq p$. From $\bm{j}_{n_{t}}\in J_{n_{t}}$ and (3.5),
@@
 $\#\mathcal{M}_{t}$ $=\#\{\eta\in J_{P}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}-\#\{\eta\in J_{P}\setminus\mathcal{M}_{t}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}$
-$\geq\#\{\eta\in J_{P}:F_{\eta}\subset F_{\bm{j}_{n_{t}}}\}-C^{-1}\underline{\lambda}^{-s}\delta_{n_{t}}^{s}\delta_{P}^{-s}$
 $\geq\frac{\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})}{\max_{\eta\in J_{P}}\mathcal{H}^{s}(F_{\eta})}-C^{-1}\underline{\lambda}^{-s}\delta_{n_{t}}^{s}\delta_{P}^{-s}$
 $\geq\frac{(\underline{\lambda}\delta_{n_{t}})^{s}}{\delta_{P}^{s}}-C^{-1}\underline{\lambda}^{-s}\delta_{n_{t}}^{s}\delta_{P}^{-s}$
-(3.9) $=(\underline{\lambda}^{s}-C^{-1}\underline{\lambda}^{-s})\delta_{n_{t}}^{s}\delta_{P}^{-s}=:C_{1}\delta_{n_{t}}^{s}\delta_{P}^{-s}.$
-It follows from our choice of $C$ that $C_{1}>0$. Since the above estimate holds for all $1\leq t\leq p$,
-$\#\mathcal{M}$ $=\sum_{t=1}^{p}\#\mathcal{M}_{t}$ (by the disjointness)
-$\geq C_{1}\delta_{N}^{-s}\sum_{t=1}^{p}\delta_{n_{t}}^{s}$ (by (3.9))
-$=C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\delta_{n_{t}}^{s}\mathcal{H}^{s}(F)$
-$\geq C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\sum_{t=1}^{p}\mathcal{H}^{s}(F_{\bm{j}_{n_{t}}})$ (since $\lambda_{\bm{j}_{n_{t}}}\leq\delta_{n_{t}}$)
-$\geq C_{1}\delta_{P}^{-s}\cdot\mathcal{H}^{s}(F)^{-1}\cdot\frac{1}{2}\mathcal{H}^{s}(f(K))$ (by (3.6))
-$=:C_{2}\delta_{P}^{-s},$
+(3.9) $=(\underline{\lambda}^{s}-C^{-1}\underline{\lambda}^{-s})\delta_{n_{t}}^{s}\delta_{P}^{-s}=:C_{1}\delta_{n_{t}}^{s}\delta_{P}^{-s},$
+where $C_{1}>0$ since $C>\underline{\lambda}^{-2s}$.
+Letting $C_{2}:=\frac{C_{1}\mathcal{H}^{s}(f(K))}{2\mathcal{H}^{s}(F)}$, we know from the disjointness of $\{\mathcal{M}_{t}\}_{t=1}^{p}$, the estimates (3.9) and (3.6) that
+$\#\mathcal{M}=\sum_{t=1}^{p}\#\mathcal{M}_{t}\geq C_{1}\delta_{P}^{-s}\sum_{t=1}^{p}\delta_{n_{t}}^{s}\geq C_{2}\delta_{P}^{-s},$
 as claimed.
-Now, by Lemma 3.3(3), $\{\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\}_{\eta\in\mathcal{M}}$ is a disjoint family.
-By the pigeonhole principle, there must exist a word $\eta\in\mathcal{M}$ such that
-$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\leq\frac{\#I_{P}}{\#\mathcal{M}}.$
+Lemma 3.3(3) also implies that $\{\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\}_{\eta\in\mathcal{M}}$ is a disjoint family.
+By the pigeonhole principle, there must exist a word $\eta_{0}\in\mathcal{M}$ such that
+$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta_{0}}\neq\varnothing\}\leq\frac{\#I_{P}}{\#\mathcal{M}}.$
 By Lemma 3.3(1), $I_{P}\approx\delta^{-Ps}$. Hence
-$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta}\neq\varnothing\}\lesssim\frac{\delta^{-Ps}}{C_{2}\delta_{P}^{-s}}=\frac{\delta^{-Ps}}{C_{2}}\cdot\Big{(}\frac{L\delta^{P}}{\Delta_{F}}\Big{)}^{s}=\frac{L^{s}}{C_{2}\Delta_{F}^{s}},$
+$\{\bm{i}\in I_{P}:f(K_{\bm{i}})\cap F_{\eta_{0}}\neq\varnothing\}\lesssim\frac{\delta^{-Ps}}{C_{2}\delta_{P}^{-s}}=\frac{\delta^{-Ps}}{C_{2}}\cdot\Big{(}\frac{L\delta^{P}}{\Delta_{F}}\Big{)}^{s}=\frac{L^{s}}{C_{2}\Delta_{F}^{s}},$
 Combining this with (3.7), we prove the proposition. ∎
 3.2. Proof of Theorem 3.2
-Now let us prove Theorem 3.2.
-Applying Proposition 3.5 to $\varepsilon=n^{-1}$, we find $\{k_{n}\}_{n=1}^{\infty}\subset\mathbb{Z}^{+}$ and $\{\bm{j}_{n}\}_{n=1}^{\infty}$ (say $\bm{j}_{n}\in J_{k_{n}}$) such that
-(3.10) $\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap f(K))>(1-n^{-1})\mathcal{H}^{s}(F_{\bm{j}_{n}})$
+Now let us prove Theorem 3.2. Applying Proposition 3.5 to $\varepsilon=n^{-1}$, we find for all $n\geq 1$ an integer $k_{n}\geq 1$ and a word $\bm{j}_{n}\in J_{k_{n}}$ such that
+(3.10) $\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap f(K))>(1-n^{-1})\mathcal{H}^{s}(F_{\bm{j}_{n}}),$
 whereas
@@
 (3.11) $\mathcal{H}^{s}\Big{(}F_{\bm{j}_{n}}\cap\bigcup_{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}}f(K_{\bm{i}})\Big{)}>(1-n^{-1})\mathcal{H}^{s}(F_{\bm{j}_{n}}).$
-For convenience, enumerate $\{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}\}=:\{\bm{i}_{n,q}\}_{q=1}^{P}$.
-With the notation, for every $n\geq 1$, $f$ sends each $K_{\bm{i}_{n,q}}$ into $F_{\bm{j}_{n}}=\psi_{\bm{j}_{n}}(F)$. This naturally generates a blow-up map
-$f_{n}:\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})$ $\to F$
+For convenience, enumerate $\{\bm{i}\in I_{k_{n}}:f(K_{\bm{i}})\subset F_{\bm{j}_{n}}\}=:\{\bm{i}_{n,t}\}_{t=1}^{P}$.
+With the notation, for every $n\geq 1$, $f$ sends each $K_{\bm{i}_{n,t}}$ into $F_{\bm{j}_{n}}=\psi_{\bm{j}_{n}}(F)$. This naturally generates a blow-up map
+$f_{n}:\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}})$ $\to F$
 $\psi_{\bm{j}_{n}}^{-1}(x)$ $\mapsto\psi_{\bm{j}_{n}}^{-1}f(x).$
-It follows that
-$\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})\Big{)}\Big{)}$ $=\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,q}})\Big{)}$
-$=\mathcal{H}^{s}(F)\cdot\lim_{n\to\infty}\frac{\mathcal{H}^{s}(F\cap\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,q}}))}{\mathcal{H}^{s}(F)}$
-$=\mathcal{H}^{s}(F)\cdot\lim_{n\to\infty}\frac{\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\bigcup_{q=1}^{P}f(K_{\bm{i}_{n,q}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n}})}$
-(3.12) $=\mathcal{H}^{s}(F),$
-where the last equality follows from (3.11).
-For $1\leq q\leq P$, write $\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})=:c_{n,q}O_{n,q}K+t_{n,q}$, where $c_{n,q}:=\lambda_{\bm{j}_{n}}^{-1}r_{\bm{i}_{n,q}}$.
-Recall that $\bm{j}_{n}\in J_{k_{n}}$ and $\bm{i}_{n,q}\in I_{k_{n}}$.
-By Lemma 3.3, it is easy to compute that $c_{n,q}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$ and
-$\min_{q\neq q^{\prime}}\operatorname{dist}(\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}}),\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q^{\prime}}}))$ $=\lambda_{\bm{j}_{n}}^{-1}\min_{q\neq q^{\prime}}\operatorname{dist}(K_{\bm{i}_{n,q}},K_{\bm{i}_{n,q^{\prime}}})$
+Clearly, $\operatorname{Lip}(f_{n})\leq\operatorname{Lip}(f)=L$ for all $n$. By (3.11),
+$\lim_{n\to\infty}\frac{\mathcal{H}^{s}\big{(}F\cap f_{n}(\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}}))\big{)}}{\mathcal{H}^{s}(F)}$ $=\lim_{n\to\infty}\frac{\mathcal{H}^{s}\big{(}F\cap\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}f(K_{\bm{i}_{n,t}})\big{)}}{\mathcal{H}^{s}(F)}$
+(3.12) $=\lim_{n\to\infty}\frac{\mathcal{H}^{s}\big{(}F_{\bm{j}_{n}}\cap\bigcup_{t=1}^{P}f(K_{\bm{i}_{n,t}})\big{)}}{\mathcal{H}^{s}(F_{\bm{j}_{n}})}=1.$
+For $1\leq t\leq P$, write $\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}})=:c_{n,t}O_{n,t}K+b_{n,t}$, where $c_{n,t}:=\lambda_{\bm{j}_{n}}^{-1}r_{\bm{i}_{n,t}}$.
+Recall that $\bm{j}_{n}\in J_{k_{n}}$ and $\bm{i}_{n,t}\in I_{k_{n}}$.
+By Lemma 3.3, it is easy to verify that $c_{n,t}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$ and
+$\min_{t\neq t^{\prime}}\operatorname{dist}(\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}}),\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t^{\prime}}}))$ $=\lambda_{\bm{j}_{n}}^{-1}\min_{t\neq t^{\prime}}\operatorname{dist}(K_{\bm{i}_{n,t}},K_{\bm{i}_{n,t^{\prime}}})$
 $>\lambda_{\bm{j}_{n}}^{-1}\Delta_{K}\delta^{k_{n}}$
 $\geq L^{-1}\Delta_{K}\Delta_{F}.$
-Applying Lemma 2.2, we can find a Lipschitz map $g:\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q})\to F$, where $c_{q}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$, $O_{q}$ is a $d\times d$ orthogonal matrix and $a_{q}\in\mathbb{R}^{d}$, such that $\operatorname{Lip}(g)\leq L$ and (recalling (3.12))
-$\mathcal{H}^{s}\Big{(}g\Big{(}\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q})\Big{)}\Big{)}\geq\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{q=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,q}})\Big{)}\Big{)}=\mathcal{H}^{s}(F).$
-So $g(\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q}))$ is a closed subset of $F$ with full measure. As a result, $F=g(\bigcup_{q=1}^{P}(c_{q}O_{q}K+a_{q}))$.
-By Baire’s theorem, there is some $q$ such that $g(cO_{q}K+a_{q})$ contains an interior part of $F$, say $F_{\bm{j}}\subset g(cO_{q}K+a_{q})$.
-Thus, similarly as in the proof of Lemma 3.1, it is easy to construct a Lipschitz map sending $cO_{q}K+a_{q}$ onto $F_{\bm{j}}$.
-Since $cO_{q}K+a_{q}$ and $F_{\bm{j}}$ are nothing but scaled copies of $K$ and $F$, respectively, this establishes Theorem 3.2.
+Applying Lemma 2.2, we can find a Lipschitz map $g:\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t})\to F$, where $c_{t}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$, $O_{t}$ is a $d\times d$ orthogonal matrix and $b_{t}\in\mathbb{R}^{d}$, such that $\operatorname{Lip}(g)\leq L$ and (recalling (3.12))
+$\mathcal{H}^{s}\Big{(}g\Big{(}\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t})\Big{)}\Big{)}\geq\lim_{n\to\infty}\mathcal{H}^{s}\Big{(}F\cap f_{n}\Big{(}\bigcup_{t=1}^{P}\psi_{\bm{j}_{n}}^{-1}(K_{\bm{i}_{n,t}})\Big{)}\Big{)}=\mathcal{H}^{s}(F).$
+Therefore, the image $g(\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t}))$ is a closed subset of $F$ with full measure. It follows that $g(\bigcup_{t=1}^{P}(c_{t}O_{t}K+b_{t}))=F$.
+By Baire’s theorem, there is some $t$ such that $g(c_{t}O_{t}K+b_{t})$ contains an interior part of $F$, say $F_{\bm{j}}\subset g(c_{t}O_{t}K+b_{t})$.
+Then, similarly as in the proof of Lemma 3.1, it is easy to construct a Lipschitz map sending $c_{t}O_{t}K+b_{t}$ onto $F_{\bm{j}}$.
+Since $c_{t}O_{t}K+b_{t}$ and $F_{\bm{j}}$ are nothing but scaled copies of $K$ and $F$, respectively, this establishes Theorem 3.2.
 4. Lipschitz surjections with measure-preserving properties
-In this section, we fix a Lipschitz embedding $f:K\to F$.
+In this section, we fix a Lipschitz embedding $f:K\to F$ with $\mathcal{H}^{s}(f(K))>0$.
 By Theorem 3.2, we may assume $f$ to be surjective.
 Again write $L:=\operatorname{Lip}(f)$ for simplicity.
-A key insight, noted in [3, 21], asserts that any bi-Lipschitz map between $K$ and $F$ (if there is any) should preserve the $s$-dimensional Hausdorff measure in some local area.
-See [14, Lemma 2.4] for a short proof.
+A key insight, noted in [3, 21], asserts that any bi-Lipschitz map between $K$ and $F$ (if there is any) should preserve the $s$-dimensional Hausdorff measure in some local area; see [14, Lemma 2.4] for a short proof.
 While the given map $f$ may fall short of this ideal, we can nevertheless produce an alternative Lipschitz surjection from $K$ to $F$ that enjoys some nice measure linearity.
@@
 with $\operatorname{Lip}(g)\leq\operatorname{Lip}(f)=L$.
-It remains to check that $\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}(K_{\bm{i}})\in\mathcal{K}$. By Lemma 3.3(1), we have
+It remains to check that $\bigcup_{\bm{i}\in\mathcal{I}}\psi_{\bm{j}}^{-1}(K_{\bm{i}})\in\mathcal{K}$. By Lemma 3.3, we have
 $\frac{\underline{r}\Delta_{F}}{L}\leq\lambda_{\bm{j}}^{-1}r_{\bm{i}}\leq\frac{\Delta_{F}}{\underline{\lambda}L},\quad\forall\bm{i}\in\mathcal{I}$
@@
 The above lemma tells us that the infimum is well defined.
-Then one can take $\{E_{n}\}\subset\mathcal{K}$ and a sequence $\{g_{n}\}$, where each $g_{n}:E_{n}\to F$ is a Lipschitz surjection with $\operatorname{Lip}(g_{n})\leq L$, such that $\lim_{n}\mathcal{H}^{s}(E_{n})$ equals the infimum.
+Then one can take $\{E_{n}\}_{n=1}^{\infty}\subset\mathcal{K}$ and a sequence $\{g_{n}\}_{n=1}^{\infty}$, where each $g_{n}:E_{n}\to F$ is a Lipschitz surjection with $\operatorname{Lip}(g_{n})\leq L$, such that $\lim_{n}\mathcal{H}^{s}(E_{n})$ equals the infimum.
 By Lemma 2.2, we can find $\widetilde{K}\in\mathcal{K}$ and a Lipschitz map $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$ such that
@@
 […]
-In the sequel, let us fix such an optimal domain $\widetilde{K}\in\mathcal{K}$, say $\widetilde{K}=\bigcup_{t=1}^{p}(a_{t}O_{t}K+b_{t})$, and a Lipschitz surjection $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$.
+In the sequel of the paper, let us fix such an optimal $\widetilde{K}\in\mathcal{K}$, say $\widetilde{K}=\bigcup_{t=1}^{p}(a_{t}O_{t}K+b_{t})$, and a Lipschitz surjection $\widetilde{f}:\widetilde{K}\to F$ with $\operatorname{Lip}(\widetilde{f})\leq L$.
 Write $\widetilde{c}:=\frac{\mathcal{H}^{s}(\widetilde{K})}{\mathcal{H}^{s}(F)}$.
+Also fix a small constant $0<\delta<\min\{1,a_{1},a_{2},\ldots,a_{p}\}$.
 […]
-(4.3) $I^{\prime}_{t,n}:=\{\bm{i}\in I^{*}:a_{t}r_{\bm{i}}\leq\delta^{n}<a_{t}r_{\bm{i}^{-}}\},\quad 1\leq t\leq p.$
+(4.3) $I^{\prime}_{t,n}:=\{\bm{i}\in I^{*}:a_{t}r_{\bm{i}}\leq\delta^{n}<a_{t}r_{\bm{i}^{-}}\},\quad 1\leq t\leq p$
 instead of $I_{n}$ (defined in (3.1)).
 Lemma 4.3.
-For every $n$ and $\bm{j}\in J_{n}$, $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for some $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$.
+For every $n$ and $\bm{j}\in J_{n}$, $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ with $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$ for all $1\leq t\leq p$.
 Moreover, if $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$, then $\psi_{\bm{j}}^{-1}(\widetilde{f}^{-1}(F_{\bm{j}}))\in\mathcal{K}$ and $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$.
@@
 $\operatorname{diam}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$ $\leq L\operatorname{diam}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
-$=La_{t}r_{\bm{i}}\leq L\delta^{n}<\min_{\eta\neq\eta^{\prime}\in J_{n}}\operatorname{dist}(F_{\eta},F_{\eta^{\prime}}),$
+(4.4) $=La_{t}r_{\bm{i}}\leq L\delta^{n}<\min_{\eta\neq\eta^{\prime}\in J_{n}}\operatorname{dist}(F_{\eta},F_{\eta^{\prime}}),$
 that is, $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ should be contained in a unique $F_{\eta}$. Therefore, one can write
-(4.4) $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
-for some $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$.
-Secondly, we show that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{m}\bigcup_{\bm{i}\in\mathcal{I}_{t}}\psi_{\bm{j}}^{-1}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ is a member of $\mathcal{K}$ provided that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$.
+(4.5) $\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$
+where $\mathcal{I}_{t}\subset I^{\prime}_{t,n}$ for all $1\leq t\leq p$.
+Secondly, we show that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}\psi_{\bm{j}}^{-1}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ is a member of $\mathcal{K}$ provided that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$.
 Let $1\leq t,t^{\prime}\leq p$, $\bm{i}\in\mathcal{I}_{t}$ and $\bm{i}^{\prime}\in\mathcal{I}_{t^{\prime}}$.
 By Lemma 3.3(1), every ratio $\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}}\in[\frac{\underline{r}\Delta_{F}}{L},\frac{\Delta_{F}}{\underline{\lambda}L}]$.
-On the other hand, if $t\neq t^{\prime}$ then since $\widetilde{K}\in\mathcal{K}$,
+Furthermore, if $t\neq t^{\prime}$ then since $\widetilde{K}\in\mathcal{K}$,
 $\operatorname{dist}(\psi_{\bm{j}}^{-1}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$ $\psi_{\bm{j}}^{-1}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))$
-$>\operatorname{dist}(a_{t}O_{t}K+b_{t},a_{t^{\prime}}O_{t^{\prime}}K+b_{t^{\prime}})\geq\gamma=L^{-1}\Delta_{F}\Delta_{K};$
+$>\operatorname{dist}(a_{t}O_{t}K_{\bm{i}}+b_{t},a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})$
+$\geq\operatorname{dist}(a_{t}O_{t}K+b_{t},a_{t^{\prime}}O_{t^{\prime}}K+b_{t^{\prime}})\geq\gamma;$
 while if $t=t^{\prime}$ and $\bm{i}\neq\bm{i}^{\prime}$ then
 $\operatorname{dist}(\psi_{\bm{j}}^{-1}(a_{t}O_{t}$ $K_{\bm{i}}+b_{t}),\psi_{\bm{j}}^{-1}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))$
-$=\lambda_{\bm{j}}^{-1}a_{t}\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})\geq\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}^{-}}\Delta_{K}\geq\frac{\Delta_{F}}{L\delta^{n}}\cdot\delta^{n}\cdot\Delta_{K}=L^{-1}\Delta_{F}\Delta_{K}$
+$=\lambda_{\bm{j}}^{-1}a_{t}\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})$
+$\geq\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}^{-}}\Delta_{K}\geq\frac{\Delta_{F}}{L\delta^{n}}\cdot\delta^{n}\cdot\Delta_{K}=L^{-1}\Delta_{F}\Delta_{K}=\gamma$
 as well.
-Assume that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$. Then
-$\Big{(}\sum_{t=1}^{p}\#\mathcal{I}_{t}\Big{)}\cdot(\underline{r}\delta^{n})^{s}\mathcal{H}^{s}(K)$ $\leq\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))$ (recall (4.4))
-$\leq\widetilde{c}\mathcal{H}^{s}(F_{\bm{j}})$
-$\leq\frac{\mathcal{H}^{s}(K)}{\mathcal{H}^{s}(F)}\cdot\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(F)$ (by Corollary 4.2)
-$=\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(K)$
-$\leq\frac{(L\delta^{n})^{s}}{\Delta_{F}^{s}}\mathcal{H}^{s}(K),$ (since $\bm{j}\in J_{n}$)
-which implies that $\sum_{t=1}^{p}\#\mathcal{I}_{t}\leq\lfloor\frac{L^{s}}{\underline{r}^{s}\Delta_{F}^{s}}\rfloor=P$. So $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
-Note that the inclusion $\widetilde{f}:\widetilde{f}^{-1}(F_{\bm{j}})\to F_{\bm{j}}$ naturally generates a Lipschitz surjection
+Now assume that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$.
+Then since $\mathcal{H}^{s}(\widetilde{K})\leq\mathcal{H}^{s}(K)$ (see Corollary 4.2) and $\bm{j}\in J_{n}$,
+$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))\leq\widetilde{c}\mathcal{H}^{s}(F_{\bm{j}})=\frac{\mathcal{H}^{s}(\widetilde{K})}{\mathcal{H}^{s}(F)}\cdot\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(F)\leq\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(K)\leq\frac{(L\delta^{n})^{s}}{\Delta_{F}^{s}}\mathcal{H}^{s}(K).$
+On the other hand, the expression (4.5) immediately implies
+$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))\geq\Big{(}\sum_{t=1}^{p}\#\mathcal{I}_{t}\Big{)}\cdot(\underline{r}\delta^{n})^{s}\mathcal{H}^{s}(K).$
+Thus $\sum_{t=1}^{p}\#\mathcal{I}_{t}\leq\lfloor\frac{L^{s}}{\underline{r}^{s}\Delta_{F}^{s}}\rfloor=P$. Altogether, we have $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
+Furthermore, the inclusion $\widetilde{f}:\widetilde{f}^{-1}(F_{\bm{j}})\to F_{\bm{j}}$ naturally generates a Lipschitz surjection
 $g:\psi_{\bm{j}}^{-1}(\widetilde{f}^{-1}(F_{\bm{j}}))$ $\to F$
@@
 which is a convex combination.
-This identity should be $\widetilde{c}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}\widetilde{c}$.
-Otherwise, since every term in that combination is positive, there exists $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}<\widetilde{c}$, which contradicts Lemma 4.3.
-In particular, $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$ for every $\bm{j}\in J_{n}$.
-This in turn implies by Lemma 4.3 that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
+The identity should be $\widetilde{c}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}\widetilde{c}$.
+Otherwise, since every term in that combination is positive, there exists $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}<\widetilde{c}$, contradicting Lemma 4.3.
+In particular, $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$ for every $\bm{j}\in J_{n}$, which in turn implies by Lemma 4.3 that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
 ∎
@@
 Proof.
-Write $r$ to be the common ratio of maps in $\Phi$.
+Write $r$ to be the common ratio of similitudes in $\Phi$.
 By Lemma 2.4, it suffices to show that $\frac{\log\lambda_{j}}{\log r}\in\mathbb{Q}$ for $j\in J$.
@@
 $j^{k_{n}}:=\underbrace{jj\cdots j}_{k_{n}\text{ terms}}\in J_{n}.$
-Since $r_{i}\equiv r$, there exists $q_{t,n}\in\mathbb{N}$ such that $I^{\prime}_{t,n}\subset I^{q_{t,n}}$, where $I^{\prime}_{t,n}$ was defined in (4.3).
-Moreover, for $t\neq t^{\prime}$, we always have $a_{t}r^{q_{t,n}}\leq\delta^{n}<a_{t^{\prime}}r^{q_{t^{\prime},n}-1}$ and hence $r^{q_{t^{\prime},n}-q_{t,n}}>a_{t}r/a_{t^{\prime}}$.
-So there is a constant $C\in\mathbb{N}$ such that
-(4.5) $|q_{t,n}-q_{t^{\prime},n}|\leq C,\quad\forall 1\leq t,t^{\prime}\leq m,\forall n\in\mathbb{Z}^{+}.$
-By (4.4) and Proposition 4.4,
-(4.6) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\Big{\{}\mathcal{H}^{s}(K)\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}:c_{n,t}\in\mathbb{N}\text{ with }\sum_{t=1}^{p}c_{n,t}\leq P\Big{\}}.$
-Note that every inner sum on the right hand side above satisfies
-$\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}$ $=r^{q_{1,n}s}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{(q_{t,n}-q_{1,n})s}$
-$\in\Big{\{}r^{q_{1,n}s}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{\beta s}:-C\leq\beta\leq C\text{ is an integer}\Big{\}}.$ (by (4.5))
-Combining this with (4.6), we see that
-$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\{\mathcal{H}^{s}(K)r^{q_{1,n}s}b:b\in\mathcal{B}\},\quad\forall n\geq 1,$
-where
-$\mathcal{B}:=\Big{\{}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{\beta s}:c_{n,t}\in\mathbb{N},\beta\in\mathbb{Z}\text{ are such that}\sum_{t=1}^{p}c_{n,t}\leq P\text{ and }-C\leq\beta\leq C\Big{\}}$
-is a finite collection. Therefore, we can find (by the pigeonhole principle) a pair of distinct integers $n,\tilde{n}$ such that
+For any $1\leq t\leq p$, let $q_{t,n}$ be the largest positive integer such that $r^{q_{t,n}}\leq\delta^{n}/a_{t}$.
+Since $r_{i}=r$ for all $i\in I$, $I^{\prime}_{t,n}=I^{q_{t,n}}$ (recall (4.3)).
+Then for $t\neq t^{\prime}$, we always have $a_{t}r^{q_{t,n}}\leq\delta^{n}<a_{t^{\prime}}r^{q_{t^{\prime},n}-1}$ and hence $r^{q_{t^{\prime},n}-q_{t,n}}>a_{t}r/a_{t^{\prime}}$.
+So there is a constant $C\geq 1$ such that
+(4.6) $|q_{t,n}-q_{t^{\prime},n}|\leq C,\quad\forall 1\leq t,t^{\prime}\leq p,\forall n\in\mathbb{Z}^{+}.$
+By Lemma 4.3 and Proposition 4.4,
+(4.7) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\Big{\{}\mathcal{H}^{s}(K)\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}:c_{n,t}\in\mathbb{Z}^{+}\cup\{0\}\text{ with }\sum_{t=1}^{p}c_{n,t}\leq P\Big{\}}.$
+Writing
+$\mathcal{B}:=\Big{\{}\sum_{t=1}^{p}c_{t}a_{t}^{s}r^{\beta_{t}s}:c_{t}\in\mathbb{Z}^{+}\cup\{0\},\beta_{t}\in\mathbb{Z}\text{ are such that}\sum_{t=1}^{p}c_{t}\leq P\text{ and}-C\leq\beta_{t}\leq C\Big{\}},$
+we see from (4.6) and (4.7) that
+$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\{\mathcal{H}^{s}(K)r^{q_{1,n}s}b:b\in\mathcal{B}\},\quad\forall n\geq 1.$
+Since $\mathcal{B}$ is a finite collection, we can find (by the pigeonhole principle) a pair of distinct integers $n,\tilde{n}$ such that
 $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))}{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{\tilde{n}}}}))}=r^{(q_{1,n}-q_{1,\tilde{n}})s}$
@@
 This reveals a subtle asymmetry: to upgrade a Lipschitz embedding to a Lipschitz isomorphism, assuming the homogeneity on $K$ is markedly easier than assuming that on $F$.
-It is noteworthy that Proposition 4.4 may relate to a characterization given by Xi [20] via finding appropriate graph-directed subsystems, but we do not know whether the measure-preserving property would directly ensure that.
+It is noteworthy that Proposition 4.4 may relate to a characterization given by Xi [20] via finding appropriate graph-directed subsystems.
+However, we do not know whether the measure-preserving property in Proposition 4.4 would directly ensure the condition in Xi’s criterion (and thereby conclude the equivalence).
 A natural step forward would be to ask how far the optimal map $\widetilde{f}$ deviates from being bi-Lipschitz.
@@
   1. (1)
-    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0$ unless $(t,\bm{i})=(t^{\prime},\bm{i}^{\prime})$;
+    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\approx\mathcal{H}^{s}(F_{\bm{j}})$ for all $1\leq t\leq p$ and $\bm{i}\in\mathcal{I}_{t}$;
   2. (2)
-    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\gtrsim\mathcal{H}^{s}(F_{\bm{j}})$ for all $1\leq t\leq m$ and $\bm{i}\in\mathcal{I}_{t}$.
+    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0$ unless $(t,\bm{i})=(t^{\prime},\bm{i}^{\prime})$.
 Proof.
-For (1), write $E:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})$ and suppose on the contrary that $\mathcal{H}^{s}(E)>0$.
-By Lemma 3.4, we can find $\eta_{k}\in J_{k}$ with $F_{\eta_{k}}\subset F_{\bm{j}}$ such that
-(4.7) $\lim_{k\to\infty}\frac{\mathcal{H}^{s}(E\cap F_{\eta_{k}})}{\mathcal{H}^{s}(F_{\eta_{k}})}=1.$
-For all large $k$, consider $A_{k}:=\widetilde{f}^{-1}(F_{\eta_{k}})$ and let
+The first statement is simple: since $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\subset F_{\bm{j}}$, the $\lesssim$ direction holds trivially; for the inverse direction, one can apply Remark 4.5 to $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ and get
+$\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$ $=\widetilde{c}^{-1}\mathcal{H}^{s}\big{(}\widetilde{f}^{-1}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\big{)}$
+$\geq\widetilde{c}^{-1}\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}}+b_{t})\approx\mathcal{H}^{s}(F_{\bm{j}}),$
+where the last estimate holds because $\lambda_{\bm{j}}\approx a_{t}r_{\bm{i}}$ when $\bm{i}\in I^{\prime}_{t,n}$.
+It remains to verify the second statement. Since $\widetilde{f}:\widetilde{K}\to F$ is a surjection,
+$F_{\bm{j}}=\widetilde{f}\big{(}\widetilde{f}^{-1}(F_{\bm{j}})\big{)}=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}).$
+Let $(t,\bm{i})\neq(t^{\prime},\bm{i}^{\prime})$.
+Write $E:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})$ and suppose on the contrary that $\mathcal{H}^{s}(E)>0$.
+Then $E$ is a Borel subset of the dust-like self-similar set $F_{\bm{j}}$ with positive $\mathcal{H}^{s}$-measure.
+By [6, Lemma 2.12], we can find positive integers $m_{1}<m_{2}<\cdots$, and $\eta_{k}\in J_{m_{k}}$ (with $F_{\eta_{k}}\subset F_{\bm{j}}$) such that
+(4.8) $\mathcal{H}^{s}(E\cap F_{\eta_{k}})>\big{(}1-k^{-1}\big{)}\mathcal{H}^{s}(F_{\eta_{k}}),\quad\forall k\geq 1.$
+For each $k\geq 1$, consider $A_{k}:=\widetilde{f}^{-1}(F_{\eta_{k}})$ and let
 $A_{k,1}:=\widetilde{f}^{-1}(F_{\eta_{k}})\cap(a_{t}O_{t}K_{\bm{i}}+b_{t})\quad\text{and}\quad A_{k,2}:=\widetilde{f}^{-1}(F_{\eta_{k}})\cap(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}).$
-Clearly, $A_{k,1}$, $A_{k,2}$ are both non-empty (since $\mathcal{H}^{s}(E\cap F_{\eta_{k}})>0$), mutually disjoint (since $(t,\bm{i})\neq(t^{\prime},\bm{i}^{\prime})$) and $A_{k,1}\cup A_{k,2}\subset A_{k}$.
-Moreover, it follows from (4.7) that
+Clearly, $A_{k,1}\cap A_{k,2}=\varnothing$ (since $(t,\bm{i})\neq(t^{\prime},\bm{i}^{\prime})$) and $A_{k,1}\cup A_{k,2}\subset A_{k}$. Note that
+(4.9) $\widetilde{f}(A_{k,1})=F_{\eta_{k}}\cap\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
+(because $\widetilde{f}$ is onto and thus $\widetilde{f}(\widetilde{f}^{-1}(A)\cap B)=A\cap\widetilde{f}(B)$ for all $A\subset F$ and $B\subset K$).
+Combining this with $\mathcal{H}^{s}(E\cap F_{\eta_{k}})>0$, we see that $A_{k,1}\neq\varnothing$.
+Similarly, $A_{k,2}\neq\varnothing$.
+Moreover, it follows from (4.8) and (4.9) that
 (4.10) $1\geq\lim_{k\to\infty}\frac{\mathcal{H}^{s}(\widetilde{f}(A_{k,1}))}{\mathcal{H}^{s}(F_{\eta_{k}})}=\lim_{k\to\infty}\frac{\mathcal{H}^{s}(F_{\eta_{k}}\cap\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))}{\mathcal{H}^{s}(F_{\eta_{k}})}\geq\lim_{k\to\infty}\frac{\mathcal{H}^{s}(F_{\eta_{k}}\cap E)}{\mathcal{H}^{s}(F_{\eta_{k}})}=1,$
 that is, (4.10) is an equality.
-On the other hand, by Lemma 4.3, we can write $\widetilde{f}^{-1}(F_{\eta_{k}})=\bigcup_{\ell=1}^{p}\bigcup_{\omega\in\mathcal{I}_{\ell,k}}(a_{\ell}O_{\ell}K_{\omega}+b_{\ell})$ for some $\mathcal{I}_{\ell,k}\subset I^{\prime}_{\ell,k}$.
-So
-$A_{k,1}=\bigcup_{\omega\in\mathcal{I}_{t,k}:K_{\omega}\subset K_{\bm{i}}}(a_{t}O_{t}K_{\omega}+b_{t})\quad\text{and}\quad A_{k,2}=\bigcup_{\omega\in\mathcal{I}_{t^{\prime},k}:K_{\omega}\subset K_{\bm{i}^{\prime}}}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}}).$
+By Lemma 4.3, we can write $\widetilde{f}^{-1}(F_{\eta_{k}})=\bigcup_{\ell=1}^{p}\bigcup_{\omega\in\mathcal{I}_{\ell,m_{k}}}(a_{\ell}O_{\ell}K_{\omega}+b_{\ell})$ for some $\mathcal{I}_{\ell,m_{k}}\subset I^{\prime}_{\ell,m_{k}}$.
+Since $F_{\eta_{k}}\subset F_{\bm{j}}$,
+$A_{k,1}=\bigcup_{\omega\in\mathcal{I}_{t,m_{k}}:K_{\omega}\subset K_{\bm{i}}}(a_{t}O_{t}K_{\omega}+b_{t})\quad\text{and}\quad A_{k,2}=\bigcup_{\omega\in\mathcal{I}_{t^{\prime},m_{k}}:K_{\omega}\subset K_{\bm{i}^{\prime}}}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}}).$
 By Proposition 4.4, $\psi_{\eta_{k}}^{-1}\widetilde{f}^{-1}(F_{\eta_{k}})\in\mathcal{K}$.
-So $\psi_{\eta_{k}}^{-1}(A_{k,1}),\psi_{\eta_{k}}^{-1}(A_{k,2})\in\mathcal{K}$ (because they are sub-unions).
-Also,
-$\mathcal{H}^{s}(A_{k,2})$ $\geq\min_{\omega\in I^{\prime}_{t^{\prime},k}}\mathcal{H}^{s}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}})$
-$\geq(\underline{r}\delta^{k})^{s}\mathcal{H}^{s}(K)$ (recall (4.3))
-$\geq\frac{\underline{r}^{s}\Delta_{F}^{s}}{L^{s}}\cdot\lambda_{\eta_{k}}^{s}\mathcal{H}^{s}(K)$ (since $\eta_{k}\in J_{k}$)
+So $\psi_{\eta_{k}}^{-1}(A_{k,1}),\psi_{\eta_{k}}^{-1}(A_{k,2})\in\mathcal{K}$ (because they are non-empty sub-unions).
+Also
+$\mathcal{H}^{s}(A_{k,2})$ $\geq\min_{\omega\in I^{\prime}_{t^{\prime},m_{k}}}\mathcal{H}^{s}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}})$
+$\geq(\underline{r}\delta^{m_{k}})^{s}\mathcal{H}^{s}(K)$ (recall (4.3))
+$\geq\frac{\underline{r}^{s}\Delta_{F}^{s}}{L^{s}}\cdot\lambda_{\eta_{k}}^{s}\mathcal{H}^{s}(K)$ (since $\eta_{k}\in J_{m_{k}}$)
 $=\frac{\underline{r}^{s}\Delta_{F}^{s}}{L^{s}\mathcal{H}^{s}(F)}\mathcal{H}^{s}(F_{\eta_{k}})\mathcal{H}^{s}(K).$
@@
 $g_{k}:\psi_{\eta_{k}}^{-1}(A_{k,1})$ $\to F$
-$\psi_{\eta_{k}}^{-1}(x)$ $\mapsto\psi_{\eta_{k}}^{-1}\widetilde{f}(x).$
-Since $\psi_{\eta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ and $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$, we can find by Lemma 2.2 a Lipschitz map $g:E\to F$ with $E\in\mathcal{K}$, $\operatorname{Lip}(g)\leq L$ such that
-$\mathcal{H}^{s}(E)\leq\sup_{k}\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}(A_{k,1}))=\mathcal{H}^{s}(F)\cdot\sup_{k}\frac{\mathcal{H}^{s}(A_{k,1})}{\mathcal{H}^{s}(F_{\eta_{k}})}<\widetilde{c}\mathcal{H}^{s}(F)$
+$\psi_{\eta_{k}}^{-1}(x)$ $\mapsto\psi_{\eta_{k}}^{-1}\widetilde{f}(x)$
+with $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$.
+Since $\psi_{\eta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$, we can find by Lemma 2.2 a set $B\in\mathcal{K}$ and a Lipschitz map $g:B\to F$ with $\operatorname{Lip}(g)\leq L$ such that
+$\mathcal{H}^{s}(B)\leq\sup_{k}\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}(A_{k,1}))=\mathcal{H}^{s}(F)\cdot\sup_{k}\frac{\mathcal{H}^{s}(A_{k,1})}{\mathcal{H}^{s}(F_{\eta_{k}})}<\widetilde{c}\mathcal{H}^{s}(F)$
 (where the last inequality is due to (4.11)) and
-$\mathcal{H}^{s}(g(E))$ $\geq\limsup_{k\to\infty}\mathcal{H}^{s}(g_{k}(\psi_{\eta_{k}}^{-1}(A_{k,1})))$
+$\mathcal{H}^{s}(g(B))$ $\geq\limsup_{k\to\infty}\mathcal{H}^{s}(g_{k}(\psi_{\eta_{k}}^{-1}(A_{k,1})))$
 $=\limsup_{k\to\infty}\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}\widetilde{f}(A_{k,1}))$ (by definition)
@@
 $=\mathcal{H}^{s}(F).$ (by (4.10))
-The latter implies that $g(E)=F$. However, by our choice of $\widetilde{K}$ (recall Corollary 4.2),
-$\widetilde{c}\mathcal{H}^{s}(F)>\mathcal{H}^{s}(E)\geq\mathcal{H}^{s}(\widetilde{K})=\widetilde{c}\mathcal{H}^{s}(F),$
-which is a contradiction. This proves (1).
-For (2), let $1\leq t\leq p$, $\bm{i}\in\mathcal{I}_{t}\subset I^{\prime}_{t,n}$ and let
-$B:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\Big{(}\bigcup_{(\bm{i}^{\prime},t^{\prime})\neq(\bm{i},t)}\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})\Big{)},$
-which is a closed set. We have seen in (1) that $\mathcal{H}^{s}(B)=0$. By Remark 4.5, $\mathcal{H}^{s}(\widetilde{f}^{-1}(B))=\widetilde{c}\mathcal{H}^{s}(B)=0$ and
-$\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$ $=\widetilde{c}^{-1}\mathcal{H}^{s}\big{(}\widetilde{f}^{-1}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\big{)}$
-$=\widetilde{c}^{-1}\mathcal{H}^{s}\big{(}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cup\widetilde{f}^{-1}(B)\big{)}$
-$=\widetilde{c}^{-1}\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
-$\gtrsim\mathcal{H}^{s}(F_{\bm{j}}),$
-where the last estimate again comes from (4.3) and Lemma 3.3(1). This proves (2). ∎
+The latter implies that $g(B)=F$, i.e., $g:B\to F$ is a Lipschitz surjection with $\operatorname{Lip}(g)\leq L$.
+However, it follows from $B\in\mathcal{K}$ and our choice of $\widetilde{K}$ (recall Corollary 4.2) that
+$\widetilde{c}\mathcal{H}^{s}(F)>\mathcal{H}^{s}(B)\geq\mathcal{H}^{s}(\widetilde{K})=\widetilde{c}\mathcal{H}^{s}(F),$
+which is a contradiction. This proves (1). ∎
 5. Lipschitz equivalence via the homogeneity of $F$
@@
 Proposition 5.1.
-There exists $\alpha\in\mathbb{Z}^{+}$ such that for all $1\leq t\leq m$, $n\geq 1$ and $\bm{i}\in I^{\prime}_{t,n}$, $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ contains at least one cell in $\{F_{\bm{j}}:\bm{j}\in J_{n+\alpha}\}$.
+There exists $\alpha\in\mathbb{Z}^{+}$ such that for all $1\leq t\leq p$, $n\geq 1$ and $\bm{i}\in I^{\prime}_{t,n}$, $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ contains at least one cell in $\{F_{\bm{j}}:\bm{j}\in J_{n+\alpha}\}$.
 In other words, the Lipschitz surjection $\widetilde{f}$ is not distributed excessively fragmented in $F$.
@@
 5.1. From Proposition 5.1 to Theorem 1.5(1)
-We first explain how Proposition 5.1 leads to the logarithmic commensurability. Assume Proposition 5.1 holds and let $i\in I$ be arbitrary.
+Assume Proposition 5.1 holds and let $i\in I$ be arbitrary.
 For $n\geq 1$, let $k_{n}$ be the unique integer such that $i^{k_{n}}\in I^{\prime}_{1,n}$.
-So $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ (recall (4.3)).
-By Proposition 5.1, $\widetilde{f}(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})\supset F_{\bm{j}_{n}}$ for some $\bm{j}_{n}\in J_{n+\alpha}$.
-Again, we can write as in Lemma 4.3 that
-(5.1) $\widetilde{f}^{-1}(F_{\bm{j}_{n}})=\bigcup_{t=1}^{m}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$
-where $\mathcal{I}_{t}\subset I^{\prime}_{t,n+\alpha}$. Notice that by Lemma 4.7 and our choice of $\bm{j}_{n}$,
-$\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0\;\textrm{ unless }(t^{\prime},\bm{i}^{\prime})=(1,i^{k_{n}}).$
-Thus, from Remark 4.5,
-$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}})\cap(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0\;\textrm{ unless }(t^{\prime},\bm{i}^{\prime})=(1,i^{k_{n}}),$
-which implies that
-$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}}))=\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}})\cap(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})).$
-Combining this with (5.1), we have
-(5.2) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}}))=\mathcal{H}^{s}\Big{(}\bigcup_{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}}(a_{1}O_{1}K_{\bm{i}}+b_{1})\Big{)}.$
+Proposition 5.1 provides a word $\bm{j}_{n}\in J_{n+\alpha}$ with $F_{\bm{j}_{n}}\subset\widetilde{f}(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})$.
+By Lemma 4.3, we can write
+(5.1) $\widetilde{f}^{-1}(F_{\bm{j}_{n}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$
+where $\mathcal{I}_{t}\subset I^{\prime}_{t,n+\alpha}$.
+For $1\leq t\leq p$, if $\mathcal{I}_{t}\not$$\emptyset$ and $\bm{i}\in\mathcal{I}_{t}$, then letting $\eta$ be the prefix of $\bm{i}$ with $\eta\in I^{\prime}_{t,n}$, we have
+$\mathcal{H}^{s}(\widetilde{f}(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})\,\cap\,$ $\widetilde{f}(a_{t}O_{t}K_{\eta}+b_{t}))$
+$\geq\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$
+$=\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$ (by (5.1))
+[…]
+By Lemma 4.7(2), this is impossible unless $t=1$ and $\eta=i^{k_{n}}$. So (5.1) should be
+(5.2) $\widetilde{f}^{-1}(F_{\bm{j}_{n}})=\bigcup_{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}}(a_{1}O_{1}K_{\bm{i}}+b_{1}).$
 Note that $\{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}\}$ can be written as $\{i^{k_{n}}\omega:\omega\in\Gamma_{n}\}$ for some $\Gamma_{n}\subset I^{*}$.
-By definition, $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ and $a_{1}r_{i}^{k_{n}}r_{\omega}\approx\delta^{n+\alpha}$ for all $\omega\in\Gamma_{n}$.
-So there exists a uniform integer $q$ such that $\Gamma_{n}\subset\bigcup_{k=1}^{q}I^{k}$.
-Then by the pigeonhole principle, we can find $n_{1}<n_{2}$ such that $\Gamma_{n_{1}}=\Gamma_{n_{2}}$.
+Since $i^{k_{n}}\in I^{\prime}_{1,n}$ and $\mathcal{I}_{1}\subset I^{\prime}_{1,n+\alpha}$, $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ and $a_{1}r_{i}^{k_{n}}r_{\omega}\approx\delta^{n+\alpha}$ for all $\omega\in\Gamma_{n}$.
+So there exists an integer $q$ independent of $n$ such that $\Gamma_{n}\subset\bigcup_{k=1}^{q}I^{k}$ for all $n$.
+By the pigeonhole principle, we may choose $n_{1}<n_{2}$ such that $\Gamma_{n_{1}}=\Gamma_{n_{2}}$.
 Since $\Psi$ is homogeneous, there are integers $z_{1},z_{2}$ such that $J_{n_{1}+\alpha}=J^{z_{1}}$ and $J_{n_{2}+\alpha}=J^{z_{2}}$.
 Recalling Proposition 4.4 and (5.2), we have
-$\widetilde{c}=\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n_{\ell}}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(a_{1}O_{1}K_{i^{k_{n_{\ell}}}\omega}+b_{1})}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{a_{1}^{s}r_{i}^{k_{n_{\ell}}s}\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(K_{\omega})}{\lambda^{z_{1}s}\mathcal{H}^{s}(F)}$
+$\widetilde{c}=\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n_{\ell}}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(a_{1}O_{1}K_{i^{k_{n_{\ell}}}\omega}+b_{1})}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{a_{1}^{s}r_{i}^{k_{n_{\ell}}s}\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(K_{\omega})}{\lambda^{z_{\ell}s}\mathcal{H}^{s}(F)}$
 for $\ell=1,2$.
 Since $\Gamma_{n_{1}}=\Gamma_{n_{2}}$, we have $r_{i}^{k_{n_{1}}-k_{n_{2}}}=\lambda^{z_{1}-z_{2}}$, i.e., $\frac{\log r_{i}}{\log\lambda}\in\mathbb{Q}$ (choosing $n_{1},n_{2}$ with a large difference, we can require that $z_{1}\neq z_{2}$).
-Since $i$ is arbitrary, this establish Theorem 1.5.
+Since $i$ is arbitrary, Theorem 1.5 is proved.
 5.2. Proof of Proposition 5.1
@@
 For every $k\geq 1$, there exists a threshold number $n(k)$ such that for $n\geq n(k)$, every subset of $\{1,2,\ldots,n\}$ with cardinality larger than $\gamma n$ contains a $k$-term arithmetic progression.
-Since $\Psi$ is homogeneous, for each $n$ we can assign a unique integer $\tau(n)$ such that $J_{n}=J^{\tau(n)}$. For $1\leq t\leq m$ and $n\geq 1$, let
-(5.3) $p(t,n):=\max\{p\geq 1:\exists\bm{i}\in I^{\prime}_{t,n}\text{ such that $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for all $\bm{j}\in J^{\tau(n)+p}$}\}$
-Suppose on the contrary that Proposition 5.1 fails. Then $\max_{1\leq t\leq m}\sup_{n}p(t,n)=\infty$. Without loss of generality, assume that $\sup_{n}p(1,n)=\infty$.
+Since $\Psi$ is homogeneous, for each $n$ we can assign a unique integer $\tau(n)$ such that $J_{n}=J^{\tau(n)}$. For $1\leq t\leq p$ and $n\geq 1$, let
+(5.3) $q(t,n):=\max\{q\geq 1:\exists\bm{i}\in I^{\prime}_{t,n}\text{ such that $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for all $\bm{j}\in J^{\tau(n)+q}$}\}.$
+Suppose on the contrary that Proposition 5.1 fails. Then $\max_{1\leq t\leq p}\sup_{n}q(t,n)=\infty$. Without loss of generality, assume that $\sup_{n}q(1,n)=\infty$.
 Let $\gamma\in(0,1)$ be a fixed constant (will be specified later).
-For $k\geq(\#J)^{6}$, by Szemerédi’s theorem, there is a large integer $n_{k}$ such that any subset of $\{1,\ldots,(\#J)^{p(1,n_{k})}\}$ with cardinality $\geq\gamma(\#J)^{p(1,n_{k})}$ contains a $k$-term arithmetic progression.
-For notational simplicity, in the rest of the proof we write $p_{k}:=p(1,n_{k})$ and enumerate $J^{p_{k}}$ as $\{\eta_{k,t}:1\leq t\leq(\#J)^{p_{k}}\}$ according to the lexicographical order, that is,
-$\eta_{k,1}=\underbrace{1\cdots 1}_{p_{k}\text{ terms}},\,\eta_{k,2}=\underbrace{1\cdots 1}_{p_{k}-1\text{ terms}}\!\!\!\!2,\,\ldots,\,\eta_{k,(\#J)^{p_{k}}}=\underbrace{(\#J)\cdots(\#J)}_{p_{k}\text{ terms}}.$
-It will be helpful to note that for any $\bm{j}\in J^{*}$ with $|\bm{j}|\leq p_{k}$,
-(5.4) $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}=\{z(\#J)^{p_{k}-|\bm{j}|}+1,z(\#J)^{p_{k}-|\bm{j}|}+2,\ldots,(z+1)(\#J)^{p_{k}-|\bm{j}|}\}$
-for some integer $z\geq 0$, i.e., it is a (#J)(\#J)-adic interval of level-$(p_{k}-|\bm{j}|)$.
-By (5.3), one can find a word $\bm{i}_{k}\in I^{\prime}_{1,n_{k}}$ with $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ for all $\bm{j}\in J^{\tau(n_{k})+p_{k}}$.
-On the other hand, recalling Lemma 3.3(3), $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\subset F_{\bm{j}_{k}}$ for some $\bm{j}_{k}\in J_{n_{k}}=J^{\tau(n_{k})}$.
-Since
-$F_{\bm{j}_{k}}=\bigcup_{\theta\in J^{p_{k}}}F_{\bm{j}_{k}\theta}=\bigcup_{1\leq t\leq(\#J)^{p_{k}}}F_{\bm{j}_{k}\eta_{k,t}},$
+For $k\geq(\#J)^{4}$, by Szemerédi’s theorem, there is a large integer $n_{k}$ such that any subset of $\{1,\ldots,(\#J)^{q(1,n_{k})}\}$ with cardinality $\geq\gamma(\#J)^{q(1,n_{k})}$ contains a $k$-term arithmetic progression.
+For notational simplicity, in the rest of the proof we write $q_{k}:=q(1,n_{k})$ and enumerate $J^{q_{k}}$ as $\{\eta_{k,t}:1\leq t\leq(\#J)^{q_{k}}\}$ according to the lexicographical order, that is,
+$\eta_{k,1}=\underbrace{1\cdots 1}_{q_{k}\text{ terms}},\,\eta_{k,2}=\underbrace{1\cdots 1}_{q_{k}-1\text{ terms}}\!\!\!\!2,\,\ldots,\,\eta_{k,(\#J)^{q_{k}}}=\underbrace{(\#J)\cdots(\#J)}_{q_{k}\text{ terms}}.$
+It is helpful to note that for any $\bm{j}\in J^{*}$ with $|\bm{j}|\leq q_{k}$,
+(5.4) $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}=\{z(\#J)^{q_{k}-|\bm{j}|}+1,z(\#J)^{q_{k}-|\bm{j}|}+2,\ldots,(z+1)(\#J)^{q_{k}-|\bm{j}|}\}$
+for some integer $z\geq 0$, i.e., a (#J)(\#J)-adic interval of level-$(q_{k}-|\bm{j}|)$.
+By (5.3), one can find a word $\bm{i}_{k}\in I^{\prime}_{1,n_{k}}$ with
+(5.5) $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1}),\quad\forall\bm{j}\in J^{\tau(n_{k})+q_{k}}.$
+On the other hand, similarly as in (4.4), one can deduce that $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\subset F_{\bm{j}_{k}}$ for some $\bm{j}_{k}\in J_{n_{k}}=J^{\tau(n_{k})}$. Since
+$F_{\bm{j}_{k}}=\bigcup_{\theta\in J^{q_{k}}}F_{\bm{j}_{k}\theta}=\bigcup_{1\leq t\leq(\#J)^{q_{k}}}F_{\bm{j}_{k}\eta_{k,t}},$
 we have
-$\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ $=\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap\bigcup_{1\leq t\leq(\#J)^{p_{k}}}F_{\bm{j}_{k}\eta_{k,t}}$
+$\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ $=\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap\bigcup_{1\leq t\leq(\#J)^{q_{k}}}F_{\bm{j}_{k}\eta_{k,t}}$
 (5.6) $=\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap\bigcup_{t\in T_{k}}F_{\bm{j}_{k}\eta_{k,t}},$
-where $T_{k}:=\{1\leq t\leq(\#J)^{p_{k}}:\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap F_{\bm{j}_{k}\eta_{k,t}}\neq\varnothing\}$. Observe that
+where $T_{k}:=\{1\leq t\leq(\#J)^{q_{k}}:\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap F_{\bm{j}_{k}\eta_{k,t}}\neq\varnothing\}$. Observe that
 $\delta^{n_{k}s}$ $\lesssim\mathcal{H}^{s}(\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1}))$ (since $\bm{i}_{k}\in I^{\prime}_{1,n_{k}}$)
@@
 $=\sum_{t\in T_{k}}\lambda_{\bm{j}_{k}}^{s}\mathcal{H}^{s}(F_{\eta_{k,t}})$
-$\lesssim(\#T_{k})\cdot\delta^{n_{k}s}\cdot\lambda^{p_{k}s}\mathcal{H}^{s}(F)$ (since $\bm{j}_{k}\in J_{n_{k}}$ and $\eta_{k,t}\in J^{p_{k}}$)
-and hence $T_{k}\gtrsim\lambda^{-p_{k}s}=(\#J)^{p_{k}}$.
-In summary, although every cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{p_{k}}\}$ is not entirely contained in $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$, the latter does intersect a large proportion of these cells.
-Picking $0<\gamma<1$ properly at the beginning, we have $T_{k}\geq\gamma(\#J)^{p_{k}}$ for all $k$.
+$\lesssim(\#T_{k})\cdot\delta^{n_{k}s}\cdot\lambda^{q_{k}s}\mathcal{H}^{s}(F)$ (since $\bm{j}_{k}\in J_{n_{k}}$ and $\eta_{k,t}\in J^{q_{k}}$)
+and hence $T_{k}\gtrsim\lambda^{-q_{k}s}=(\#J)^{q_{k}}$.
+In summary, although every cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{q_{k}}\}$ is not entirely contained in $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$, the latter does intersect a large proportion of these cells.
+Picking $0<\gamma<1$ properly at the beginning, we have $T_{k}\geq\gamma(\#J)^{q_{k}}$ for all $k$.
 As a consequence, $T_{k}$ should contain a $k$-term arithmetic progression, say $AP_{k}$.
-For the purpose of blowing up on a good sub-cell of $F_{\bm{j}_{k}}$, we perform the following truncation: pick $\theta_{k}\in J^{*}$ such that $F_{\bm{j}_{k}\theta_{k}}$ is a largest cell contained entirely in $\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\bm{j}_{k}\eta_{k,t}}$ and set
+Pick $\theta_{k}\in J^{*}$ such that $F_{\theta_{k}}$ is a largest cell contained entirely in $\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\eta_{k,t}}$ (so $|\theta_{k}|\leq q_{k}$) and set
 (5.7) $AP^{\prime}_{k}:=\{t\in AP_{k}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\},$
-which is a subprogression of $AP_{k}$. By (5.4), $\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}$ is simply a #J\#J-adic interval of the form
-(5.7) $\{z_{k}(\#J)^{p_{k}-|\theta_{k}|}+1,z_{k}(\#J)^{p_{k}-|\theta_{k}|}+2,\ldots,(z_{k}+1)(\#J)^{p_{k}-|\theta_{k}|}\}$
+which is a subprogression of $AP_{k}$. By (5.4), $\{1\leq t\leq(\#J)^{q_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}$ is simply a #J\#J-adic interval of the form
+(5.8) $\{z_{k}(\#J)^{q_{k}-|\theta_{k}|}+1,z_{k}(\#J)^{q_{k}-|\theta_{k}|}+2,\ldots,(z_{k}+1)(\#J)^{q_{k}-|\theta_{k}|}\}$
 for some $z_{k}\geq 0$. Write $d_{k}$ to be the common difference of $AP_{k}$. Clearly,
-(5.8) $AP^{\prime}_{k}\geq\frac{\#\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{d_{k}}-2=\frac{(\#J)^{p_{k}-|\theta_{k}|}}{d_{k}}-2.$
-It is straightforward to obtain the following lower bound estimate of $AP_{k}^{\prime}$.
+(5.9) $AP^{\prime}_{k}\geq\frac{\#\{1\leq t\leq(\#J)^{q_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{d_{k}}-1=\frac{(\#J)^{q_{k}-|\theta_{k}|}}{d_{k}}-1.$
 Lemma 5.3.
-$AP^{\prime}_{k}\geq(\#J)^{-4}k$.
+$AP^{\prime}_{k}\geq(\#J)^{-3}k$.
 Proof.
 Write $\ell_{k}:=\lfloor\frac{\log(k-1)d_{k}}{\log\#J}\rfloor$. Then
-$\max AP_{k}-\min AP_{k}\geq(\#AP_{k}-1)d_{k}=(k-1)d_{k}\geq(\#J)^{\ell_{k}}>3\cdot(\#J)^{\ell_{k}-2}.$
-Therefore, $\min AP_{k},\max AP_{k}]$ must contain a #J\#J-adic interval
-$\{z(\#J)^{\ell_{k}-2}+1,\ldots,(z+1)(\#J)^{\ell_{k}-2}\}$
+$\max AP_{k}-\min AP_{k}=(\#AP_{k}-1)d_{k}=(k-1)d_{k}\geq(\#J)^{\ell_{k}}\geq 2\cdot(\#J)^{\ell_{k}-1}.$
+Therefore, the interval $\min AP_{k},\max AP_{k}]$ must contain a #J\#J-adic interval of the form
+$\big{\{}z(\#J)^{\ell_{k}-1}+1,\ldots,(z+1)(\#J)^{\ell_{k}-1}\big{\}}$
 for some $z\geq 0$.
-By (5.4), this is just $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}$ for some $\bm{j}\in J^{p_{k}-(\ell_{k}-2)}$.
+By (5.4), this set corresponds to $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}$ for some $\bm{j}\in J^{q_{k}-(\ell_{k}-1)}$.
 In particular, $F_{\bm{j}}\subset\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\eta_{k,t}}$.
-Recalling our choice of $\theta_{k}$ (above (5.6)), we have $|\theta_{k}|\leq|\bm{j}|=p_{k}-(\ell_{k}-2)$ and hence
-$\#AP^{\prime}_{k}$ $\geq\frac{(\#J)^{\ell_{k}-2}-2d_{k}}{d_{k}}$ (by (5.8))
-$=(\#J)^{-2}\cdot(\#J)^{\lfloor\frac{\log(k-1)d_{k}}{\log\#J}\rfloor}\cdot d_{k}^{-1}-2$
-$\geq(\#J)^{-3}(k-1)d_{k}\cdot d_{k}^{-1}-2$
-$>k(\#J)^{-3}-3\geq k(\#J)^{-4},$
-where the last inequality holds since $k\geq(\#J)^{6}$. ∎
+Recalling the choice of $\theta_{k}$ (above (5.7)), we have $|\theta_{k}|\leq|\bm{j}|=q_{k}-(\ell_{k}-1)$ and it follows from (5.9) that
+$\#AP^{\prime}_{k}$ $\geq\frac{(\#J)^{\ell_{k}-1}}{d_{k}}-1$
+$=(\#J)^{-1}\cdot(\#J)^{\lfloor\frac{\log(k-1)d_{k}}{\log\#J}\rfloor}\cdot d_{k}^{-1}-1$
+$\geq(\#J)^{-2}(k-1)d_{k}\cdot d_{k}^{-1}-1$
+$>k(\#J)^{-2}-2,$
+which is at least $k(\#J)^{-3}$ because $k\geq(\#J)^{4}$. ∎
 To derive a contradiction, write
@@
 Proof.
-For (1), it follows from Lemma 5.3 that
-$d_{k}$ $\leq\frac{\#\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{\#AP^{\prime}_{k}-1}$ (by (5.6))
-$=\frac{(\#J)^{p_{k}-|\theta_{k}|}}{\#AP^{\prime}_{k}-1}$ (by (5.7))
-$\leq(\#J)^{p_{k}-|\theta_{k}|-\lfloor\frac{\log k}{\log\#J}\rfloor-c}$
-[…]
-$\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})\leq\lambda^{|\bm{j}_{k}|+p_{k}-\widetilde{p}_{k}}=\lambda_{\bm{j}_{k}}\cdot\lambda^{|\theta_{k}|+\lfloor\frac{\log k}{\log\#J}\rfloor+c}=\lambda_{\bm{j}_{k}\theta_{k}}\cdot\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor+c},$
+For (1), write $\widetilde{q}_{k}:=q_{k}-|\theta_{k}|-\lfloor\frac{\log k}{\log\#J}\rfloor+4$. Since $k\geq(\#J)^{4}$, Lemma 5.3 implies
+$AP_{k}^{\prime}-1\geq(\#J)^{-1}\cdot\#AP_{k}^{\prime}\geq(\#J)^{-4}k\geq(\#J)^{\lfloor\frac{\log k}{\log\#J}\rfloor-4}$
+and hence
+$d_{k}$ $\leq\frac{\#\{1\leq t\leq(\#J)^{q_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{\#AP^{\prime}_{k}-1}$ (by (5.7))
+$=\frac{(\#J)^{q_{k}-|\theta_{k}|}}{\#AP^{\prime}_{k}-1}$ (by (5.8))
+$\leq(\#J)^{q_{k}-|\theta_{k}|-\lfloor\frac{\log k}{\log\#J}\rfloor+4}=(\#J)^{\tilde{q}_{k}}.$
+So $AP_{k}$ should intersect every block of $(\#J)^{\widetilde{q}_{k}}$ consecutive integers in $\min AP_{k},\max AP_{k}]$ (and in particular, every #J\#J-adic interval of level-$(q_{k}-\widetilde{q}_{k})$ in $\min AP_{k},\max AP_{k}]$).
+Since $AP_{k}\subset T_{k}$, by the definition of $T_{k}$ and (5.4), $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ (and hence $\widetilde{f}(A_{k,1})$) should intersect every level-$(|\bm{j}_{k}|+q_{k}-\widetilde{q}_{k})$ cell contained in $F_{\bm{j}_{k}\theta_{k}}$.
+From this, we deduce that
+$\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})\leq\lambda^{|\bm{j}_{k}|+q_{k}-\widetilde{q}_{k}}=\lambda_{\bm{j}_{k}}\cdot\lambda^{|\theta_{k}|+\lfloor\frac{\log k}{\log\#J}\rfloor-4}=\lambda_{\bm{j}_{k}\theta_{k}}\cdot\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor-4},$
 where $\operatorname{d_{H}}$ denotes the Hausdorff distance. Then
-$\operatorname{d_{H}}(\psi_{\bm{j}_{k}\theta_{k}}\widetilde{f}(A_{k,1}),F)=\lambda_{\bm{j}_{k}\theta_{k}}^{-1}\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})=\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor+c},$
+$\operatorname{d_{H}}(\psi^{-1}_{\bm{j}_{k}\theta_{k}}\widetilde{f}(A_{k,1}),F)=\lambda_{\bm{j}_{k}\theta_{k}}^{-1}\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})=\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor-4},$
 which vanishes as $k\to\infty$. This proves (1).
-For (2), recall that $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ does not contain any cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{p_{k}}\}$ and $|\theta_{k}|\leq p_{k}$.
-So there are $1\leq t\leq m$, $\bm{i}^{\prime}_{k}\in I^{\prime}_{t,n_{k}}$ with $(t,\bm{i}^{\prime}_{k})\neq(1,\bm{i}_{k})$ such that
-$\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}}+b_{t})\cap F_{\bm{j}_{k}\theta_{k}}\neq\varnothing.$
-Pick  [^1] For general $\delta$ the word $\bm{j}_{k}\theta_{k}$ might not belong to any $J_{n}$.
-But one can simply pick $\delta=\lambda$ at the beginning to fix this problem.
-an integer $q_{k}\in\mathbb{Z}^{+}$ so that $\bm{j}_{k}\theta_{k}\in J_{q_{k}}$.
-In the spirit of Lemma 4.3, there exists a word $\omega_{k}$ such that $\bm{i}^{\prime}_{k}\omega_{k}\in I^{\prime}_{t,q_{k}}$ and $\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\subset F_{\bm{j}_{k}\theta_{k}}$.
+For (2), note that there are $1\leq t\leq p$, $\bm{i}^{\prime}_{k}\in I^{\prime}_{t,n_{k}}$ with $(t,\bm{i}^{\prime}_{k})\neq(1,\bm{i}_{k})$ such that
+$\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}}+b_{t})\cap F_{\bm{j}_{k}\theta_{k}}\neq\varnothing;$
+otherwise, $F_{\bm{j}_{k}\theta_{k}}\subset\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$, which contradicts (5.5) because $|\theta_{k}|\leq q_{k}$.
+Pick any point $x\in a_{t}O_{t}K_{\bm{i}^{\prime}_{k}}+b_{t}$ with $\widetilde{f}(x)\in F_{\bm{j}_{k}\theta_{k}}$.
+Similarly as in (4.4), we can find a word $\omega_{k}\in I^{*}$ such that $x\in a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t}$, $r_{\bm{i}^{\prime}_{k}\omega_{k}}\approx\lambda_{\bm{j}_{k}\theta_{k}}$ and $\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\subset F_{\bm{j}_{k}\theta_{k}}$.
 Note that
-(5.10) $\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\approx\delta^{q_{k}s}\approx\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})=\widetilde{c}^{-1}\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}})),$
+(5.11) $\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\approx\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})=\widetilde{c}^{-1}\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}})),$
 where the last equality comes from Proposition 4.4. Then
@@
 $\leq(1-c_{*})\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}}))$ (by (5.11))
-for some $0<c_{*}<1$ independent of $k$. This proves (2). ∎
+for some $0<c_{*}<1$ independent of $k$, which completes the proof. ∎
 Now we are ready to obtain a contradiction. Again, each Lipschitz inclusion $\widetilde{f}:A_{k,1}\to F_{\bm{j}_{k}\theta_{k}}$ naturally generates a blow-up map
@@
 with $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$.
-Similar to the second paragraph of the proof of Lemma 4.7, we have $\psi_{\bm{j}_{k}\theta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ for all $k$.
+Similar to the fourth paragraph of the proof of Lemma 4.7, we have $\psi_{\bm{j}_{k}\theta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ for all $k$.
 Then, combining Lemma 2.2, Remark 2.3 and Lemma 5.4(1), we obtain a Lipschitz surjection $g:E\to F$ for some $E\in\mathcal{K}$ such that $\operatorname{Lip}(g)\leq L$ and
@@
 $\leq\mathcal{H}^{s}(F)\cdot(1-c_{*})\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}}))}{\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})}$ (by Lemma 5.4(2))
+$<\widetilde{c}\mathcal{H}^{s}(F)$ (by Proposition 4.4)
 […]
@@
 In particular, $\mu(K_{\bm{i}})=r_{\bm{i}}^{s}$ for all $\bm{i}\in I^{*}$.
-To state a mass decomposition lemma in [22], write $\mathbb{Z}[r^{s}]$ be the ring generated by $r^{s}$ and the integer set $\mathbb{Z}$.
-With a slight abuse of notation, here we pick $\delta=r$ and let
-$\left\{\begin{array}[]{l}I_{n}:=\{\bm{i}\in I^{*}:r_{\bm{i}}\leq\delta^{n}<r_{\bm{i}^{-}}\},\\ J_{n}:=\{\bm{j}\in J^{*}:\lambda_{\bm{j}}\leq\delta^{n}<\lambda_{\bm{j}^{-}}\}.\end{array}\right$.
+To state a mass decomposition lemma in [22], let $\mathbb{Z}[r^{s}]$ be the ring generated by $r^{s}$ and the integer set $\mathbb{Z}$. With a slight abuse of notation, here we pick $\delta=r$ and let
+(6.1) $\left\{\begin{array}[]{l}I_{n}:=\{\bm{i}\in I^{*}:r_{\bm{i}}\leq\delta^{n}<r_{\bm{i}^{-}}\},\\ J_{n}:=\{\bm{j}\in J^{*}:\lambda_{\bm{j}}\leq\delta^{n}<\lambda_{\bm{j}^{-}}\}.\end{array}\right$.
 Lemma 6.1 ([22, Lemma 3.4]).
@@
 More precisely, we can decompose $I^{3}=\bigcup_{j\in J}\tilde{I}_{j}$ such that $\sum_{\bm{i}\in\tilde{I}_{j}}r_{\bm{i}}^{s}=\lambda_{j}^{s}$ for each $j\in J$.
-In this example, the ratio root $r$ equals $3^{-1}$. So $\delta=3^{-1}$ and $J_{1}=J$. Write
+In this example, the ratio root $r$ equals $3^{-1}$, so $\delta=3^{-1}$ and $J_{1}=J$. Write
 $\Omega:=\{1,3^{-s},3^{-2s},\ldots,3^{-5s}\}.$
 For such parameters $r$ and $\Omega$, there exists a large integer $c$ as specified in Lemma 6.1.
-We claim that for every $k\geq 1$, there exists a partition $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$ such that $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1+(k-1)c}$.
-The claim can be proved by induction.
-Clearly, picking $c$ large enough at the beginning, every cell in $\{K_{\bm{i}}:\bm{i}\in I^{3}\}$ is a finite union of cells in $\{K_{\omega}:\omega\in I_{1+c}\}$.
-So we can transform the previous partition of $I^{3}$ into another partition $I_{1+c}=\bigcup_{\bm{j}\in J_{1}}\tilde{I}_{\bm{j}}$ with $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1}$.
+We claim that for every $k\geq 1$, there exists a partition $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$ such that $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1+(k-1)c}$, where $I_{n}$ and $J_{n}$ are as defined in (6.1).
+The claim is proved by induction.
+Clearly, picking $c$ large enough initially, every cell in $\{K_{\bm{i}}:\bm{i}\in I^{3}\}$ can be expressed as a finite union of cells in $\{K_{\omega}:\omega\in I_{1+c}\}$.
+Since $J_{1}=J$, we can transform the previous partition of $I^{3}$ into another partition $I_{1+c}=\bigcup_{\bm{j}\in J_{1}}\tilde{I}_{\bm{j}}$ with $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1}$.
 This settles the case $k=1$.
@@
 (6.2) $\lambda_{\eta}^{s}\in 3^{-(1+kc)s}\{1,3^{-s},\ldots,3^{-5s}\}=3^{-(1+kc)s}\Omega=r^{(1+kc)s}\Omega,\quad\forall\eta\in J_{1+kc}.$
-Let $\bm{j}\in J_{1+(k-1)c}$. Note that by the induction hypothesis, $\bm{j}$ corresponds to some $\tilde{I}_{\bm{j}}\subset I_{1+kc}$ with
+Let $\bm{j}\in J_{1+(k-1)c}$. By the induction hypothesis, $\bm{j}$ corresponds to some $\tilde{I}_{\bm{j}}\subset I_{1+kc}$ with
 $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}=\sum_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\lambda_{\eta}^{s},$
 where $\bm{j}\prec\eta$ means that $\bm{j}$ is a prefix of $\eta$.
-In other words, writing $\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}, EE$ is simply an element in $\mathcal{B}_{1+kc}$ (as in Lemma 5.4) with $\mu(E)=\sum_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\lambda_{\eta}^{s}$, where each $\lambda_{\eta}^{s}$ is a member of $r^{(1+kc)s}\Omega$ (recall (6.1)).
-It then follows from Lemma 6.1 that there exists a partition $E=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}$, where every $E_{\eta}$ is a member of $\mathcal{B}_{1+(k+1)c}$ with $\mu(E_{\eta})=\lambda_{\eta}^{s}$.
-By definition, there is some $\tilde{I}_{\eta}\subset I_{1+(k+1)c}$ such that $E_{\eta}=\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}$.
+In other words, writing $E_{\bm{j}}:=\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}$, $E_{\bm{j}}$ is simply an element in $\mathcal{B}_{1+kc}$ (as in Lemma 5.4) with $\mu(E_{\bm{j}})=\sum_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\lambda_{\eta}^{s}$, where each $\lambda_{\eta}^{s}$ is a member of $r^{(1+kc)s}\Omega$ (recall (6.2)).
+It then follows from Lemma 6.1 that there exists a partition $E_{\bm{j}}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}$, where every $E_{\eta}$ is a member of $\mathcal{B}_{1+(k+1)c}$ with $\mu(E_{\eta})=\lambda_{\eta}^{s}$.
+By definition, there is some $\tilde{I}_{\eta}\subset I_{1+(k+1)c}$ such that $E_{\eta}=\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi}$.
 We conclude that
-(6.2) $\bigcup_{\omega\in\tilde{I}_{\bm{j}}}K_{\omega}=E=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa},$
+(6.3) $\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}=E_{\bm{j}}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
 where each equality represents a partition. It follows that
-$\bigcup_{\kappa\in I_{1+(k+1)c}}K_{\kappa}=K$ $=\bigcup_{\omega\in I_{1+kc}}K_{\omega}$
-$=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\omega\in\tilde{I}_{\bm{j}}}K_{\omega}$ (since $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$)
-$=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}$ (by (6.2))
-$=\bigcup_{\eta\in J_{1+kc}}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa},$
+(6.4) $\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi}=\bigcup_{\eta\in J_{1+kc}}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
+because $J_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\{\eta\in J_{1+kc}:\bm{j}\prec\eta\}$.
+On the other hand, since $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$ (the induction hypothesis),
+$\bigcup_{\bm{i}\in I_{1+kc}}K_{\bm{i}}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}.$
+Combining this with (6.4), we have
+$\bigcup_{\xi\in I_{1+(k+1)c}}K_{\xi}=K=\bigcup_{\bm{i}\in I_{1+kc}}K_{\bm{i}}=\bigcup_{\eta\in J_{1+kc}}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
 which in turn induces a partition (since each $\tilde{I}_{\eta}\subset I_{1+(k+1)c}$)
 $I_{1+(k+1)c}=\bigcup_{\eta\in J_{1+kc}}\tilde{I}_{\eta}$
-such that $\sum_{\kappa\in\tilde{I}_{\eta}}r_{\kappa}^{s}=\mu(E_{\eta})=\lambda_{\eta}^{s}$ for all $\eta\in J_{1+kc}$. This completes the induction.
-The mass distribution in our claim naturally induces a Lipschitz surjection from $K$ to $F$ as follows.
+such that $\sum_{\xi\in\tilde{I}_{\eta}}r_{\xi}^{s}=\mu(E_{\eta})=\lambda_{\eta}^{s}$ for all $\eta\in J_{1+kc}$. This completes the induction.
+The claim naturally induces a Lipschitz surjection from $K$ to $F$ as follows.
 For $x\in K$ and $k\geq 1$, let $\bm{i}_{k}$ be the unique word in $I_{1+kc}$ such that $x\in K_{\bm{i}_{k}}$.
@@
 By our claim, there exists $\bm{j}_{k}\in J_{1+(k-1)c}$ such that $\bm{i}_{k}\in\tilde{I}_{\bm{j}_{k}}$.
-Moreover, it follows from (6.2) that
-$K_{\bm{i}_{k+1}}\subset K_{\bm{i}_{k}}\subset\bigcup_{\omega\in\tilde{I}_{\bm{j}_{k}}}K_{\omega}=\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}.$
-In other words, $\bm{i}_{k+1}\in\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\tilde{I}_{\eta}$.
-So $\bm{j}_{k}\prec\bm{j}_{k+1}$, that is, $\bm{j}_{1}\prec\bm{j}_{2}\prec\cdots$ as well.
-Then we just set $g(x)$ to be the single point in $\bigcap_{k}F_{\bm{j}_{k}}$.
-In this way, it is not hard to see that $g(K)=F$.
+Moreover,
+(6.5) $K_{\bm{i}_{k+1}}\subset K_{\bm{i}_{k}}\subset\bigcup_{\omega\in\tilde{I}_{\bm{j}_{k}}}K_{\omega}=\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
+where the last equality follows from (6.3). Since $\bm{i}_{k+1}\in I_{1+(k+1)c}$ and $\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\tilde{I}_{\eta}\subset I_{1+(k+1)c}$, (6.5) implies that
+$\bm{i}_{k+1}\in\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\tilde{I}_{\eta}.$
+Combining this with our choice of $\bm{j}_{k+1}$ (i.e., the word in $J_{1+kc}$ such that $\bm{i}_{k+1}\in\widetilde{I}_{{\bm{j}}_{k+1}}$), we see that $\bm{j}_{k+1}\in\{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta\}$.
+In particular, $\bm{j}_{k}\prec\bm{j}_{k+1}$.
+So $\bm{j}_{1}\prec\bm{j}_{2}\prec\cdots$ as well and we set $g(x)$ to be the single point in $\bigcap_{k}F_{\bm{j}_{k}}$.
+To see the surjectivity of $g$, let $y\in F$ be arbitrary.
+For each $k\geq 1$, let $\xi_{k}$ be the unique word in $\widetilde{J}_{1+(k-1)c}$ such that $y\in F_{\xi_{k}}$.
+From the previous induction argument, it follows that $B_{k}:=\bigcup_{\bm{i}\in\widetilde{I}_{\xi_{k}}}K_{\bm{i}}$ is non-empty, compact and nested decreasing in $k$ (see (6.3)).
+Moreover, $g$ maps the non-empty intersection $\bigcap_{k=1}^{\infty}B_{k}$ to the singleton $\{y\}$.
+So $g(K)=F$.
 It remains to prove that $g$ is Lipschitz.
@@
 So $g$ is Lipschitz.
-Acknowledgements. The research of the authors is partially supported by NSFC grant 12371089, and the Fundamental Research Funds for the Central Universities of China grant 2024FZZX02-01-01.
+Acknowledgements.
+The research of the authors is partially supported by NSFC grants 12371089, 12501113, 12531004, and the Fundamental Research Funds for the Central Universities of China grant 2024FZZX02-01-01.